### v1

# Design-Based Inference under Random Potential Outcomes via Riesz Representation

Abstract
We introduce a general framework for design-based causal inference that accommodates stochastic potential outcomes, thereby extending the classical Neyman-Rubin setup in which outcomes are treated as fixed. In our formulation, each unit’s potential outcome is modelled as a function $\tilde{y}_{i}(z,\omega)$, where $\omega$ denotes latent randomness external to the treatment assignment. Building on recent work that connects design-based estimation with the Riesz representation theorem, we construct causal estimators by embedding potential outcomes in a Hilbert space and defining treatment effects as linear functionals. This allows us to derive unbiased and consistent estimators, even when potential outcomes exhibit random variation. The framework retains the key advantage of design-based analysis, namely, the use of a known randomisation scheme for identification, while enabling inference in settings with inherent stochasticity. We establish large-sample properties under local dependence, provide a variance estimator compatible with sparse dependency structures, and illustrate the method through a simulation. Our results unify design-based reasoning with random-outcome modelling, broadening the applicability of causal inference in complex experimental environments.
1 Introduction
Randomised experiments remain central to the identification of causal effects. In the classical design-based perspective, pioneered by Neyman (1990) and Rubin (1974), potential outcomes are treated as fixed quantities, with all randomness attributed to the treatment assignment mechanism. This approach has given rise to influential estimators such as the Horvitz-Thompson estimator, which rely solely on the experimental design for identification.
Building on this foundation, several strands of recent literature have extended design-based inference to more complex settings. In particular, a substantial body of work addresses interference and dependence among units, including network and spatial designs (Aronow and Samii, 2017; Sävje et al., 2021; Athey et al., 2021). These methods typically retain the fixed-potential-outcomes assumption and model dependence through exposure mappings or stratified interference structures. Other related contributions focus on inference under model misspecification or robust estimation in observational studies (Abadie et al., 2020; Imbens, 2004), but rely on outcome modelling rather than randomisation for identification.
Recent contributions, notably by Harshaw, Wang, and Sävje (2022, hereafter HWS), have reframed this design-based approach using the Riesz representation theorem. By modelling potential outcomes as elements of a Hilbert space induced by the design, and defining treatment effects as linear functionals on that space, they construct a unified framework that encompasses standard weighting estimators and extends naturally to settings with interference or continuous treatments.
A limitation of this framework, however, is that it assumes potential outcomes are deterministic. In many applied settings, outcomes exhibit stochastic variation even under fixed treatment assignments. For instance, in biological experiments or sensor-based field studies, outcomes may be subject to measurement noise or random physiological responses. This variability can be naturally modelled via a latent random element $\omega$ that affects the outcome but not the treatment assignment. In such contexts, it is more appropriate to model potential outcomes as random functions of a latent variable $\omega$, while maintaining the integrity of the randomisation mechanism.
This paper introduces a design-based framework that accommodates random potential outcomes, thereby generalising HWS’s approach. Unlike HWS, who assume fixed potential outcomes, we allow them to be random conditional on a latent variable $\omega$. This extension is motivated both by practical and conceptual considerations. In many applied settings, such as biological experiments or sensor-based field studies, outcomes exhibit intrinsic randomness even under identical treatment assignments—due, for example, to measurement noise or stochastic physiological responses. More fundamentally, the very notion of an average treatment effect implicitly assumes a distribution over units or outcome-generating processes, suggesting that potential outcomes should be treated as random quantities rather than fixed values. Our formulation makes this randomness explicit while preserving the design-based logic of identification through randomised treatment assignment. We retain the assumption that the treatment vector is drawn from a known design distribution $\mu$, and model each potential outcome as a function $y_{i}(z,\omega)$. The Riesz representation theorem continues to play a central role: it ensures that, under minimal regularity conditions, treatment effects can still be represented as inner products with a unique representer function, even in the presence of outcome-level randomness.
Our analysis is confined to experimental settings with known randomisation schemes. We do not consider observational designs, where unconfoundedness must be justified via covariate adjustment. The framework allows for spillover effects and continuous treatments.
The main contribution of this paper is to extend the Riesz representation framework for design-based causal inference to settings where potential outcomes are themselves random functions. By embedding these outcome functions in a Hilbert space defined over both treatment assignments and latent randomness, we show that treatment effects can still be expressed as continuous linear functionals and estimated accordingly. The resulting Riesz estimator remains unbiased and consistent under local dependence assumptions and exhibits desirable asymptotic properties. In addition, we develop a variance estimation method that accommodates sparse dependency structures, such as those arising from network interference. Finally, we discuss practical implementation using basis expansions and illustrate the method through a simulation study.
The paper is organised as follows. Section 2 introduces the model and basic assumptions. Section 3 establishes large-sample properties of the estimator under local dependence. Section 4 presents a feasible variance estimator and discusses its consistency. Section 5 covers computational aspects of estimating the Riesz representer in both finite- and infinite-dimensional settings. Section 6 provides a simulation study. Section 7 concludes with a discussion of the broader implications and possible extensions.
2 Model Setup
2.1 The Potential Outcomes
We begin by considering a setting with $n$ units, where
Assumption 1 (The Setting).
each unit’s potential outcome is modelled as a general measurable function $y_{i}(z,x_{i},\epsilon_{i})$ with the following arguments: $z=(z_{1},\dots,z_{n})$ denoting the vector of treatment assignments; $x_{i}=x_{i}(\omega)$ representing the possibly observed covariates for unit $i$; $\epsilon_{i}=\epsilon_{i}(\omega)$ denoting the idiosyncratic error or unobserved heterogeneity for unit $i$. Here, $\omega\in\Omega$ is a latent random element defined on a common probability space $(\Omega,\mathcal{F}_{\omega},P)$ shared by all units.
Furthermore, we define the structural mapping from the latent space $\Omega$ to variables $(x_{i},\epsilon_{i})$ and represent the potential outcome via the composition
$\tilde{y}_{i}(z,\omega)=y_{i}(z,x_{i}(\omega),\epsilon_{i}(\omega)),$ (2.1)
since $x_{i}$ and $\epsilon_{i}$ are measurable functions of $\omega$.
Although $\omega$ is not unit-specific, its structure permits different subcomponents to influence different units, thereby being able to preserve unit-specific variability and independence and accommodating heterogeneous potential outcomes.
The potential outcome for unit $i$ depends on the entire treatment vector $z$, so that interference (or spillover effects) is permitted. Notably, one may also consider the special case of no spillover effects, whereby
$y_{i}(z,x_{i},\epsilon_{i})=y_{i}(z_{i},x_{i},\epsilon_{i}),$ (2.2)
which corresponds to a version of the Stable Unit Treatment Value Assumption (SUTVA) incorporating consistency, as introduced by Rubin (1980). In the present work we do not impose the no-spillover condition a priori; rather, we allow for general interference and propose to test for its presence.
The structural mapping (2.1) is the primary object of interest in what follows.
Denoting $\mathcal{Z}$ as the treatment assignment space, equipped with its Borel sigma-algebra $\mathcal{F}_{z}$, we presume that the structural mapping $\tilde{y}_{i}:\mathcal{Z}\times\Omega\to\mathbb{R}$ is measurable with respect to the sigma-algebra $\mathcal{F}_{z}\otimes\mathcal{F}_{\omega}$ on the domain and the Borel sigma-algebra $\mathcal{B}(\mathbb{R})$ on the codomain. This ensures that the relevant random variables and their expectations are well-defined, and it also serves as a prerequisite for defining inner products and applying Hilbert space tools in the sequel. Note that the measurability of the covariates is sometimes required to ensure that conditional expectations and probabilities given $x_{i}$ are well-defined within the underlying probability space, although we do not explicitly rely on this property in the present paper.
Following randomisation principles, we assume that
Assumption 2 (Randomisation).
the treatment assignment vector $z$ is drawn from a known randomisation distribution that is independent of the outcome-generating process given the latent variable $\omega$.
Assumption 2 is not required for the unbiasedness or asymptotic validity of the Riesz estimator. Its purpose is operational rather than inferential: it ensures that the design distribution $\mu$ remains valid when conditioning on $\omega$, which is essential for defining the Riesz representer via the inner product induced by $\mu$. Once the representer $\psi_{i}$ is constructed, all subsequent inference, including estimation and asymptotic analysis, proceeds without requiring independence between $z$ and $\omega$. This assumption will be invoked again in Section 5, where the Gram matrix is computed in order to estimate the Riesz representer of the potential outcomes.
It is worth noting that this condition corresponds to the assumption of ignorability or unconfoundedness, as articulated in Rubin (1978) and further developed in Rosenbaum and Rubin (1983) for observational studies. Under this formulation, $\omega$ governs all outcome-related stochasticity across units, while the randomisation mechanism for $z$ remains externally controlled and separate (see Neyman (1990) and Fisher (1935)).
Conceptually, once a particular realisation of $\omega$ is fixed, each unit’s potential outcome function $\tilde{y}_{i}(z,\omega)$ becomes deterministic in $z$; hence, the only source of randomness in the observed outcomes from the randomised assignment of treatments. This approach preserves the fundamental idea that each unit has a well-defined potential outcome under every treatment assignment—capturing the core principle of SUTVA (or, more generally, the notion of well-defined potential outcomes)—even as we permit these outcomes to vary stochastically across realisations of $\omega$.
This formulation is flexible and expressive in that it captures both unconfoundedness and confounding, depending on the relationship between $z$ and the latent variable $\omega$. If $z$ and $\omega$ are independent, then $\tilde{y}_{i}(z^{0},\omega)$ and $\tilde{y}_{i}(z^{1},\omega)$ are independent of $z$. More generally, if $z$ and $\omega$ are independent given $x\in X_{0}$, then $y_{i}(z^{0},x,\epsilon_{i})$ and $y_{i}(z^{1},x,\epsilon_{i})$ are independent of $z\mid x\in X_{0}$, for any measurable $X_{0}\subset\mathcal{F}_{x}:=\{A\subseteq x_{i}(\Omega):x_{i}^{-1}(A)\in\mathcal{F}_{\omega}\}$. In randomized experiments, the design ensures that $z$ is independent of $(x_{i},\epsilon_{i})$, mitigating confounding. In observational studies, however, it is imperative to adjust for $x_{i}$ to control for confounding. In either case, this framework models potential outcomes as random rather than fixed.
2.2 Model Space and Inner Product
Let $L^{p}(\mathcal{Z}\times\Omega)$ denote the space of all measurable functions $u:\mathcal{Z}\times\Omega\to\mathbb{R}$, satisfying $\mathbb{E}[|u(z,\omega)|^{p}]<\infty$, for some integer $p\geq 1$. For notational convenience, we henceforth write $L^{p}$ in place of $L^{p}(\mathcal{Z}\times\Omega)$, whenever the meaning is clear from context. Furthermore, We denote the $p$-norm of $u\in L^{p}$ by
$\|u\|_{p}:=\left(\int\left|u(z,\omega)\right|^{p}\mu(\mathop{}\!\mathrm{d}z)P(\mathop{}\!\mathrm{d}\omega)\right)^{1/p}=\left(\mathbb{E}\left|u(z,\omega)\right|^{p}\right)^{1/p},$ (2.3)
where $\mu$ denotes the probability measure for randomisation. We write $\|u\|:=\|u\|_{2}$ for brevity if $u\in L^{2}$. Note that $p$-norm characterises the $p$-moment of the random variable $u$.
We model each unit’s potential outcome function, via the structural mapping (2.1), as an element of a model space $\mathcal{M}_{i}$, which is a subspace of $L^{2}$. Formally, we write
Assumption 3 (Model Space).
$\tilde{y}_{i}(z,\omega)\in\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega).$ (2.4)
The framework of HWS can be viewed as a special case of our model, corresponding to the degenerate setting where $\Omega=\{\omega_{0}\}$ for some $\omega_{0}$. In this case, the latent variable $\omega_{0}$ is effectively fixed, and the realised data serve as the entire outcome-generating mechanism. Thus, the potential outcome function (2.1) becomes a deterministic function of $z$, recovering the setting considered in their work.
We equip $\mathcal{M}_{i}$ with the inner product
$\langle u,v\rangle:=\mathbb{E}_{\omega}\Bigl{[}\mathbb{E}_{z}\bigl{[}u(z,\omega)\,v(z,\omega)\bigr{]}\Bigr{]}=\mathbb{E}\bigl{[}u(z,\omega)\,v(z,\omega)\bigr{]},$ (2.5)
where the second equality follows from the law of iterated expectations. The associated norm is given by $\|u\|=\sqrt{\langle u,u\rangle}$, which coincides with the standard $L^{2}$-norm.
In the degenerate case where $\Omega=\{\omega_{0}\}$, the inner product reduces to
$\langle u,v\rangle^{0}=\mathbb{E}_{z}\bigl{[}u(z,\omega_{0})\,v(z,\omega_{0})\bigr{]},$ (2.6)
which reduces to the inner product used in HWS. This structure enables us to apply the Riesz representation theorem to construct estimators for treatment effects, even in the presence of interference.
2.3 Treatment Effect Functionals and Riesz Representation
While our model space $\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega)$ may not be complete under the inner product defined in (2.5), this poses no issue in what follows. Any inner product space admits a Hilbert space completion, and the Riesz representation theorem applies in that completed space. Hence, throughout this paper, whenever necessary, we implicitly work with the closure of $\mathcal{M}_{i}$ in $L^{2}(\mathcal{Z}\times\Omega)$.
Let $\mathcal{M}_{i}^{\prime}$ denote the dual space of $\mathcal{M}_{i}$.
Assumption 4 (Dual Representability of the Treatment Effect).
The treatment effect $\theta_{i}$ for each unit $i$ can be represented as a continuous linear functional on the model space of potential outcomes, viz., $\theta_{i}:\mathcal{M}_{i}\to\mathbb{R}$ with $\theta_{i}\in\mathcal{M}_{i}^{\prime}$.
Note that functions in an $L^{2}$ space are identified through equivalence classes, meaning two functions differing only on a set of measure zero are considered identical. Consequently, a well-defined linear functional must assign the same value to all functions within the same equivalence class.
A potential pitfall arises when attempting to use evaluation functionals, such as $\theta(u)=u(x_{0})$, that assign values based on a single point. Such evaluation maps generally are not well-defined in an $L^{2}$ context because they may yield different values for functions identical almost everywhere, violating linearity and the identification of elements in the same equivalence class.
In our analysis, we explicitly avoid these issues by requiring that the treatment effect functional $\theta_{i}$ be defined consistently on equivalence classes. More specifically, if two functions differ by a function of zero $L^{2}$-norm (i.e., distance zero), the functional $\theta_{i}$ shall assign them the same value.
From Assumption 4, we also require that the linear functional $\theta_{i}$ on a normed space is continuous or equivalently bounded. A heuristic justification is that requiring the treatment effect functional to be continuous ensures that small perturbations in the potential outcome function (as measured by the $L^{2}$ norm) result in small changes in the treatment effect. This continuity is essential for invoking the Riesz representation theorem, which in turn guarantees that $\theta$ can be uniquely represented by an element of the Hilbert space.
Researchers need not be concerned about such technicalities in practice, as they are always free to define or adjust their treatment effect functional to comply with the requirements in Assumptions 4. For instance, one might consider the following examples of continuous linear functionals.
Expected treatment effect for binary treatment
$\theta_{i}(u)\;=\;\mathbb{E}_{\omega}\bigl{[}u(z^{(1)},\omega)-u(z^{(0)},\omega)\bigr{]}$ (2.7)
representing the expected outcome difference when shifting treatment assignment from $z^{(0)}$ to $z^{(1)}$.
Expected treatment effect for treatment intervals or regions
$\theta_{i}(u)\;=\;\mathbb{E}_{\omega}\Bigl{[}\mathbb{E}_{z\in A}\bigl{[}u(z,\omega)\bigr{]}\Bigr{]}-\mathbb{E}_{\omega}\Bigl{[}\mathbb{E}_{z\in B}\bigl{[}u(z,\omega)\bigr{]}\Bigr{]},$ (2.8)
where $A,B\subset\mathcal{Z}$ and $A\cap B=\emptyset$, representing the expected outcome difference between treatment groups $A$ and $B$.
Expected partial derivative
$\theta_{i}(u)=\mathbb{E}_{\omega}\left[\frac{\partial u}{\partial z}(z_{0},\omega)\right],$ (2.9)
at some point $z_{0}\in\mathcal{Z}$, representing the expected marginal effect at $z_{0}$.
The next two propositions show that these functionals, under certain conditions, can be linear and continuous.
Proposition 1.
Suppose that $\mu(A)>0$ and $\mu(B)>0$. The functional (2.8) is linear and continuous on $\mathcal{M}_{i}$.
Note that Example (2.7) is a special case of Example (2.8), obtained by setting $A=\{z^{(1)}\}$ and $B=\{z^{(0)}\}$. The condition $\mu(A)>0$ and $\mu(B)>0$ is commonly referred to as the positivity or overlap assumption in the causal inference literature. It ensures that both treatment groups are represented in the randomisation distribution, which guarantees that the corresponding estimand is well-defined and estimable. In our case, the positivity assumption plays a crucial role in verifying that the functional is both linear and continuous – properties that are essential for the treatment effet to be well-defined and for the Riesz representation to hold.
It is more subtle to establish the linearity and continuity of the functional (2.9), as it requires additional structure from the model space.
Proposition 2.
[…]
This result implies that, for the functional to satisfy the necessary continuity condition, the model space must be further restricted to $\mathcal{M}_{i}\subset L^{2}(\Omega;H^{1}(\mathcal{Z}))\subset L^{2}$. Within this space, weak derivatives are well-defined and continuous, thereby ensuring the applicability of the Riesz representation theorem.
Theorem 2.1 (Riesz Representation Theorem in $\mathcal{M}_{i}$).
Under Assumptions 1, 3, and 4, for every $\theta_{i}\in\mathcal{M}_{i}^{\prime}$, there exists a unique element $\psi_{i}\in\mathcal{M}_{i}$ such that
$\theta_{i}(u)=\langle u,\psi_{i}\rangle\quad\text{for all }u\in\mathcal{M}_{i},$ (2.10)
and
$\|\theta_{i}\|_{\mathcal{M}_{i}^{\prime}}=\|\psi_{i}\|,$ (2.11)
where the dual norm is defined by
$\left\lVert\theta_{i}\right\rVert_{\mathcal{M}_{i}^{\prime}}:=\sup_{\|u\|\leq 1}\left\lvert\theta_{i}(u)\right\rvert.$ (2.12)
The motivation for this theorem is partly rooted in the approach of HWS, who proposed a design-based framework applying the Riesz representation theorem to characterise treatment effects as continuous linear functionals. Building on this foundational idea, we develop a broader formulation that accommodates random potential outcomes. In particular, the unique element $\psi_{i}$ representing the functional $\theta_{i}$ via the inner product, hereafter referred to as the Riesz representer, plays a central role in our estimation procedure. This generalisation retains the structural clarity of the original framework while introducing additional flexibility through stochastic heterogeneity in outcomes.
By constructing the Riesz representer $\psi_{i}$ corresponding to the treatment effect functional $\theta_{i}$, we only consider those treatment effects that are representable in the dual space $\mathcal{M}_{i}^{\prime}$. Consequently, we estimate the treatment effect
$\theta_{i}(\tilde{y}_{i})=\langle\tilde{y}_{i},\psi_{i}\rangle=\mathbb{E}\Bigl{[}\tilde{y}_{i}(z,\omega)\,\psi_{i}(z,\omega)\Bigr{]},$ (2.13)
by using the corresponding Riesz estimator
$\hat{\theta}_{i}(z,\omega)=\tilde{y}_{i}(z,\omega)\,\psi_{i}(z,\omega),$ (2.14)
as, in practice, we observe only one realisation of the outcome $\tilde{y}_{i}(z,\omega)$ for certain $z$ and $\omega$. It follows immediately that the Riesz estimator in (2.14) for each unit is unbiased for its treatment effect given in (2.13).
Definition 1 (The Estimand).
The finite-sample estimand we consider in this paper is defined as a weighted average of individual treatment effects
$\tau_{n}=\sum_{i=1}^{n}\nu_{ni}\theta_{i}(\tilde{y}_{i}),$ (2.15)
where the weights $\nu_{ni}\in\mathbb{R}$ satisfy $\nu_{ni}\geq 0$.
This quantity corresponds to the average treatment effect under a general weighting scheme and fits naturally within the broader context of causal inference, where estimands are often defined as averages over unit-level causal effects under various designs and sampling schemes.
Our formulation generalises the estimand considered by HWS, which corresponds to the special case:
$\tau_{n}^{0}=\frac{1}{n}\sum_{i=1}^{n}\langle\tilde{y}_{i},\psi_{i}\rangle^{0}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{z}\Bigl{[}\tilde{y}_{i}(z,\omega_{0})\,\psi_{i}(z,\omega_{0})\Bigr{]},$ (2.16)
where $\Omega=\{\omega_{0}\}$ is a singleton probability space. In this special case, the treatment effect is deterministic and the average is equally weighted. It is worth noting that the estimand in HWS does not correspond to the average treatment effect in the usual sense. Since their framework omits randomness in the covariates, the resulting estimand reflects a fixed-realisation functional rather than an expectation over the full distribution of potential outcomes. In contrast, our formulation integrates over both observed and unobserved heterogeneity, thereby recovering an estimand that aligns with the average treatment effect as commonly defined in the broader causal inference literature.
A common choice for the weights is the uniform weighting $\nu_{ni}=1/n$, which yields the sample average treatment effect (SATE). More generally, this formulation accommodates subgroup-specific or covariate-adjusted averages, depending on the structure of the weights. We impose the following assumption to control the magnitude of the weights $\nu_{ni}$.
Assumption 5 (Uniform Weight Bound).
There exist a constant $\bar{\nu}>0$ such that
$n\nu_{ni}\leq\bar{\nu}\quad\text{for all }i\text{ and all }n\in\mathbb{N}.$ (2.17)
It guarantees that each weight is bounded by a constant multiple of $1/n$, uniformly in $i$, without requiring the individual sequences $\{n\nu_{ni}\}_{n\geq i}$ to converge. Consequently, no finite subset of indices can receive an asymptotically dominant share of the total weight. Note that Assumption 5 implies $\sup_{n\to\infty}\sum_{i=1}^{n}\nu_{ni}<\infty$ but not necessarily one, which is quite general. In the continuous analogue this corresponds to the integrability of the weight function.
Given the unit-level Riesz estimators (2.14) and the weights,
Definition 2 (The Estimator).
we define the corresponding aggregate Riesz estimator for the finite-sample estimand
$\hat{\tau}_{n}(z,\omega)=\sum_{i=1}^{n}\nu_{ni}\hat{\theta}_{i}(z,\omega).$ (2.18)
Thus we have the unbiasedness of the Riesz estimator in the following.
Theorem 2.2 (Unbiasedness of the Riesz Estimator).
Under Assumptions1, 3, and 4, the Riesz estimator defined in (2.14) is unbiased for the individual treatment effect, i.e., $\mathbb{E}\bigl{[}\hat{\theta}_{i}(z,\omega)\bigr{]}=\theta_{i}(\tilde{y}_{i})$. Consequently, the aggregate Riesz estimator in (2.18) is unbiased for the finite-sample estimand (2.15), i.e., $\mathbb{E}\Bigl{[}\hat{\tau}_{n}(z,\omega)\Bigr{]}=\tau_{n}$.
3 Large Sample Properties of the Aggregate Riesz Estimator
In this section, we proceed by clarifying the notion of local independence required in our framework. Informally, local dependence refers to the idea that subsets of random variables are conditionally independent of those outside their respective neighbourhoods. This structure is often represented using a dependency graph, in which vertices correspond to units and edges encode potential statistical dependence. For further information about the dependence structure of the variables, see for example Chen and Shao (2004).
Definition 3 (Dependency Neighbourhoods).
For each unit $i$, let $\{M_{ik}\}_{k\in I_{i}}$ be the collection of all subsets $M_{ik}\subset\{1,2,\dots,n\}$ such that the pair of random variables $(\tilde{y}_{i},\psi_{i})$ is independent of the set $\left\{(\tilde{y}_{j},\psi_{j}):j\notin M_{ik}\right\}$. We define the dependency neighbourhood of unit $i$ as the intersection over all such sets:
$N_{i}:=\bigcap_{k\in I_{i}}M_{ik}.$ (3.1)
This definition ensures that $N_{i}$ is the smallest subset of units for which independence holds. Formally, for all such $j\notin N_{i}$, and for all measurable sets $A_{i}\in\mathcal{F}_{\tilde{y}_{i}}$, $B_{i}\in\mathcal{F}_{\psi_{i}}$, $A_{j}\in\mathcal{F}_{\tilde{y}_{j}}$, and $B_{j}\in\mathcal{F}_{\psi_{j}}$, we have
$\mathbb{P}\left((A_{i}\cap B_{i})\cap(A_{j}\cap B_{j})\right)=\mathbb{P}(A_{i}\cap B_{i})\,\,\mathbb{P}(A_{j}\cap B_{j}).$
This definition corresponds to Definition 3.1 in Ross (2011) and is a weaker version of Definition 5.1 in HWS, as it only requires independence of the potential outcome $\tilde{y}_{i}$ and the Riesz representer $\psi_{i}$ from those outside $N_{i}$, rather than full functional independence across the spaces $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$.
In contrast to HWS, who assume independence of the full model spaces across units, our framework only requires independence between the observable random pairs $(\tilde{y}_{i},\psi_{i})$. This relaxation is justified by the introduction of a latent random variable $\omega$, under which potential outcomes are modelled as conditionally random functions. Importantly, our formulation permits cases where the model spaces coincide, i.e. $\mathcal{M}_{i}=\mathcal{M}_{j}$, so long as the associated outcome representer pairs remain independent. Consequently, we do not assume functional independence at the space level, but merely require that the realised outcomes and their associated Riesz representers be independent across sufficiently separated units. This distinction both broadens the scope of our analysis and clarifies the assumptions necessary for variance control and asymptotic normality in a design-based setting.
We denote by $\left\lvert N_{i}\right\rvert$ the number of units in the dependency neighbourhood of unit $i$, which we refer to as the neighbourhood size for unit $i$. Let $D_{n}=\max_{1\leq i\leq n}\left\lvert N_{i}\right\rvert$ and $d_{n}=n^{-1}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert$ denote, respectively, the maximum and average neighbourhood sizes across the sample. These quantities will play a central role in the asymptotic analysis to follow.
In practice, the specification of neighbourhoods $N_{i}$ depends on substantive knowledge of the experimental context. In network experiments, for example, $N_{i}$ might include unit $i$ and its observed social or physical connections. In spatial designs, it is common to define $N_{i}$ via fixed-radius rules or $k$-nearest-neighbour relationships based on geographic distance. When no such structure is explicitly available, one can conservatively define $N_{i}$ to include all units within the same covariate stratum or treatment block as unit $i$. These strategies allow researchers to encode plausible channels of dependence without imposing strong model-based assumptions.
The following two assumptions restrict the growth rate of the dependency neighbourhoods.
Assumption 6 (Maximum Neighbourhood Size).
The maximum dependency neighbourhood size $D_{n}$ satisfies $D_{n}=o(n^{d})$ for some $d>0$.
Assumption 7 (Average Neighbourhood Size).
The average dependency neighbourhood size $d_{n}$ satisfies
  1. (a)
    $d_{n}=o(n)$;
  2. (b)
    $\sup_{n\in\mathbb{N}}d_{n}<\infty$.
In addition, for a sequence of functions $u_{1},\dots,u_{n}\in L^{p}$, where $p>1$, we define the max-$p$ norm as
$\left\lVert u\right\rVert_{\max,p}^{n}:=\max_{1\leq i\leq n}\left\lVert u_{i}\right\rVert_{p}.$ (3.2)
The sequence $\left\lVert u\right\rVert_{\max,p}^{n}$ is clearly non-decreasing in $n$, and, as such, is bounded if and only if it converges as $n\to\infty$.
We now state a proposition that provides an upper bound on the asymptotic variance, denoted by $\sigma_{n}^{2}:=\mathbb{V}[\hat{\tau}_{n}(z,\omega)]$, of the aggregate Riesz estimator $\hat{\tau}_{n}(z,\omega)$ of the estimand $\tau_{n}$. This result follows from the setup and notation introduced above, and it serves to justify or at least suggest the consistency of the estimator under appropriate conditions.
Proposition 3 (Variance Upper Bound).
Under Assumptions 1 and 3-5, the following inequality holds
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\,\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2},$ (3.3)
for any $p,q\in[1,\infty]$ satisfying $21/p+1/q=1/$ 2. The inequality holds trivially if either $\|\tilde{y}_{i}\|_{p}$ or $\|\psi_{i}\|_{q}$ is infinite.
Furthermore, we define $S_{2}:=\left\{(p,q)\in[1,\infty]^{2}:\,\frac{1}{p}+\frac{1}{q}=\frac{1}{2}\right\}$, and then we have
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}d_{n}}{n}\left(\inf_{p,q\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$ (3.4)
The inequality hold trivially if $\inf_{p,q\in S}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}$ is infinite.
This proposition differs from Proposition 5.1 in HWS, as it allows for more general weighting schemes for individual treatment effects. Nonetheless, it conveys the same essential message. Specifically, it reveals how the upper bound on the variance depends on the cardinalities of the dependency neighbourhoods, the weights $\nu_{ni}$, and the moments of $\tilde{y}_{i}$ and $\psi_{i}$.
In particular, the result implies that the variance may decay as the sample size increases, provided that the norms $\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}$ and $\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}$ remain bounded if there exists such a pair $p,q\in[1,\infty]$ satisfying $21/p+1/q=1/$ 2, and that the neighbourhood sizes $\left\lvert N_{i}\right\rvert$ are uniformly small. The bound highlights the importance of controlling both the complexity of dependence and the magnitude of unit-level variation in order to obtain valid inference.
Assumption 8 (Boundedness of Norms).
For some $r\geq 1$, there exist real numbers $p,q\in[r,\infty]$ satisfying $1/p+1/q=1/r$ such that
$\sup_{i\in\mathbb{N}}\left\lVert\tilde{y}_{i}\right\rVert_{p}<\infty\quad\text{and}\quad\sup_{i\in\mathbb{N}}\left\lVert\psi_{i}\right\rVert_{q}<\infty.$ (3.5)
This assumption is formulated to allow flexibility in the choice of $r$. Intuitively, the condition imposes uniform boundedness on the $p$th moments of $\tilde{y}_{i}$ and the $q$th moments of $\psi_{i}$, thereby excluding heavy-tailed behaviour.
Note that the max-$p$ norm satisfies $\left\lVert u\right\rVert_{max,p}^{n}\leq\sup_{i\in\mathbb{N}}\left\lVert u_{i}\right\rVert_{p}$. Therefore, Assumption 8 is equivalent to requiring that the four quantities $\|\tilde{y}\|_{\max,p_{1}}^{n}$, $\|\psi\|_{\max,q_{1}}^{n}$, $\|\tilde{y}\|_{\max,p_{2}}^{n}$, and $\|\psi\|_{\max,q_{2}}^{n}$ remain bounded and converge as $n\to\infty$.
Then we have the following theorem about the asymptotic consistency of the aggregate Riesz estimator.
Theorem 3.1 (Consistency).
Under Assumptions 1, 3-5, 7$(a)$, and 8(r=2), the aggregate Riesz estimator in (2.18) is consistent in mean square. Moreover, if 7$(b)$ holds, then the the aggregate Riesz estimator satisfies $O_{p}(n^{-1/2})$.
We should interpret $O_{p}(n^{-1/2})$ with care. This notation means that the sequence $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ is bounded in probability, but it does not imply convergence in distribution to a single non-degenerate law. In fact, the sequence may vary between different limiting distributions with bounded variances.
So far, we have shown that, under certain conditions, the aggregate Riesz estimator is consistent. However, these conditions alone are not sufficient for establishing asymptotic normality; an observation that aligns with the fact that consistency, in general, does not imply asymptotic normality.
The following non-degeneracy assumption imposes a lower bound on the variance of the aggregate Riesz estimator and plays a critical role in ensuring asymptotic normality. When combined with the previous assumptions, it provides sufficient conditions for the aggregate Riesz estimator to satisfy a central limit theorem.
Assumption 9 (Variance Lower Bound).
$\inf_{n\in\mathbb{N}}\sqrt{n}\,\sigma_{n}\geq\sigma_{0},\quad\text{for some }\sigma_{0}>0.$ (3.6)
This condition corresponds to Assumption 2 in HWS. It allows the asymptotic variance $\sigma_{n}^{2}$ to decay with $n$, but not too rapidly. Intuitively, it ensures that the root-$n$ scaled estimator $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ retains a non-degenerate variance in the limit.
In the proof of the theorem, we aim to bound the Wasserstein distance between a sequence of measures and the standard normal. As shown in (D.5), the upper bound on this distance involves inverse powers of the variance $\sigma_{n}^{2}$. Without imposing a lower bound on $\sqrt{n}\sigma_{n}$, the upper bound on the Wasserstein distance derived in (D.5) may fail to converge to zero. While this does not necessarily imply that convergence in distribution to the normal law fails, it means that we can no longer guarantee asymptotic normality using this bounding argument. The assumption therefore provides a sufficient condition to ensure that the Wasserstein bound vanishes, thereby supporting convergence in distribution to a non-degenerate normal law.
We require further the following assumption to establish the asymptotic normality result, which concerns the existence of finite fourth moments.
Assumption 10 (Finite Fourth Moment).
The individual treatment effect estimator satisfies:
  1. (a)
    $\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)^{4}\right]<\infty$;
  2. (b)
    $\sup_{i\in\mathbb{N}}\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)^{4}\right]<\infty$.
Note that condition $(b)$ imposes a stronger requirement than $(a)$, as it demands uniform boundedness of the fourth moments across all units.
Theorem 3.2 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, and 10$(a)$, the aggregate Riesz estimator defined in (2.18) satisfies the following asymptotic normality
[…]
Both Assumptions 8(r=3) and 8(r=4) are required here, as boundedness of the max-$p$ norms at $r=4$ does not necessarily imply boundedness at $r=3$ across the joint probability space $(\mathcal{Z},\Omega)$ when it is unbounded.
Theorem 3.2 provides the foundation for statistical inference based on the aggregate Riesz estimator. In particular, it enables asymptotic testing of the null hypothesis $H_{0}:\tau_{n}=0$ against the alternative $H_{1}:\tau_{n}\neq 0$, using the standardised statistic $T_{n}=\hat{\tau}_{n}(z,\omega)/\sigma_{n}$, provided that both $\hat{\tau}_{n}(z,\omega)$ and $\sigma_{n}$ can be consistently estimated from a single realisation of the assignment vector $z$ and the latent variable $\omega$.
It is important to emphasise that the estimand $\tau_{n}$ represents a population-level quantity: it is defined as a weighted average of individual treatment effects, where each individual effect is itself an expectation taken over the joint randomness in both z and $\omega$. Thus, $\tau_{n}$ depends on the full data-generating distribution, not just on a particular observed sample.
The remarkable implication of Theorem 3.2 is that, under suitable regularity conditions, we can construct valid asymptotic tests and confidence intervals for this population estimand based solely on a single realisation of $z$ and $\omega$. This result ensures that standard Wald-type inference procedures remain valid, despite the underlying potential outcomes and treatment effects being defined via expectations over unobserved randomness.
The required rate condition on the maximum dependency neighbourhood size, $D_{n}=o(n^{1/4})$, mirrors that of HWS. However, the proof here proceeds by bounding Wasserstein distance under mixed expectations over both the treatment assignments $z$ and latent heterogeneity $\omega$. This highlights a key observation: the dependence-degree requirement does not tighten despite the added randomness in potential outcomes.
This result is particularly relevant in empirical settings where potential outcomes are inherently stochastic. For instance, in field experiments involving sensor-based measurements, observed outcomes may be contaminated by random noise due to hardware limitations or environmental fluctuations. Similarly, in panel experiments with repeated units, treatment effects may be estimated via machine learning algorithms that introduce randomness in the form of first-stage fitted values. In both cases, the stochastic component of the outcome arises naturally and cannot be ignored. The asymptotic normality result ensures that inference remains valid under such randomness.
4 Variance Estimation under Local Dependence
The test statistic supported by Theorem 3.2 requires an estimate of the variance of the aggregate Riesz estimator $\hat{\tau}_{n}$, which is generally unknown in finite samples. In the presence of local dependence, such as that arising from network interference, exposure mappings, or latent factors, standard i.i.d. variance estimators are invalid, and specialised methods must be employed.
Several approaches have been proposed. Some yield consistent estimators, but only under strong structural assumptions. For example, Aronow and Samii (2017) rely on known exposure mappings and joint inclusion probabilities; Yu et al. (2022) assume bounded-degree network structures; and Liu and Hudgens (2014) consider two-stage designs with bounded group sizes. Other methods construct variance bounds rather than consistent estimators: Chin (2019) derive conservative expressions using Stein’s method, while Chen and Shao (2004) propose block bootstrap procedures that are valid under fixed neighbourhood size and finite moment assumptions.
Of particular relevance to our setting, HWS construct an unbiased estimator for an upper bound on the variance using a tensor-product Riesz representation. More recently, Harshaw et al. (2024) develop a constrained optimisation framework to tighten such conservative bounds across a wide range of experimental designs. However, neither method formally proves that their bound converges to the true variance, except in special cases. Without sharpness, these bounds may overestimate the variance and lead to overly conservative inference.
In this paper, we take a different approach. Rather than estimating a variance upper bound, we propose a direct estimator for the true asymptotic variance of the aggregate Riesz estimator under local dependence. This estimator is shown to be consistent under conditions that essentially match those required for asymptotic normality. The key requirement is knowledge of the dependency neighbourhood structure: not only the maximal neighbourhood size $D_{n}$, but also which specific units are likely to be dependent. This may seem restrictive, but it is consistent with the broader literature, where many authors (including HWS) note that the variance is generally non-identifiable from a single realisation without further assumptions.
Our recommendation is practical: given estimates $\zeta_{i}=\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$, we define the empirical variance estimator by retaining only the terms that correspond to known (or assumed) dependent pairs. The rest are set to zero. This enables consistent estimation of the variance and allows us to construct feasible and asymptotically valid test statistics.
The challenge of variance estimation under local dependence can be seen as a non-identifiability problem. Let $\vec{\zeta}_{n}=(\zeta_{1},\dots,\zeta_{n})^{\prime}$ denote a vector of centred random variables with covariance matrix $\Sigma_{n}=\mathbb{E}[\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}]$. From a single realisation, we observe only the outer product $\hat{\Sigma}_{n}=\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}$, and cannot estimate $\Sigma_{n}$ consistently without further structure. Even under independence, each $\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}$ typically has variance of order one, so
$\mathbb{E}\bigl{[}\|\hat{\Sigma}_{n}-\Sigma_{n}\|_{F}^{2}\bigr{]}\;\geq\;cn^{2}$ (4.1)
for some constant $c>0$, where $\left\lVert\cdot\right\rVert_{F}$ stands for the Frobenius norm for a matrix, indicating that the Frobenius error diverges with $n$.
Consistency can be recovered by imposing stronger assumptions, for example: observing multiple independent copies of $\vec{\zeta}_{n}$, enabling standard sample covariance estimators; time ordering and weak stationarity allowing the use of banded estimators or kernel-smoothed autocovariances; structural models such as factor structures (low-rank plus sparse decomposition), enabling regularised estimation (e.g., POET); vanishing signal assuming that $\mathbb{V}[\zeta_{n}]\to 0$ as $n\to\infty$, so that the norm $\|\Sigma_{n}\|_{F}$ itself converges to zero; among others.
In the absence of such assumptions, the full matrix $\Sigma_{n}$ is not consistently estimable. Our proposed approach sidesteps this by targeting only those components of the variance that correspond to known dependent pairs.
Suppose the dependency structure is given by
$\mathcal{E}_{n}=\{(i,j)\in\{1,\dots,n\}^{2}:\zeta_{i}\text{ and }\zeta_{j}\text{ are dependent}\}.$ (4.2)
We define the plug-in variance estimator
$\hat{\sigma}_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j},$ (4.3)
and the population variance
$\sigma_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\mathbb{E}[\zeta_{i}\zeta_{j}].$ (4.4)
The following theorem shows that, under mild regularity conditions, $\hat{\sigma}_{n}^{2}$ consistently estimates $\sigma_{n}^{2}$.
Theorem 4.1 (Consistency of the Variance Estimator).
Under Assumptions 1, 3-5, 6(d=1/3), and 10$(b)$, the variance estimator $n\hat{\sigma}_{n}^{2}$ converges to $n\sigma_{n}^{2}$ in mean square.
Corollary 1 (Asymptotic Normality).
Under Assumptions 1, 3-5, 6(d=1/4), 8(r=3), 8(r=4), 9, and 10$(b)$, the aggregate Riesz estimator defined in (2.18) satisfies the following asymptotic normality
[…]
Note that Assumption 6(d=1/4) is stronger than Assumption 6(d=1/3) and therefore suffices to imply both consistency and asymptotic normality.
Corollary 1 establishes that the aggregate Riesz estimator is asymptotically normal under local dependence, provided that the dependency neighbourhoods are sufficiently sparse and the variance estimator $\hat{\sigma}_{n}^{2}$ is properly constructed.
In practice, estimating the variance $\sigma_{n}^{2}$ requires specifying which unit pairs are believed to be dependent. Suppose that for each sample, the dependency structure is approximated by the index set $\mathcal{E}_{n}\subset\{1,\dots,n\}^{2}$ defined in (4.2). We construct a sparse matrix $\hat{\Sigma}_{n}^{d}\in\mathbb{R}^{n\times n}$ by retaining only the terms associated with dependent pairs:
$(\hat{\Sigma}_{n}^{d})_{ij}\;=\;\begin{cases}\zeta_{i}\zeta_{j}&\text{if }(i,j)\in\mathcal{E}_{n},\\ 0&\text{otherwise}.\end{cases}$ (4.6)
The corresponding plug-in variance estimator is then given by
$\hat{\sigma}_{dn}^{2}\;=\;\nu_{n}^{\prime}\hat{\Sigma}_{n}^{d}\nu_{n},$ (4.7)
where $\nu_{n}=(\nu_{n1},\dots,\nu_{nn})^{\top}$ is the vector of aggregation weights. This estimates the true sampling variance
$\sigma_{n}^{2}=\nu_{n}^{\prime}\Sigma_{n}\nu_{n},$ (4.8)
as defined in (4.4).
The consistency result in Theorem 4.1 ensures that the plug-in estimator $n\hat{\sigma}_{dn}^{2}$ converges to $n\sigma_{n}^{2}$, provided that the approximated variance $n\hat{\sigma}_{dn}^{2}$ is sufficiently close to the ideal estimator $n\hat{\sigma}_{n}^{2}$, which in turn requires that the underlying dependency structure is at least approximately known. When such an approximation is feasible, the resulting test statistic based on (4.5) is valid and interpretable.
If the units are believed to be independent, then each $\zeta_{i}$ is uncorrelated with every other $\zeta_{j}$, and the dependency set $\mathcal{E}_{n}$ consists only of diagonal pairs $(i,i)$. In this case, the plug-in matrix reduces to a diagonal matrix and the variance estimator becomes
$\hat{\sigma}_{n}^{2}=\sum_{i=1}^{n}\nu_{ni}^{2}\zeta_{i}^{2},\quad\sigma_{n}^{2}=\sum_{i=1}^{n}\nu_{ni}^{2}\mathbb{E}[\zeta_{i}^{2}],$
which corresponds to the classical form used in standard inference under independence. Corollary 1 thus recovers the familiar asymptotic normality result as a special case of the more general locally dependent setting treated here.
Taken together, Theorems 3.14.1 and Corollary 1 convey a central message: although the estimand $\tau_{n}$ is defined as an average of individual treatment effects evaluated over the joint probability space $(\mathcal{Z},\Omega)$, the entire inference procedure, from estimation to variance approximation, can be carried out using just a single realisation of $z\in\mathcal{Z}$ and $\omega\in\Omega$. This is made possible by the structure of the Riesz representation framework and the use of carefully constructed estimators that remain valid under local dependence. As the sample size $n$ grows, one draw from the treatment and outcome distribution contains sufficient information for reliable estimation and inference, provided that the underlying conditions hold. In this way, the results provide a rigorous justification for making population-level claims based on one observed experiment.
5 Riesz Representer Estimation in Practice
The theoretical results in earlier sections show that the aggregate Riesz estimator can be constructed using only a single realisation of the underlying data-generating process. Specifically, inference is based on one realisation $\omega_{0}\in\Omega$ of the latent variable $\omega$, and one draw of the treatment vector $z$.
In our framework, expectations are taken over the randomisation distribution of $z$, conditional on $\omega_{0}$, and, crucially, this suffices to identify the estimand, which is defined as an expectation over both $z$ and $\omega$.
This structure has concrete implications for how we estimate the individual Riesz representers $\psi_{i}(z,\omega)$. By Theorem 2.1, for each unit $i$, there exists a unique representer satisfying (2.10) with (2.6). Hence, the representer $\psi_{i}(z,\omega_{0})$ is identified conditionally and can be estimated directly from the realised function $\tilde{y}_{i}(z,\omega_{0})$.
Note that this formulation generalises the setting considered by HWS, which corresponds to the special case where $\Omega=\{\omega_{0}\}$ is a singleton—i.e., the potential outcomes are deterministic. While HWS recover $\psi_{i}$ using similar inner-product logic, our framework accommodates stochastic potential outcomes and makes clear that conditional estimation is still sufficient for recovering population-level effects.
5.1 Computation in Finite-Dimensional Model Spaces
In finite-dimensional model spaces $\mathcal{M}_{i}$, we estimate $\psi_{i}(z,\omega_{0})$ via basis expansion and moment matching. Let $\{g_{i,1},\dots,g_{i,m}\}$ be a basis for $\mathcal{M}_{i}$ and $\mathrm{dim}(\mathcal{M}_{i})=m$. We solve for coefficients $\bm{\beta}_{i}\in\mathbb{R}^{m}$ such that
$\hat{\psi}_{i}(z,\omega_{0})=\sum_{k=1}^{m}\beta_{i,k}\,g_{i,k}(z,\omega_{0})$ (5.1)
satisfies the identity $\theta_{i}(\tilde{y}_{i})=\langle\tilde{y}_{i},\hat{\psi}_{i}\rangle^{0}$. Algorithm 1 summarises this procedure.
Algorithm 1 follows the same procedure as described by HWS (preprint version), in which Riesz representers are computed via basis expansion and moment matching under the randomisation distribution. While we adopt the same general approach, we provide here a few concrete examples of basis choices to clarify practical implementation.
When the treatment space $\mathcal{Z}$ is finite with $K$ distinct levels, a simple and interpretable basis consists of treatment indicators. In this case, one may define each basis function $g_{i,k}(z,\omega)$ to equal 1 if $z$ corresponds to the $k$-th treatment condition and 0 otherwise. This yields a saturated model with a separate coefficient for each treatment arm and does not rely on any structural assumptions about the response surface.
When $z$ is continuous or high-dimensional, one can instead use low-order basis functions such as polynomials or splines in $z$, possibly interacted with pre-specified functions of $\omega$. For instance, if $z\in\mathbb{R}$, a natural choice is the span of $\{1,z,z^{2},\omega,z\cdot\omega\}$, which captures basic nonlinearities and interactions between treatment and latent outcome variation. These basis sets offer a balance between expressiveness and computational tractability and can be tailored to the structure of the experiment.
This procedure is fully offline in the sense that it depends only on the design distribution for $z$, the basis functions, and the known functional $\theta_{i}$. It does not require access to observed outcomes $Y_{i}=\tilde{y}_{i}(z_{0},\omega_{0})$.
This construction relies on Assumption 2, which ensures that the treatment assignment $z$ is independent of the latent variable $\omega$. In particular, this allows expectations over $z$ conditional on a fixed realisation $\omega_{0}$ to be taken with respect to the known design distribution $\mu$. Without this assumption, the conditional distribution of $z\mid\omega_{0}$ could differ from $\mu$, and the estimated representer would no longer correspond to the true Riesz representation defined under the design. It is at this stage, during the offline construction of $\hat{\psi}_{i}(z,\omega_{0})$, that Assumption 2 is invoked. All subsequent inference proceeds without requiring independence between treatment assignment and outcome-level randomness.
In symmetric settings with common $\mathcal{M}_{i}$ and $\theta_{i}$, the same representer $\psi$ can be reused across units.
5.2 Approximation in Infinite-Dimensional Model Spaces
In many practical applications, the model space $\mathcal{M}_{i}$ is infinite-dimensional. This arises naturally when either the treatment assignment $z$ or the latent variable $\omega$ lies in a continuous domain, or when potential outcomes depend on complex functional relationships. To compute the Riesz representer in such settings, we project the problem onto a sequence of growing finite-dimensional subspaces.
Our setting explicitly accounts for a random latent variable $\omega$, and the Riesz representation is defined with respect to an inner product $\langle\cdot,\cdot\rangle^{0}$ taken over the treatment assignment $z$, conditional on the realised value $\omega_{0}\in\Omega$. Although our estimation routine shares a computational structure with existing methods, constructing a Gram matrix and solving a linear system, our framework generalises the representer computation to a richer class of problems involving stochastic potential outcomes.
Assumption 11 (Separable Model Space).
The model space $\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega)$, equipped with the inner product $\langle\cdot,\cdot\rangle^{0}$ defined in (2.6) for some $\omega_{0}\in\Omega$, is a separable Hilbert space.
By Assumption 11, the space $\mathcal{M}_{i}$ admits a countable orthonormal basis $\{e_{k}\}_{k\geq 1}$. For each dimension $m\in\mathbb{N}$, define the subspace
$\mathcal{M}_{i}^{(m)}=\operatorname{span}\{e_{1},\dots,e_{m}\},$ (5.2)
and let $P_{m}:\mathcal{M}_{i}\to\mathcal{M}_{i}^{(m)}$ denote the orthogonal projection. Then the projected representer
$\psi_{i}^{(m)}:=P_{m}\psi_{i}=\sum_{k=1}^{m}\langle\psi_{i},e_{k}\rangle^{0}\,e_{k}$ (5.3)
converges to $\psi_{i}$ in norm as $m\to\infty$, by the completeness of the Hilbert space $\mathcal{M}_{i}$, since $\{e_{k}\}_{k\geq 1}$ forms a total orthonormal system and $P_{m}$ is the orthogonal projection onto the span of the first $m$ basis elements.
The estimation of $\psi_{i}^{(m)}$ proceeds by solving the finite-dimensional moment equation system that results from projecting the Riesz representation identity onto the span of the first $m$ basis functions. This yields a linear system involving a Gram matrix and a target vector, as shown in Algorithm 2.
This procedure is still fully offline in the sense that both $G_{i}^{(m)}$ and $T_{i}^{(m)}$ can be computed offline from the design distribution and the known linear functional $\theta_{i}$. It does not require access to observed outcomes $Y_{i}=\tilde{y}_{i}(z_{0},\omega_{0})$. In symmetric settings with common $\mathcal{M}_{i}$ and $\theta_{i}$, the same representer $\psi$ can be reused across units.
The procedure for estimating $\psi_{i}(z,\omega_{0})$ described above relies on projecting onto a finite-dimensional subspace spanned by a pre-specified basis and solving empirical moment equations derived from the known randomisation design and linear functional $\theta_{i}$. This approach follows the classical sieve estimation framework and is theoretically well understood; see, e.g., Chen (2007); Chen and Pouzo (2012); van der Vaart (2000).
While consistency of $\hat{\psi}_{i}^{(m)}\to\psi_{i}$ under appropriate regularity conditions can be rigorously established using a standard bias-variance decomposition, this is not the focus of the present paper. Our main contribution lies in the theory developed in earlier sections, namely, the design-based asymptotic normality of the aggregate estimator and the construction of consistent variance estimators under local dependence.
For completeness, we note that under standard conditions, such as an orthonormal basis, eigenvalues of the Gram matrix bounded away from zero, and sufficient smoothness of the representer, the estimation error satisfies
$\|\widehat{\psi}_{i}^{(m)}-\psi_{i}\|=O_{p}\left(m^{-s}+\sqrt{m/n}\right)$ (5.4)
for some $s>0$. This rate is minimised by setting the truncation level $m$ proportional to $n^{1/(2s+1)}$, which balances the bias and variance contributions. We omit the full proof, as it follows standard arguments from the sieve estimation literature.
The orthonormal system $\{e_{k}\}$ can either be specified a priori, using classical bases such as Fourier series, wavelets, or splines, or derived from the data through methods like principal components or kernel eigenfunctions. Additional practical examples of basis construction in deterministic settings are discussed in the preprint by HWS, to which we refer the interested reader for further details.
6 Simulation Study
We carry out a simulation study to evaluate the finite-sample performance of the Riesz estimator and the plug-in variance estimator. The simulation is designed to reflect a typical experimental setting with local dependence induced by latent block-level effects. All potential outcomes, covariates, and errors are modelled as functions of a common latent variable, consistent with our framework.
For each unit $i\in\{1,\dots,n\}$, we simulate treatment assignment $z_{i}\sim\mathrm{Bernoulli}(0.5)$ independently. Covariates $x_{i}(\omega)$ and outcome-level parameters $(\alpha_{i},\beta_{i},\delta_{i})$ are drawn i.i.d. from $\mathcal{N}(0,1)$. The potential outcome is generated via
$\tilde{y}_{i}(z,\omega)=\alpha_{i}+\beta_{i}z+\delta_{i}zx_{i}+\varepsilon_{i}(\omega),$ (6.1)
where $\varepsilon_{i}(\omega)$ is a locally dependent noise term described below. The observed outcome is $Y_{i}=\tilde{y}_{i}(z_{i},\omega)$.
Local dependence is introduced by partitioning the sample into blocks, where units in the same block share a common latent shock. Specifically, the number of blocks is $B_{n}=\lfloor n^{1-d}\rfloor$ for a fixed parameter $d\in[0,0.3]$. The error for unit $i$ is
$\varepsilon_{i}(\omega)=\gamma_{i}\eta_{b(i)}+\nu_{i},$
where $b(i)$ is the block to which unit $i$ is assigned, $\eta_{b}\sim\mathcal{N}(0,\sigma^{2})$ is the shared block-level shock (with $\sigma^{2}=1$), $\nu_{i}\sim\mathcal{N}(0,1)$ is an idiosyncratic noise term, and $\gamma_{i}\in\{-1,1\}$ is a random sign (independent Rademacher) to allow both positive and negative correlations between units in the same block. This construction yields a dependency graph $\mathcal{E}_{n}$ with within-block dependence and across-block independence.
We estimate the average treatment effect using the aggregate Riesz estimator. The representer $\psi_{i}$ is computed as
$\psi_{i}(z,\omega_{0})=\frac{z}{\mu_{1}}-\frac{1-z}{\mu_{0}},\quad\text{with }\mu_{1}=\mathbb{E}[z_{i}]=0.5,$
which corresponds to the efficient estimator under the known design. The estimator is
$\hat{\tau}_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}\psi_{i}(z_{i},\omega_{0}),$
and the plug-in variance estimator follows (4.3)
$\hat{\sigma}_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j},\quad\text{where }\zeta_{i}=Y_{i}\psi_{i}(z_{i})-\theta_{i}(\tilde{y}_{i})\text{ and }\nu_{ni}=1/n.$
That is, we sum pairwise products of residuals over all unit pairs within the same block.
We consider sample sizes $n\in\{100,200,500,1000\}$ and dependency growth parameters $d\in\{0.0,0.1,0.2,0.25,0.3\}$. For each $(n,d)$ pair, we run $2000$ independent replications. For each replication, we compute the estimator $\hat{\tau}_{n}$, the plug-in standard error $\hat{\sigma}_{n}$, and test statistics for Wald-type inference. We record the following metrics:
  • Empirical bias: $\mathbb{E}[\hat{\tau}_{n}-\tau_{n}]$;
  • Root mean squared error (RMSE): $\sqrt{\mathbb{E}[(\hat{\tau}_{n}-\tau_{n})^{2}]}$;
  • Coverage of 95% confidence intervals: $\mathbb{P}(|\hat{\tau}_{n}-\tau_{n}|\leq 1.96\hat{\sigma}_{n})$;
  • Rejection rates at 1%, 5%, and 10% significance levels.
The complete codebase, including replication scripts and plotting routines, is publicly available at github.com/yukai-yang/RieszRE_Experiments under the MIT license.
Figure 1 depicts the empirical bias and root mean squared error (RMSE) of the Riesz estimator as functions of the dependency growth rate $d$, stratified by sample size. Across all settings, the estimator exhibits negligible bias: empirical values remain close to zero, typically below 0.01 in absolute magnitude. The RMSE decreases steadily as the sample size increases, from approximately 0.37 at $n=100$ to around 0.10 at $n=1000$, indicating improved precision in larger samples. These results offer strong empirical support for the consistency of $\hat{\tau}_{n}$ under local dependence, as established in Theorem 3.1.
Figure 2 shows the empirical coverage rates of nominal 95% confidence intervals across varying levels of dependency growth $d$ and sample sizes $n$. Coverage rates are uniformly at or above the nominal level in all scenarios. For $n=100$, coverage ranges from 94.9% to 95.6%, while for $n=500$ and $n=1000$, it stabilises in the range of 95.5% to 96.8%. These results indicate that the plug-in variance estimator is slightly conservative but well-calibrated, yielding confidence intervals with reliable coverage even in the presence of moderate local dependence.
Table 2 reports the empirical rejection rates of two-sided Wald tests at the 1%, 5%, and 10% significance levels across all combinations of $n$ and $d$. At the 5% level, rejection rates generally lie within the 4%–5% range. For instance, when $n=1000$, the empirical size ranges from 3.4% to 4.5%, depending on the degree of local dependence. Similar patterns hold at the 1% and 10% levels, with empirical rejection rates closely aligned with nominal values across the board. These findings indicate that the standardised test statistic
$T_{n}=\frac{\hat{\tau}_{n}-\tau_{n}}{\hat{\sigma}_{n}}$
is approximately standard normal under the null, consistent with the asymptotic normality established in Corollary 1.
The simulation confirms the unbiasedness, consistency, and inferential validity of the Riesz estimator in finite samples. The plug-in variance estimator successfully accounts for local dependence induced by block-level shocks, even when correlations are signed and heterogeneous. The method performs well across all sample sizes and dependency regimes considered, providing empirical support for its practical application.
7 Conclusion
This paper extends the design-based framework for causal inference to accommodate random potential outcomes. By introducing latent randomness through a variable $\omega$, we allow for stochastic variation in outcomes while preserving the core identification principle of randomised treatment assignment. Our formulation maintains the key advantage of design-based methods, minimal reliance on outcome modelling, while generalising the standard framework to address the inherent stochasticity of modern experimental data.
We develop a general asymptotic theory for the aggregate Riesz estimator under local dependence. A central strength of the results lies in their generality: neither the estimator $\hat{\tau}_{n}(z,\omega)$ nor the variance $\sigma_{n}^{2}$ is required to converge to a fixed distribution or limiting value. Instead, the asymptotic theory accommodates growing network complexity, heterogeneous variance behaviour, and non-stationary dependence structures.
Theorems 3.1, 3.2, and 4.1, along with Corollary 1, establish a robust framework for inference using the Riesz estimator. For instance, Theorem 3.2 shows that the standardised statistic
$\sigma_{n}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)$
is asymptotically standard normal without requiring convergence of the variance sequence. Similarly, Theorem 4.1 demonstrates that $n\hat{\sigma}_{n}^{2}$ consistently estimates $n\sigma_{n}^{2}$, even when the latter does not stabilise. These results enable valid inference in experimental settings with increasing dependence, network size, or covariate complexity.
Theorem 3.1 establishes mean-square consistency under mild moment and neighbourhood growth assumptions. In cases where the average neighbourhood size is uniformly bounded, the estimator achieves an $O_{p}(n^{-1/2})$ rate of convergence, without relying on classical asymptotic distributions. Corollary 1 provides a feasible and fully data-driven test statistic that converges to the standard normal, provided regularity conditions are satisfied and the plug-in variance estimator is consistent.
To evaluate the finite-sample behaviour of the method, we conduct a comprehensive simulation study based on locally dependent data generated via signed block-level shocks. The results confirm the theoretical properties: the Riesz estimator is unbiased and consistent, with RMSE declining as sample size increases. Confidence intervals based on the plug-in standard error exhibit accurate or slightly conservative coverage across all dependency settings. Moreover, empirical rejection rates for two-sided Wald tests closely match nominal levels, even in moderate sample sizes. These findings validate the practical applicability of the proposed estimator in realistic experimental designs.
In summary, the Riesz representation approach remains valid and powerful in the presence of outcome-level randomness and structured dependence. The framework is robust and flexible, enabling valid inference without restrictive asymptotic assumptions. It is particularly well-suited to modern experiments involving sensors, adaptive interventions, high-dimensional treatments, or biological systems, where stochastic variation is fundamental rather than incidental.
Future research may extend the framework to incorporate non-linear or non-smooth functionals, covariate adjustment in observational designs, or high-dimensional and non-parametric model spaces. We believe the tools developed here provide a foundation for a new class of design-based estimators that remain grounded in randomisation principles while adapting to the complexity of real-world data.
Appendix A Notation Summary
Table 1: Summary of commonly used notation throughout the paper.
Symbol Description
$n$ Number of units in the experiment
$z\in\mathcal{Z}$ Treatment assignment vector or design variable
$\mu$ Treatment vector drawn from known randomisation distribution $\mu$
$\omega\in\Omega$ Latent variable representing outcome-level randomness
$\tilde{y}_{i}(z,\omega)$ Potential outcome of unit $i$ under treatment $z$ and latent variable $\omega$
$\theta_{i}$ Linear functional defining the treatment effect for unit $i$
$\tau_{n}$ Population- or sample-level estimand (e.g., average treatment effect)
$\hat{\tau}_{n}$ Aggregate Riesz estimator
$\psi_{i}(z,\omega)$ Riesz representer for unit $i$
$\hat{\psi}_{i}(z,\omega)$ Estimated Riesz representer (e.g., via basis expansion)
$\mathcal{M}_{i}$ Model space for unit $i$’s outcome function
$N_{i}$ Dependency neighbourhood of unit $i$
$D_{n}$ Maximum neighbourhood size across units
$d_{n}$ Average neighbourhood size across units
$\sigma_{n}^{2}$ Variance of $\hat{\tau}_{n}$
$\hat{\sigma}_{n}^{2}$ Estimated variance of $\hat{\tau}_{n}$
Appendix B Algorithms
Algorithm 1 Computing the Riesz Representer $\psi_{i}(z,\omega_{0})$
Inputs:
$\bullet$ Basis functions $\{g_{i,1},\dots,g_{i,m}\}$ for $\mathcal{M}_{i}$
$\bullet$ Known linear functional $\theta_{i}$
$\bullet$ Randomisation distribution for $z$
Step 1: Compute Gram Matrix $\hat{\mathbf{G}}_{i}$
$[\hat{\mathbf{G}}_{i}]_{\ell,k}=\mathbb{E}_{z}\bigl{[}g_{i,\ell}(z,\omega_{0})g_{i,k}(z,\omega_{0})\bigr{]}$
Step 2: Compute Target Vector $\hat{\mathbf{T}}_{i}$
$[\hat{\mathbf{T}}_{i}]_{\ell}=\theta_{i}(g_{i,\ell})$
Step 3: Solve Linear System
$\hat{\mathbf{G}}_{i}\bm{\beta}_{i}=\hat{\mathbf{T}}_{i}\quad\Rightarrow\quad\bm{\beta}_{i}=\hat{\mathbf{G}}_{i}^{-1}\hat{\mathbf{T}}_{i}$
Output: Estimated Riesz representer $\hat{\psi}_{i}(z,\omega_{0})$
Algorithm 2 Approximating the Riesz Representer $\psi_{i}(z,\omega_{0})$ in Infinite-Dimensional Settings
Inputs:
$\bullet$ Truncated orthonormal basis $\{e_{1},\dots,e_{m}\}$ of $\mathcal{M}_{i}$
$\bullet$ Known linear functional $\theta_{i}$
$\bullet$ Randomisation distribution for $z$
$\bullet$ Regularisation parameter $\lambda>0$
Step 1: Compute Gram Matrix $\hat{\mathbf{G}}_{i}^{(m)}$
$[\hat{\mathbf{G}}_{i}^{(m)}]_{\ell,k}=\mathbb{E}_{z}\bigl{[}e_{\ell}(z,\omega_{0})\,e_{k}(z,\omega_{0})\bigr{]}$
Step 2: Compute Target Vector $\hat{\mathbf{T}}_{i}^{(m)}$
$[\hat{\mathbf{T}}_{i}^{(m)}]_{\ell}=\theta_{i}(e_{\ell})$
Step 3: Solve Regularised Linear System
$\hat{\mathbf{G}}_{i}^{(m)}\bm{\beta}_{i}=\hat{\mathbf{T}}_{i}^{(m)}\quad\Rightarrow\quad\bm{\beta}_{i}=\left(\hat{\mathbf{G}}_{i}^{(m)}\right)^{-1}\hat{\mathbf{T}}_{i}^{(m)}$
Output: Estimated Riesz representer
$\hat{\psi}_{i}^{(m)}(z,\omega_{0})=\sum_{k=1}^{m}\beta_{i,k}\,e_{k}(z,\omega_{0})$
Appendix C Figures and Table
Figure 1: Bias and RMSE of estimators across varying neighbourhood growth rates $d$.
[…]
Figure 2: Coverage of 95% confidence intervals under varying dependency growth rates $d$.
0.920.940.960.9800.10.20.250.3Dependency growth rate $d$$n=$1002005001000
Table 2: Empirical rejection rates at 1%, 5%, and 10% significance levels across sample sizes $n$ for each dependency growth rate $d$. Based on 2000 replications using the plug-in variance estimator.
Sig. levels $d$ $n=100$ $n=200$ $n=500$ $n=1000$
1% 0.00 0.0075 0.0060 0.0060 0.0050
0.10 0.0100 0.0050 0.0050 0.0070
0.20 0.0100 0.0065 0.0045 0.0085
0.25 0.0115 0.0100 0.0075 0.0055
0.30 0.0155 0.0100 0.0100 0.0060
5% 0.00 0.0470 0.0345 0.0325 0.0380
0.10 0.0440 0.0380 0.0410 0.0335
0.20 0.0450 0.0410 0.0390 0.0450
0.25 0.0455 0.0480 0.0425 0.0345
0.30 0.0510 0.0480 0.0415 0.0435
10% 0.00 0.0930 0.0820 0.0785 0.0825
0.10 0.0870 0.0810 0.0830 0.0795
0.20 0.0865 0.0855 0.0855 0.0850
0.25 0.0865 0.0950 0.0930 0.0760
0.30 0.0955 0.0905 0.0940 0.0920
Appendix D Lemmas and Proofs of Propositions and Theorems
Proof of Proposition 1.
The linearity for both the two examples follows directly from the facts that integrals are linear and that we are taking differences of linear functionals.
For continuity, we need to rewrite (2.8) clearly
$\theta_{i}(u)=\iint_{\mathcal{Z}\times\Omega}\left(\frac{\mathbf{I}_{A}(z)}{\mu(A)}-\frac{\mathbf{I}_{B}(z)}{\mu(B)}\right)u(z,\omega)\,\mu(\mathrm{d}z)\,P(\mathrm{d}\omega)=\left\langle u,\,\varphi\right\rangle,$
where $\mathbf{I}$ stands for indicator function, and
$\varphi(z,\omega)=\frac{\mathbf{I}_{A}(z)}{\mu(A)}-\frac{\mathbf{I}_{B}(z)}{\mu(B)}.$
We can see that $\varphi\in L^{2}(\mathcal{Z}\times\Omega)$ as
$\|\varphi\|_{L^{2}}^{2}$ $=$ $\int_{\Omega}\int_{\mathcal{Z}}\left|\frac{\mathbf{I}_{A}(z)}{\mu(A)}-\frac{\mathbf{I}_{B}(z)}{\mu(B)}\right|^{2}\,\mu(\mathop{}\!\mathrm{d}z)\,P(\mathop{}\!\mathrm{d}\omega)$
$=$ $\int_{\mathcal{Z}}\left|\frac{\mathbf{I}_{A}(z)}{\mu(A)}-\frac{\mathbf{I}_{B}(z)}{\mu(B)}\right|^{2}\,\mu(\mathrm{d}z)=\frac{1}{\mu(A)}+\frac{1}{\mu(B)}<\infty.$
Now continuity is immediate due to the Cauchy-Schwarz inequality
$|\theta_{i}(u)|=|\langle u,\varphi\rangle|\leq\left\lVert u\right\rVert\,\left\lVert\varphi\right\rVert.$
Proof of Proposition 2.
Derivatives and expectation are linear, which establishes linearity.
Since $u\in L^{2}(\Omega;H^{1}(\mathcal{Z}))$, it follows that for almost every $\omega\in\Omega$, the function $z\mapsto u(z,\omega)\in H^{1}(\mathcal{Z})$, and hence the weak derivative $\partial u/\partial z$ exists and is continuous almost everywhere in $z$. Moreover, for any interior point $z_{0}\in\mathcal{Z}$, the mapping $\omega\mapsto\partial u/\partial z(z_{0},\omega)$ belongs to $L^{2}(\Omega)$, with
$\left\|\frac{\partial u}{\partial z}(z_{0},\cdot)\right\|_{L^{2}(\Omega)}\leq C\left\|\frac{\partial u}{\partial z}\right\|,$
for some constant $C>0$, by the Sobolev embedding theorem and the continuity of the point evaluation map on $H^{1}(\mathcal{Z})$. See Adams and Fournier (2003, Theorem 5.36) and related discussion of pointwise evaluation in $H^{1}$ spaces.
Applying the Cauchy–Schwarz inequality gives
$|\theta_{i}(u)|=\left|\int_{\Omega}\frac{\partial u}{\partial z}(z_{0},\omega)\,P(\mathrm{d}\omega)\right|\leq\left\|\frac{\partial u}{\partial z}(z_{0},\cdot)\right\|_{L^{2}(\Omega)},$
and this completes the proof. ∎
Proof of Theorem 2.1.
See Riesz (1907). ∎
Proof of Theorem 2.2.
Assumptions 1, 3-4 imply the existence of $\psi_{i}\in\mathcal{M}_{i}$ (Theorem 2.1), and (2.13) and (2.14) imply $\mathbb{E}\bigl{[}\hat{\theta}_{i}(z,\omega)\bigr{]}=\theta_{i}(\tilde{y}_{i})$, which further implies that
$\mathbb{E}\Bigl{[}\hat{\tau}_{n}(z,\omega)\Bigr{]}=\sum_{i=1}^{n}\nu_{ni}\,\mathbb{E}\bigl{[}\hat{\theta}_{i}(z,\omega)\bigr{]}=\sum_{i=1}^{n}\nu_{ni}\,\theta_{i}(\tilde{y}_{i})=\tau_{n}.$
Lemma 1.
Suppose the weights $\nu_{ni}$ satisfy Assumption 5. Then for any integer $r\geq 1$, the following bound holds:
$\sum_{i=1}^{n}\nu_{ni}^{r}=O(n^{1-r}).$ (D.1)
Proof.
By Assumption 5, we have
$\sum_{i=1}^{n}\nu_{ni}^{r}\leq\sum_{i=1}^{n}\left(\frac{\bar{\nu}}{n}\right)^{r}=n\cdot\left(\frac{\bar{\nu}}{n}\right)^{r}=\bar{\nu}^{r}\cdot n^{1-r},$
which implies the claimed order bound. ∎
Lemma 2.
Under Assumptions 1, 3, and 4, the $r$th central moment of the treatment effect Riesz estimator, for some integer $r\geq 1$, if it exists, satisfies
$\mathbb{E}\left|\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right|^{r}\leq\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p}\,\left\lVert\psi_{i}\right\rVert_{q}\right)^{r},$ (D.2)
for any $p,q\in[1,\infty]$ satisfying $1/p+1/q=1/r$. In particular, when $r=2$,
$\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{2}\leq\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2}.$ (D.3)
The inequalities (D.2) and (D.3) hold trivially if either $\|\tilde{y}_{i}\|_{p}$ or $\|\psi_{i}\|_{q}$ is infinite.
Proof.
Inequality (D.2) is a direct extension of Lemma C.3 in HWS, which handles the case without the additional randomness $\omega$. The proof carries through immediately by applying Hölder’s inequality on the product space $\mathcal{Z}\times\Omega$, since the additional integration over $\omega$ preserves the inequality structure.
Regarding inequality (D.3),
$\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{2}$ $=\mathbb{E}[\hat{\theta}_{i}(z,\omega)]^{2}-(\theta_{i}(\tilde{y}_{i}))^{2}\leq\mathbb{E}\left[\tilde{y}_{i}(z,\omega)\psi_{i}(z,\omega)\right]^{2}$
$=\left\lVert\tilde{y}_{i}(z,\omega)\psi_{i}(z,\omega)\right\rVert^{2}\leq\left(\left\lVert\tilde{y}_{i}(z,\omega)\right\rVert_{p}\,\left\lVert\psi_{i}(z,\omega)\right\rVert_{q}\right)^{2}.$ (Hölder)
Proof of Proposition 3.
Let $\zeta_{i}=\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$ and we observe that $\mathbb{E}[\zeta_{i}]=0$ by Theorem 2.2. Then,
$\sigma_{n}^{2}$ $=\mathbb{V}[\hat{\tau}_{n}(z,\omega)-\tau_{n}]=\mathbb{V}\left[\sum_{i=1}^{n}\nu_{ni}\zeta_{i}\right]=\mathbb{E}\left[\sum_{i=1}^{n}\nu_{ni}\zeta_{i}\right]^{2}$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\nu_{ni}\nu_{nj}\mathbf{I}(i\in N_{j})\mathbf{I}(j\in N_{i})\mathbb{E}\left[\zeta_{i}\zeta_{j}\right]$
$\leq\sum_{i=1}^{n}\sum_{j=1}^{n}\nu_{ni}\nu_{nj}\mathbf{I}(i\in N_{j})\mathbf{I}(j\in N_{i})\left\lVert\zeta_{i}\right\rVert\,\left\lVert\zeta_{j}\right\rVert$ (Cauchy-Schwartz)
$\leq\sum_{i=1}^{n}\nu_{ni}\sum_{j=1}^{n}\frac{1}{2}\nu_{nj}\left(\mathbf{I}(j\in N_{i})\left\lVert\zeta_{i}\right\rVert^{2}+\mathbf{I}(i\in N_{j})\left\lVert\zeta_{j}\right\rVert^{2}\right)$ (AM-GM)
$=\sum_{i=1}^{n}\nu_{ni}\left(\sum_{j\in N_{i}}\nu_{nj}\right)\left\lVert\zeta_{i}\right\rVert^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\left\lVert\zeta_{i}\right\rVert^{2}$ (Lemma 1 (2.17))
$\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\,\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2}$ (Lemma 2 (D.3))
Note that AM-GM stands for arithmetic-geometric mean inequality.
To obtain the uniform bound, we apply the max-$p$ norm, as defined in (3.2), and note that $d_{n}=n^{-1}\sum_{i=1}^{n}|N_{i}|$, yielding
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\left(\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert=\frac{\bar{\nu}^{2}d_{n}}{n}\left(\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$
Since this inequality holds for any $p,q\in[1,\infty]$ satisfying $21/p+1/q=1/$ 2, we have a more conservative upper bound
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}d_{n}}{n}\left(\inf_{p,q\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$
Proof of Theorem 3.1.
The result follows directly from Proposition 3. Specifically, the bound in (3.4), together with Assumption 8 with $r=2$, implies that
$\mathbb{E}\left[\hat{\tau}_{n}(z,\omega)-\tau_{n}\right]^{2}\to 0,$
as $n\to\infty$, establishing mean square consistency.
Moreover, suppose $\sup_{n\in\mathbb{N}}d_{n}<\infty$, that is, there exists some $d_{0}\in(0,\infty)$ such that $d_{n}\leq d_{0}$ for all $n$. Since $\hat{\tau}_{n}-\tau_{n}$ follows some distribution with zero mean and variance $\sigma_{n}^{2}$, it follows that $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ has variance $n\sigma_{n}^{2}$. By the bound in Proposition 3, we obtain
$n\sigma_{n}^{2}\leq\bar{\nu}^{2}d_{0}\left(\inf_{p,q\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}<\infty,$
under the stated assumptions. Hence, the aggregate Riesz estimator satisfies $O_{p}(n^{-1/2})$. ∎
Lemma 3.
[…]
Proof.
We will show that $\mathbb{E}[f(W_{n})]\to\mathbb{E}[f(Z)]$ for all bounded uniformly continuous functions $f:X\to\mathbb{R}$, which characterizes weak convergence; see Billingsley (1999, Theorem 2.1).
Let $f\in\mathrm{BUC}(X)$, the space of all bounded uniformly continuous functions. Since Lipschitz functions are dense in $\mathrm{BUC}(X)$ under the supremum norm, for any $\varepsilon>0$, there exists a Lipschitz function $f_{\varepsilon}:X\to\mathbb{R}$ such that
$\|f-f_{\varepsilon}\|_{\infty}<\frac{\varepsilon}{3}.$
The function $f_{\varepsilon}$ is Lipschitz with some finite constant $L_{\varepsilon}$. Therefore,
$\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|\leq L_{\varepsilon}\cdot d_{W}(W_{n},Z).$
Since $d_{W}(W_{n},Z)\to 0$, there exists $N(\varepsilon)\in\mathbb{N}$ such that
$\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|<\frac{\varepsilon}{3},\quad\text{for all }n>N(\varepsilon).$
Using the triangle inequality, we have
$\left|\mathbb{E}[f(W_{n})]-\mathbb{E}[f(Z)]\right|$ $\leq\left|\mathbb{E}[f(W_{n})]-\mathbb{E}[f_{\varepsilon}(W_{n})]\right|$
$\quad+\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|$
$\quad+\left|\mathbb{E}[f_{\varepsilon}(Z)]-\mathbb{E}[f(Z)]\right|$
$<\frac{\varepsilon}{3}+\frac{\varepsilon}{3}+\frac{\varepsilon}{3}=\varepsilon.$
[…]
Proof of Theorem 3.2.
The proof consists of two steps. First, we apply Stein’s method in the seminal paper Stein (1972), and Theorem 3.5 from Ross (2011) to establish Wasserstein convergence of a sequence of measures, following the approach of HWS but under a different setup. Second, we show that this Wasserstein convergence implies weak convergence of the sequence of measures.
Clearly $\mathbb{E}[\hat{\theta}_{i}(z,\omega)]^{4}<\infty$ is equivalent to $\mathbb{E}[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})]^{4}<\infty$. Define $X_{i}=\nu_{ni}(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i}))$, and then $\mathbb{E}[X_{i}^{4}]<\infty$. By Theorem 2.2, we have also $\mathbb{E}[X_{i}]=0$.
Note that $\sum_{i=1}^{n}X_{i}=\hat{\tau}_{n}(z,\omega)-\tau_{n}$. Define $W_{n}=\sum_{i=1}^{n}X_{i}/\sigma_{n}$. Since the collection $\{X_{1},X_{2},...,X_{n}\}$ have dependency neighborhoods $N_{i}$, $i=1,...,n$, with $D_{n}=\max_{1\leq i\leq n}|N_{i}|$. Then we can now apply Theorem 3.5 in Ross (2011) and obtain
$d_{W}(W_{n},Z)\leq\frac{D_{n}^{2}}{\sigma_{n}^{3}}\sum_{i=1}^{n}\mathbb{E}|X_{i}|^{3}+\frac{\sqrt{26}D_{n}^{3/2}}{\sqrt{\pi}\sigma_{n}^{2}}\sqrt{\sum_{i=1}^{n}\mathbb{E}[X_{i}^{4}]},$ (D.4)
where $d_{W}(W_{n},Z)$ denotes the Wasserstein distance between $W_{n}$ and a standard normal one $Z$. Observe that (D.4) gives an upper bound of the distance.
First, we examine the first term in (D.4)
$\sum_{i=1}^{n}\mathbb{E}\left\lvert X_{i}\right\rvert^{3}$ $=\sum_{i=1}^{n}\mathbb{E}\left\lvert\nu_{ni}\left(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right)\right\rvert^{3}$
$=\sum_{i=1}^{n}\nu_{ni}^{3}\mathbb{E}\left\lvert\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right\rvert^{3}$
$\leq\sqrt{\sum_{i=1}^{n}\nu_{ni}^{6}}\cdot\sqrt{\sum_{i=1}^{n}\left(\mathbb{E}\left\lvert\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right\rvert^{3}\right)^{2}}$ (Cauchy-Schwartz)
$\leq n^{-5/2}\bar{\nu}^{3}\cdot\sqrt{\sum_{i=1}^{n}\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p_{1}}\,\left\lVert\psi_{i}\right\rVert_{q_{1}}\right)^{6}}$ (Lemma 1 and 2)
$\leq n^{-2}\bar{\nu}^{3}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}\right)^{3}.$
Next, we examine the second term in (D.4).
$\sum_{i=1}^{n}\mathbb{E}[X_{i}]^{4}$ $=\sum_{i=1}^{n}\mathbb{E}\left[\nu_{ni}\left(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right)\right]^{4}$
$=\sum_{i=1}^{n}\nu_{ni}^{4}\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{4}$
$\leq\sqrt{\sum_{i=1}^{n}\nu_{ni}^{8}}\cdot\sqrt{\sum_{i=1}^{n}\left(\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{4}\right)^{2}}$ (Cauchy-Schwartz)
$\leq n^{-7/2}\bar{\nu}^{4}\cdot\sqrt{\sum_{i=1}^{n}\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p_{2}}\,\left\lVert\psi_{i}\right\rVert_{q_{2}}\right)^{8}}$ (Lemma 1 and 2)
$\leq n^{-3}\bar{\nu}^{4}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}\right)^{4}.$
Thus, we obtain
$d_{W}(W_{n},Z)$ $\leq\frac{\bar{\nu}^{3}D_{n}^{2}}{n^{2}\sigma_{n}^{3}}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}\right)^{3}+\frac{\sqrt{26}\bar{\nu}^{2}D_{n}^{3/2}}{\sqrt{\pi}n^{3/2}\sigma_{n}^{2}}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}\right)^{2}$
$=O\left(\frac{D_{n}^{2}}{n^{2}\sigma_{n}^{3}}\right)+O\left(\frac{D_{n}^{3/2}}{n^{3/2}\sigma_{n}^{2}}\right),$ (D.5)
under Assumption 8, which states that $\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}$, $\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}$, $\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}$, and $\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}$ converge for some $p_{1},q_{1},p_{2},q_{2}\in[1,\infty]$ satisfying $31/p_{1}+1/q_{1}=1/$ 3 and $41/p_{2}+1/q_{2}=1/$ 4.
Given Assumption 9, (D.5) can be further bounded by
$d_{W}(W_{n},Z)\leq O\left(\frac{D_{n}^{2}}{n^{1/2}}\right)+O\left(\frac{D_{n}^{3/2}}{n^{1/2}}\right),$ (D.6)
which implies that, if $D_{n}=o(n^{1/4})$, then $d_{W}(W_{n},Z)\to 0$.
Lemma 3 shows that Wasserstein convergence implies weak convergence, which completes the proof. ∎
Lemma 4.
Let $\{\zeta_{i}\}_{i=1}^{n}$ be random variables such that
$\mathbb{E}[\zeta_{i}]=0\quad\text{and}\quad\sup_{i\in\mathbb{N}}\mathbb{E}[\zeta_{i}^{4}]<\infty.$
For deterministic weights $\nu_{ni}$ satisfying Assumption 5, define
$\Delta_{n}\;=\;n\sum_{(i,j)\in\mathcal{E}_{n}}\Bigl{(}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}-\nu_{ni}\nu_{nj}\,\mathbb{E}[\zeta_{i}\zeta_{j}]\Bigr{)}.$
Then
$\Delta_{n}\;=\;O_{p}\bigl{(}n^{-1/2}\,D_{n}^{3/2}\bigr{)}.$
Proof.
Set
$X_{ij}\;:=\;n\,\nu_{ni}\nu_{nj}\bigl{(}\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}]\bigr{)},\qquad(i,j)\in\mathcal{E}_{n}.$
Because $\nu_{ni}\leq\bar{\nu}/n$, $n\nu_{ni}\nu_{nj}\leq\bar{\nu}^{2}/n.$ Then for any two dependent pairs $(i,j),(l,k)\in\mathcal{E}_{n}$, we apply Cauchy-Schwarz and the uniform fourth moment condition $\sup_{i}\mathbb{E}[\zeta_{i}^{4}]<\infty$ to get
$\operatorname{Cov}[X_{ij},X_{lk}]\leq\Bigl{(}\tfrac{\bar{\nu}^{2}}{n}\Bigr{)}^{2}\mathbb{E}\bigl{[}(\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}])(\zeta_{l}\zeta_{k}-\mathbb{E}[\zeta_{l}\zeta_{k}])\bigr{]}\leq\frac{\bar{\nu}^{4}C}{n^{2}}$
for some constant $C>0$. Note that $(i,j),(l,k)$ can be the same pair.
Fix $(i,j)\in\mathcal{E}_{n}$. A second pair $(k,\ell)$ can be dependent with $(i,j)$ only if it lies within the corresponding neighbourhoods, that is, $(k,\ell)\cap(N_{i}\cup N_{j})\neq\emptyset$. Each unit has at most $D_{n}$ neighbours, so there are at most $D_{n}^{2}-1=O(D_{n}^{2})$ such pairs. Hence every $X_{ij}$ is dependent with at most $C_{1}D_{n}^{2}$ other summands for some constant $C_{1}>0$.
There are at most $nD_{n}$ ordered pairs in $\mathcal{E}_{n}$, so
$\operatorname{Var}(\Delta_{n})\;\leq\;C_{1}D_{n}^{2}\cdot nD_{n}\cdot\frac{\bar{\nu}^{4}C}{n^{2}}\;=\;C_{2}\,\bar{\nu}^{4}\,\frac{D_{n}^{3}}{n},$
for a constant $C_{2}>0$.
By Chebyshev’s inequality, for any $\varepsilon>0$,
$\mathbb{P}\!\bigl{(}|\Delta_{n}|>\varepsilon\bigr{)}\;\leq\;\frac{\operatorname{Var}(\Delta_{n})}{\varepsilon^{2}}\;\leq\;\frac{C_{2}\,\bar{\nu}^{4}}{\varepsilon^{2}}\,\frac{D_{n}^{3}}{n}.$
Since the right-hand side is of order $n^{-1}\,D_{n}^{3},$ we have
$\Delta_{n}\;=\;O_{p}\bigl{(}n^{-1/2}\,D_{n}^{3/2}\bigr{)},$
which completes the proof. ∎
Proof of Theorem 4.1.
By Theorem 2.2, we have $\mathbb{E}[\zeta_{i}]=0$, where $\zeta_{i}:=\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$. Assumption 10$(b)$ ensures that $\sup_{i\in\mathbb{N}}\mathbb{E}[\hat{\theta}_{i}(z,\omega)^{4}]<\infty$, which is equivalent to $\sup_{i\in\mathbb{N}}\mathbb{E}[\zeta_{i}^{4}]<\infty$, since $\theta_{i}(\tilde{y}_{i})$ is constant.
Lemma 4 together with Assumption 5 implies that
$n\hat{\sigma}_{n}^{2}-n\sigma_{n}^{2}=n\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\left(\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}]\right)=O_{p}(n^{-1/2}D_{n}^{3/2}).$
If Assumption 6 holds with $D_{n}=o(n^{1/3})$, this expression converges to zero in mean square. ∎
Proof of Corollary 1.
The result follows directly from Theorems 3.2 and 4.1. We write
$\hat{\sigma}_{n}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)=\frac{\sqrt{n}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)}{\sqrt{n}\sigma_{n}}\cdot\frac{\sqrt{n}\sigma_{n}}{\sqrt{n}\hat{\sigma}_{n}}.$
The first term converges in distribution to $\mathcal{N}(0,1)$ by Theorem 3.2, and the second term converges to 1 in probability by Theorem 4.1, together with Assumption 9, which guarantees that $\sqrt{n}\sigma_{n}$ is bounded away from zero.
Since the product of a sequence converging in distribution and another converging in probability (to a constant) also converges in distribution, Slutsky’s theorem yields the desired result. ∎