### v5

# Design-Based Inference under Random Potential Outcomes via Riesz Representation

Abstract
We introduce a design-based framework for causal inference that accommodates random potential outcomes, thereby extending the classical Neyman-Rubin model in which outcomes are treated as fixed. Each unit’s potential outcome is modelled as a structural mapping $\tilde{y}_{i}(z,\omega)$, where $z$ denotes the treatment assignment and $\omega$ represents latent outcome-level randomness. Inspired by recent connections between design-based inference and the Riesz representation theorem, we embed potential outcomes in a Hilbert space and define treatment effects as linear functionals, yielding estimators constructed via their Riesz representers. This approach preserves the core identification logic of randomised assignment while enabling valid inference under stochastic outcome variation. We establish large-sample properties under local dependence and develop consistent variance estimators that remain valid under weaker structural assumptions, including partially known dependence. A simulation study illustrates the robustness and finite-sample behaviour of the estimators. Overall, the framework unifies design-based reasoning with stochastic outcome modelling, broadening the scope of causal inference in complex experimental settings.
Keywords: Design-based causal inference; Random potential outcomes; Riesz representation; Local dependence; Variance estimation; Simulation study.
1 Introduction
Randomised experiments remain central to the identification of causal effects. In the classical design-based perspective, pioneered by Neyman (1990) and Rubin (1974), potential outcomes are treated as fixed quantities, with all randomness attributed to the treatment assignment mechanism. This approach has given rise to influential estimators such as the Horvitz-Thompson estimator, which rely solely on the experimental design for identification.
Building on this foundation, several strands of recent literature have extended design-based inference to more complex settings. In particular, a substantial body of work addresses interference and dependence among units, including network and spatial designs (Aronow and Samii, 2017; Sävje et al., 2021; Athey et al., 2021). These methods typically retain the fixed-potential-outcomes assumption and model dependence through exposure mappings or stratified interference structures. Other related contributions focus on inference under model misspecification or robust estimation in observational studies (Abadie et al., 2020; Imbens, 2004), but rely on outcome modelling rather than randomisation for identification.
A limitation of the classical design-based framework is its assumption that potential outcomes are deterministic. However, in many applied settings, outcomes may vary randomly even when treatment assignments are fixed. This variability can arise from factors such as measurement noise or random physiological responses.
We propose a new design-based framework in which potential outcomes are modelled as random functions of the treatment assignment $z$ and a latent variable $\omega$, written as $\tilde{y}_{i}(z,\omega)$. This formulation is motivated by both practical and conceptual considerations. Many modern experiments, such as in biology or sensor-based fieldwork, exhibit intrinsic outcome-level variability even under identical treatments. More fundamentally, the very notion of an average treatment effect presumes a distribution over units or outcomes, suggesting that potential outcomes should be treated as random rather than fixed.
Our framework retains the design-based identification logic through randomised treatment assignment, assuming the treatment vector is randomly drawn from a known distribution $\mu$. By making outcome-level randomness explicit, the approach extends the classical Neyman-Rubin model and bridges design-based reasoning with super-population perspectives, offering a unified structure for causal inference in complex experimental environments.
Our analysis is confined to experimental settings with known randomisation schemes. We do not consider observational designs, where unconfoundedness must be justified via covariate adjustment. The framework allows for spillover effects and continuous treatments.
Recent contributions, notably by Harshaw, Wang, and Sävje (2022, hereafter HWS), have reframed the classical design-based approach through the Riesz representation theorem. By modelling potential outcomes as elements of a Hilbert space induced by the design, and defining treatment effects as linear functionals on that space, they construct a unified framework that encompasses standard weighting estimators and extends naturally to settings with interference or continuous treatments. Their asymptotic results on consistency and normality draw upon foundational ideas from Stein’s method framework (Stein, 1972; Ross, 2011) and the analysis of dependency graphs as formalised by Ross (2011), which they adapt to the design-based context.
This paper draws methodological inspiration from the Riesz-based framework introduced by HWS. While our approach broadly follows the logic of their estimator, namely, representing treatment effect functionals via projection in a Hilbert space, we work within a fundamentally different modelling context. In particular, we consider random potential outcomes defined over a latent probability space, in contrast to the fixed-outcome formulation adopted by HWS. As a result, all assumptions, estimands, and inferential results are formulated independently to suit the stochastic framework introduced here.
The Riesz representation theorem remains central: under minimal regularity conditions, it ensures that treatment effects can still be expressed as inner products with a unique representer function, even when potential outcomes are random.
The resulting Riesz estimator remains unbiased and consistent under local dependence. In addition, it satisfies asymptotic normality, enabling valid inference from a single experimental realisation. These asymptotic results draw upon ideas from Stein’s method and dependency graph theory, as also employed in HWS, but are adapted here to a random potential outcome setting.
To implement valid hypothesis testing based on the Riesz estimator, one must input the true variance for the corresponding test statistic. Previous work has focused on deriving upper bounds on the variance that are sharp under strong conditions, but these bounds are not directly usable for inference and tend to be overly conservative in practice. In contrast, we develop a local-dependence variance estimator that does not require assumptions stronger than those already needed for asymptotic normality. The only additional requirement is some structural knowledge about local dependence, which is often realistic in networked, spatial, or blocked experimental designs.
In addition to the local-dependence variance estimator, we propose correlation-based and conservative set-based alternatives that yield valid inference when uncorrelated unit pairs can be reasonably approximated. This version is computationally simpler and circumvents the need to estimate weak or negligible cross-terms, making it especially practical in settings with limited structural information.
Notably, these variance estimators are also applicable in fixed-outcome settings, thereby strengthening previous frameworks by providing a practical means of conducting inference based on a single realisation of the experimental design.
A central insight of our framework is that it formalises a form of sample-based ergodicity under local dependence: averaging across units within a single experimental realisation can substitute for averaging across multiple hypothetical repetitions of the stochastic environment. Despite the presence of latent randomness in potential outcomes, our results show that one need not observe multiple parallel worlds to conduct valid inference. Instead, under appropriate moment and dependence assumptions, a single realisation of the experiment suffices for consistent estimation and asymptotically valid uncertainty quantification.
We further discuss implementation via basis expansions, and conduct a simulation study to evaluate the finite-sample performance of the estimator. The results confirm key theoretical properties, such as unbiasedness, consistency, and valid coverage, and demonstrate that the method remains robust across a range of dependency structures.
Throughout, we use “random” to refer to outcome-level variation due to latent variables, and “stochastic” to describe modelling frameworks that explicitly incorporate such randomness.
The paper is organised as follows. Section 2 introduces the model and basic assumptions. Section 3 establishes large-sample properties of the estimator under local dependence. Section 4 presents feasible variance estimators and discusses their consistency. Section 5 covers computational aspects of estimating the Riesz representer in both finite- and infinite-dimensional settings. Section 6 provides a simulation study. Section 7 concludes with a discussion of the broader implications and possible extensions.
2 Model Setup
2.1 The Potential Outcomes
We begin by considering a setting with $n$ units, where
Assumption 1 (The Stochastic Setting).
each unit’s potential outcome is modelled as a general measurable function $y_{i}(z,x_{i},\epsilon_{i})$ with the following arguments: $z=(z_{1},\dots,z_{n})$ denoting the vector of treatment assignments; $x_{i}=x_{i}(\omega)$ representing the possibly observed covariates for unit $i$; $\epsilon_{i}=\epsilon_{i}(\omega)$ denoting the idiosyncratic error or unobserved heterogeneity for unit $i$. Here, $\omega\in\Omega$ is a latent random element defined on a common probability space $(\Omega,\mathcal{F}_{\omega},P)$ shared by all units.
Furthermore, we define the structural mapping from the latent space $\Omega$ to variables $(x_{i},\epsilon_{i})$ and represent the potential outcome via the composition
$\tilde{y}_{i}(z,\omega)=y_{i}(z,x_{i}(\omega),\epsilon_{i}(\omega)),$ (2.1)
since $x_{i}$ and $\epsilon_{i}$ are measurable functions of $\omega$.
We refer to this as the stochastic setting because the potential outcomes $\tilde{y}_{i}(z,\omega)$ are explicitly modelled as random functions of a latent variable $\omega$, introducing outcome-level randomness into the design-based framework. This formulation enables the analysis of causal effects under a stochastic data-generating process and bridges the gap between classical design-based inference and super-population perspectives.
Although $\omega$ is not unit-specific, its structure permits different subcomponents to influence different units, thereby being able to preserve unit-specific variability and independence and accommodating heterogeneous potential outcomes.
The potential outcome for unit $i$ depends on the entire treatment vector $z$, so that interference (or spillover effects) is permitted. Notably, one may also consider the special case of no spillover effects, whereby
$y_{i}(z,x_{i},\epsilon_{i})=y_{i}(z_{i},x_{i},\epsilon_{i}),$ (2.2)
which corresponds to a version of the Stable Unit Treatment Value Assumption (SUTVA) incorporating consistency, as introduced by Rubin (1980). In the present work we do not impose the no-spillover condition a priori; rather, we allow for general interference and propose to test for its presence.
In practice, it typically suffices to allow heterogeneity across units only through $x_{i}$ and $\epsilon_{i}$, making it unnecessary to distinguish $y_{i}$ explicitly. In such cases, one may work with a common structural function $y$, while still obtaining heterogeneous mappings $\tilde{y}_{i}(z,\omega)=y(z,x_{i},\epsilon_{i})$ across units.
The structural mapping (2.1) is the primary object of interest in what follows.
Denoting $\mathcal{Z}$ as the treatment assignment space, equipped with its Borel sigma-algebra $\mathcal{F}_{z}$, we presume that the structural mapping $\tilde{y}_{i}:\mathcal{Z}\times\Omega\to\mathbb{R}$ is measurable with respect to the sigma-algebra $\mathcal{F}_{z}\otimes\mathcal{F}_{\omega}$ on the domain and the Borel sigma-algebra $\mathcal{B}(\mathbb{R})$ on the codomain. This ensures that the relevant random variables and their expectations are well-defined, and it also serves as a prerequisite for defining inner products and applying Hilbert space tools in the sequel. Note that the measurability of the covariates is sometimes required to ensure that conditional expectations and probabilities given $x_{i}$ are well-defined within the underlying probability space, although we do not explicitly rely on this property in the present paper.
Following randomisation principles, we assume that
Assumption 2 (Randomisation).
the treatment assignment vector $z$ is drawn from a known randomisation distribution that is independent of the outcome-generating process given the latent variable $\omega$.
Assumption 2 is not required for the unbiasedness or asymptotic validity of the Riesz estimator. Its purpose is operational rather than inferential: it ensures that the design distribution $\mu$ remains valid when conditioning on $\omega$, which is essential for defining the Riesz representer via the inner product induced by $\mu$. Once the representer $\psi_{i}$ is constructed, all subsequent inference, including estimation and asymptotic analysis, proceeds without requiring independence between $z$ and $\omega$. This assumption will be invoked again in Section 5, where the Gram matrix is computed in order to estimate the Riesz representer of the potential outcomes.
This assumption also justifies the use of the product measure $\mu(dz)\,P(d\omega)$ in the corresponding integrals. It allows expectations of the form $\mathbb{E}_{z,\omega}[\cdot]$ to be written as iterated expectations $\mathbb{E}_{z}\mathbb{E}_{\omega}[\cdot]$. Throughout, when writing integrals or expectations in this way, we implicitly invoke this assumption.
It is worth noting that this assumption corresponds to the assumption of ignorability or unconfoundedness, as articulated in Rubin (1978) and further developed in Rosenbaum and Rubin (1983) for observational studies. Under this formulation, $\omega$ governs all outcome-related randomness across units, while the randomisation mechanism for $z$ remains externally controlled and separate (see Neyman (1990) and Fisher (1935)).
Conceptually, once a particular realisation of $\omega$ is fixed, each unit’s potential outcome function $\tilde{y}_{i}(z,\omega)$ becomes deterministic in $z$; hence, the only source of randomness in the observed outcomes from the randomised assignment of treatments. This approach preserves the fundamental idea that each unit has a well-defined potential outcome under every treatment assignment, capturing the core principle of SUTVA (or, more generally, the notion of well-defined potential outcomes), even as we permit these outcomes to vary randomly across realisations of $\omega$.
This formulation is flexible and expressive in that it captures both unconfoundedness and confounding, depending on the relationship between $z$ and the latent variable $\omega$. If $z$ and $\omega$ are independent, then $\tilde{y}_{i}(z^{0},\omega)$ and $\tilde{y}_{i}(z^{1},\omega)$ are independent of $z$. More generally, if $z$ and $\omega$ are independent given $x\in X_{0}$, then $y_{i}(z^{0},x,\epsilon_{i})$ and $y_{i}(z^{1},x,\epsilon_{i})$ are independent of $z\mid x\in X_{0}$, for any measurable $X_{0}\subset\mathcal{F}_{x}:=\{A\subseteq x_{i}(\Omega):x_{i}^{-1}(A)\in\mathcal{F}_{\omega}\}$. In randomised experiments, the design ensures that $z$ is independent of $(x_{i},\epsilon_{i})$, mitigating confounding. In observational studies, however, it is imperative to adjust for $x_{i}$ to control for confounding. In either case, this framework models potential outcomes as random rather than fixed.
2.2 Model Space and Inner Product
Let $L^{p}(\mathcal{Z}\times\Omega)$ denote the space of all measurable functions $u:\mathcal{Z}\times\Omega\to\mathbb{R}$, satisfying $\mathbb{E}[|u(z,\omega)|^{p}]<\infty$, for some integer $p\geq 1$. For notational convenience, we henceforth write $L^{p}$ in place of $L^{p}(\mathcal{Z}\times\Omega)$, whenever the meaning is clear from context. Furthermore, We denote the $p$-norm of $u\in L^{p}$ by
$\|u\|_{p}:=\left(\int_{\mathcal{Z}\times\Omega}\left|u(z,\omega)\right|^{p}\mu(\mathop{}\!\mathrm{d}z)P(\mathop{}\!\mathrm{d}\omega)\right)^{1/p}=\left(\mathbb{E}\left|u(z,\omega)\right|^{p}\right)^{1/p},$ (2.3)
where $\mu$ denotes the probability measure for randomisation. The equality in (2.3) relies on Assumption 2. If $z$ and $\omega$ are not independent, the intermediate expression involving the product measure should be omitted, and the $L^{p}$ norm should be defined directly in terms of the joint distribution of $(z,\omega)$. We write $\|u\|:=\|u\|_{2}$ for brevity if $u\in L^{2}$. Note that $p$-norm characterises the $p$-moment of the random variable $u$.
We model each unit’s potential outcome function, via the structural mapping (2.1), as an element of a model space $\mathcal{M}_{i}$, which is a subspace of $L^{2}$. Formally, we write
Assumption 3 (Model Space).
$\tilde{y}_{i}(z,\omega)\in\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega).$ (2.4)
Our construction is formulated in a distinct functional setting that explicitly separates treatment assignments from latent variation induced by unobserved random elements. The model space incorporates both an observable assignment vector $z$ and a latent random element $\omega$, capturing two sources of variation. This separation underpins our interpretation of potential outcomes as random functions and motivates a stochastic modelling framework. In particular, the space $\mathcal{M}_{i}$ is defined over random potential outcomes, and the associated inner product reflects variation due to both assignment and latent structure. Importantly, the spaces from which $z$ and $\omega$ are drawn may differ in structure: while $\omega$ is modelled as a general element of a latent probability space, $z$ is often a finite-dimensional or even discrete vector representing treatment assignments. This structural heterogeneity is explicitly accommodated in our formulation of the product space $\mathcal{Z}\times\Omega$ and the associated function space $\mathcal{M}_{i}$.
The framework of HWS can be viewed as a special case of our model, corresponding to the degenerate setting where $\Omega=\{\omega_{0}\}$ for some $\omega_{0}$. In this case, the latent variable $\omega_{0}$ is effectively fixed, and the realised data serve as the entire outcome-generating mechanism. Thus, the potential outcome function (2.1) becomes a deterministic function of $z$, recovering the setting considered in their work.
We equip $\mathcal{M}_{i}$ with the inner product
$\langle u,v\rangle:=\mathbb{E}_{\omega}\Bigl{[}\mathbb{E}_{z}\bigl{[}u(z,\omega)\,v(z,\omega)\bigr{]}\Bigr{]}=\mathbb{E}\bigl{[}u(z,\omega)\,v(z,\omega)\bigr{]},$ (2.5)
where the second equality follows from the law of iterated expectations. The associated norm is given by $\|u\|=\sqrt{\langle u,u\rangle}$, which coincides with the standard $L^{2}$-norm.
In the degenerate case where $\Omega=\{\omega_{0}\}$, the inner product reduces to
$\langle u,v\rangle^{0}=\mathbb{E}_{z}\bigl{[}u(z,\omega_{0})\,v(z,\omega_{0})\bigr{]},$ (2.6)
which coincides with the inner product employed in HWS. This structure enables us to apply the Riesz representation theorem to construct estimators for treatment effects, even in the presence of interference.
2.3 Treatment Effect Functionals and Riesz Representation
While our model space $\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega)$ may not be complete under the inner product defined in (2.5), this poses no issue in what follows. Any inner product space admits a Hilbert space completion, and the Riesz representation theorem applies in that completed space. Hence, throughout this paper, whenever necessary, we implicitly work with the closure of $\mathcal{M}_{i}$ in $L^{2}(\mathcal{Z}\times\Omega)$.
Let $\mathcal{M}_{i}^{\prime}$ denote the dual space of $\mathcal{M}_{i}$.
Assumption 4 (Dual Representability of the Treatment Effect).
The treatment effect $\theta_{i}$ for each unit $i$ can be represented as a continuous linear functional on the model space of potential outcomes, viz., $\theta_{i}:\mathcal{M}_{i}\to\mathbb{R}$ with $\theta_{i}\in\mathcal{M}_{i}^{\prime}$.
This assumption reflects a standard construction in functional analysis, where a treatment effect is represented as a continuous linear functional on a function space of potential outcomes. Such a formulation enables the use of the Riesz representation theorem, which yields an estimator expressed as an inner product with a representer function. This approach applies to both classical fixed-outcome frameworks and to stochastic settings where potential outcomes depend on latent variables.
Note that functions in an $L^{2}$ space are identified through equivalence classes, meaning two functions differing only on a set of measure zero are considered identical. Consequently, a well-defined linear functional must assign the same value to all functions within the same equivalence class.
A potential pitfall arises when attempting to use evaluation functionals, such as $\theta(u)=u(x_{0})$, that assign values based on a single point. Such evaluation maps generally are not well-defined in an $L^{2}$ context because they may yield different values for functions identical almost everywhere, violating linearity and the identification of elements in the same equivalence class.
In our analysis, we explicitly avoid these issues by requiring that the treatment effect functional $\theta_{i}$ be defined consistently on equivalence classes. More specifically, if two functions differ by a function of zero $L^{2}$-norm (i.e., distance zero), the functional $\theta_{i}$ shall assign them the same value.
From Assumption 4, we also require that the linear functional $\theta_{i}$ on a normed space is continuous or equivalently bounded. A heuristic justification is that requiring the treatment effect functional to be continuous ensures that small perturbations in the potential outcome function (as measured by the $L^{2}$ norm) result in small changes in the treatment effect. This continuity is essential for invoking the Riesz representation theorem, which in turn guarantees that $\theta$ can be uniquely represented by an element of the Hilbert space.
Researchers need not be concerned about such technicalities in practice, as they are always free to define or adjust their treatment effect functional to comply with the requirements in Assumptions 4. For instance, one might consider the following examples of continuous linear functionals.
Expected treatment effect for binary treatment
$\theta_{i}(u)\;=\;\mathbb{E}_{\omega}\bigl{[}u(z^{(1)},\omega)-u(z^{(0)},\omega)\bigr{]}$ (2.7)
representing the expected outcome difference when shifting treatment assignment from $z^{(0)}$ to $z^{(1)}$.
Expected treatment effect for treatment intervals or regions
$\theta_{i}(u)\;=\;\mathbb{E}_{\omega}\Bigl{[}\mathbb{E}_{z\in A}\bigl{[}u(z,\omega)\bigr{]}\Bigr{]}-\mathbb{E}_{\omega}\Bigl{[}\mathbb{E}_{z\in B}\bigl{[}u(z,\omega)\bigr{]}\Bigr{]},$ (2.8)
where $A,B\subset\mathcal{Z}$ and $A\cap B=\emptyset$, representing the expected outcome difference between treatment groups $A$ and $B$.
Expected partial derivative
$\theta_{i}(u)=\mathbb{E}_{\omega}\left[\frac{\partial u}{\partial z}(z_{0},\omega)\right],$ (2.9)
at some point $z_{0}\in\mathcal{Z}$, representing the expected marginal effect at $z_{0}$.
The next two propositions show that these functionals, under certain conditions, can be linear and continuous.
Proposition 1.
Suppose that $\mu(A)>0$ and $\mu(B)>0$. The functional (2.8) is linear and continuous on $\mathcal{M}_{i}$.
Note that Example (2.7) is a special case of Example (2.8), obtained by setting $A=\{z^{(1)}\}$ and $B=\{z^{(0)}\}$. The condition $\mu(A)>0$ and $\mu(B)>0$ is commonly referred to as the positivity or overlap assumption in the causal inference literature. It ensures that both treatment groups are represented in the randomisation distribution, which guarantees that the corresponding estimand is well-defined and estimable. In our case, the positivity assumption plays a crucial role in verifying that the functional is both linear and continuous – properties that are essential for the treatment effect to be well-defined and for the Riesz representation to hold.
It is more subtle to establish the linearity and continuity of the functional (2.9), as it requires additional structure from the model space.
Proposition 2.
[…]
This result implies that, for the functional to satisfy the necessary continuity condition, the model space must be further restricted to $\mathcal{M}_{i}\subset L^{2}(\Omega;H^{1}(\mathcal{Z}))\subset L^{2}$. Within this space, weak derivatives are well-defined and continuous, thereby ensuring the applicability of the Riesz representation theorem.
Theorem 1 (Riesz Representation Theorem in $\mathcal{M}_{i}$).
Under Assumptions 1, 3, and 4, for every $\theta_{i}\in\mathcal{M}_{i}^{\prime}$, there exists a unique element $\psi_{i}\in\mathcal{M}_{i}$ such that
$\theta_{i}(u)=\langle u,\psi_{i}\rangle\quad\text{for all }u\in\mathcal{M}_{i},$ (2.10)
and
$\|\theta_{i}\|_{\mathcal{M}_{i}^{\prime}}=\|\psi_{i}\|,$ (2.11)
where the dual norm is defined by
$\left\lVert\theta_{i}\right\rVert_{\mathcal{M}_{i}^{\prime}}:=\sup_{\|u\|\leq 1}\left\lvert\theta_{i}(u)\right\rvert.$ (2.12)
The resulting Riesz representer $\psi_{i}$ plays a central role in our estimation framework.
This approach retains the mathematical elegance of Riesz-based identification while extending it to accommodate outcome-level stochasticity. It lays the foundation for valid estimation in experiments where heterogeneity arises not only from treatment variation but also from latent random effects.
By constructing the Riesz representer $\psi_{i}$ corresponding to the treatment effect functional $\theta_{i}$, we only consider those treatment effects that are representable in the dual space $\mathcal{M}_{i}^{\prime}$. Consequently, we estimate the treatment effect
$\theta_{i}(\tilde{y}_{i})=\langle\tilde{y}_{i},\psi_{i}\rangle=\mathbb{E}\Bigl{[}\tilde{y}_{i}(z,\omega)\,\psi_{i}(z,\omega)\Bigr{]},$ (2.13)
by using the corresponding Riesz estimator
$\hat{\theta}_{i}(z,\omega)=\tilde{y}_{i}(z,\omega)\,\psi_{i}(z,\omega),$ (2.14)
as, in practice, we observe only one realisation of the outcome $\tilde{y}_{i}(z,\omega)$ for certain $z$ and $\omega$. It follows immediately that the Riesz estimator in (2.14) for each unit is unbiased for its treatment effect given in (2.13).
Definition 1 (Aggregate Estimand).
The finite-sample estimand we consider in this paper is defined as a weighted average of individual treatment effects
$\tau_{n}=\sum_{i=1}^{n}\nu_{ni}\theta_{i}(\tilde{y}_{i}),$ (2.15)
where the weights $\nu_{ni}\in\mathbb{R}$ satisfy $\nu_{ni}\geq 0$.
This quantity corresponds to an average treatment effect under a general weighting scheme, aligning with the broader framework of causal inference, where estimands are typically defined as weighted averages of unit-level causal effects across different designs and sampling schemes. Our setup allows for general fixed weighting schemes $\nu_{i}$, providing greater flexibility in defining estimands.
More importantly, $\tau_{n}$ reflects an integration over latent randomness $\omega$, and thus captures variation across a distribution of potential outcomes, effectively averaging over “parallel worlds” indexed by $\omega$. In contrast to the classical Neyman-Rubin framework, where potential outcomes are fixed and comparisons are defined across units within a single universe, our formulation treats outcomes as random functions and estimands as expectations over both treatment assignments and latent heterogeneity. This shift from fixed to random potential outcomes leads to a richer class of estimands, grounded in the observed design but expressive enough to reflect underlying stochastic complexity. It also conceptually aligns with super-population inference while maintaining the rigour of design-based identification.
The estimand $\tau_{n}$ does not rely on Assumption 2; it is mathematically well-defined regardless of whether $z$ and $\omega$ are independent. However, the dependence structure between $z$ and $\omega$ influences the potential outcomes $\tilde{y}_{i}$ and thus affects the value of the estimand. In the causal inference literature, selecting the right estimand involves specifying a target quantity that captures the scientific question of interest under the assumed data-generating process. Different assumptions about the relationship between $z$ and $\omega$ lead to different estimands, each carrying its own causal or associational interpretation.
The estimand considered by HWS corresponds to the special case of our formulation,
$\tau_{n}^{0}=\frac{1}{n}\sum_{i=1}^{n}\langle\tilde{y}_{i},\psi_{i}\rangle^{0}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{z}\Bigl{[}\tilde{y}_{i}(z,\omega_{0})\,\psi_{i}(z,\omega_{0})\Bigr{]},$ (2.16)
where $\Omega=\{\omega_{0}\}$ is a singleton probability space. In this special case, the treatment effect is deterministic and the average is equally weighted.
A common choice for the weights is the uniform weighting $\nu_{ni}=1/n$, which yields the sample average treatment effect (SATE). More generally, this formulation accommodates subgroup-specific or covariate-adjusted averages, depending on the structure of the weights. We impose the following assumption to control the magnitude of the weights $\nu_{ni}$.
Assumption 5 (Uniform Weight Upper Bound).
There exist a constant $\bar{\nu}>0$ such that
$n\nu_{ni}\leq\bar{\nu}\quad\text{for all }i\text{ and all }n\in\mathbb{N}.$ (2.17)
It guarantees that each weight is bounded by a constant multiple of $1/n$, uniformly in $i$, without requiring the individual sequences $\{n\nu_{ni}\}_{n\geq i}$ to converge. Consequently, no finite subset of indices can receive an asymptotically dominant share of the total weight. Note that Assumption 5 implies $\sup_{n\to\infty}\sum_{i=1}^{n}\nu_{ni}<\infty$ but not necessarily one, which is quite general. In the continuous analogue this corresponds to the integrability of the weight function.
Given the unit-level Riesz estimators (2.14) and the weights,
Definition 2 (Aggregate Riesz Estimator).
we define the corresponding aggregate Riesz estimator for the finite-sample estimand
$\hat{\tau}_{n}(z,\omega)=\sum_{i=1}^{n}\nu_{ni}\hat{\theta}_{i}(z,\omega).$ (2.18)
Thus we have the unbiasedness of the Riesz estimator in the following.
Theorem 2 (Unbiasedness of the Riesz Estimator).
The Riesz estimator defined in (2.14) is unbiased for the individual treatment effect, i.e., $\mathbb{E}\bigl{[}\hat{\theta}_{i}(z,\omega)\bigr{]}=\theta_{i}(\tilde{y}_{i})$. Consequently, the aggregate Riesz estimator in (2.18) is unbiased for the finite-sample estimand (2.15), i.e., $\mathbb{E}\Bigl{[}\hat{\tau}_{n}(z,\omega)\Bigr{]}=\tau_{n}$.
This is an immediate consequence of the fact that a single realisation of a random variable is, by definition, unbiased for its expectation.
3 Large Sample Properties of the Aggregate Riesz Estimator
In this section, we clarify the notion of local independence required in our framework, which is a concept that has been widely adopted in recent literature on causal inference under interference. Informally, local dependence refers to the idea that subsets of random variables are conditionally independent of those outside their respective neighbourhoods. This structure is often represented using a dependency graph, in which vertices correspond to units and edges encode potential statistical dependence. For further information about the dependence structure of the variables, see for example Chen and Shao (2004).
Definition 3 (Dependency Neighbourhoods).
For each unit $i$, let $\{M_{ik}\}_{k\in I_{i}}$ be the collection of all subsets $M_{ik}\subset\{1,2,\dots,n\}$ such that the pair of random variables $(\tilde{y}_{i},\psi_{i})$ is independent of the set $\left\{(\tilde{y}_{j},\psi_{j}):j\notin M_{ik}\right\}$. We define the dependency neighbourhood of unit $i$ as the intersection over all such sets:
$N_{i}:=\bigcap_{k\in I_{i}}M_{ik}.$ (3.1)
This definition ensures that $N_{i}$ is the smallest subset of units for which independence holds. Formally, for all such $j\notin N_{i}$, and for all measurable sets $A_{i}\in\mathcal{F}_{\tilde{y}_{i}}$, $B_{i}\in\mathcal{F}_{\psi_{i}}$, $A_{j}\in\mathcal{F}_{\tilde{y}_{j}}$, and $B_{j}\in\mathcal{F}_{\psi_{j}}$, we have
$\mathbb{P}\left((A_{i}\cap B_{i})\cap(A_{j}\cap B_{j})\right)=\mathbb{P}(A_{i}\cap B_{i})\,\,\mathbb{P}(A_{j}\cap B_{j}).$
This definition corresponds to Definition 3.1 in Ross (2011) and is a weaker version of Definition 5.1 in HWS, as it only requires independence of the potential outcome $\tilde{y}_{i}$ and the Riesz representer $\psi_{i}$ from those outside $N_{i}$, rather than full functional independence across the spaces $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$.
In contrast to HWS, who assume independence of the full model spaces across units, our framework only requires independence between the observable random pairs $(\tilde{y}_{i},\psi_{i})$. This relaxation is justified by the introduction of a latent random variable $\omega$, under which potential outcomes are modelled as conditionally random functions. Importantly, our formulation permits cases where the model spaces coincide, i.e. $\mathcal{M}_{i}=\mathcal{M}_{j}$, so long as the associated outcome representer pairs remain independent. Consequently, we do not assume functional independence at the space level, but merely require that the realised outcomes and their associated Riesz representers be independent across sufficiently separated units. This distinction both broadens the scope of our analysis and clarifies the assumptions necessary for variance control and asymptotic normality in a design-based setting.
We denote by $\left\lvert N_{i}\right\rvert$ the number of units in the dependency neighbourhood of unit $i$, which we refer to as the neighbourhood size for unit $i$. Let $D_{n}=\max_{1\leq i\leq n}\left\lvert N_{i}\right\rvert$ and $d_{n}=n^{-1}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert$ denote, respectively, the maximum and average neighbourhood sizes across the sample. These quantities will play a central role in the asymptotic analysis to follow.
In practice, the specification of neighbourhoods $N_{i}$ depends on substantive knowledge of the experimental context. In network experiments, for example, $N_{i}$ might include unit $i$ and its observed social or physical connections. In spatial designs, it is common to define $N_{i}$ via fixed-radius rules or $k$-nearest-neighbour relationships based on geographic distance. When no such structure is explicitly available, one can conservatively define $N_{i}$ to include all units within the same covariate stratum or treatment block as unit $i$. These strategies allow researchers to encode plausible channels of dependence without imposing strong model-based assumptions.
The following two assumptions restrict the growth rate of the dependency neighbourhoods.
Assumption 6 (Maximum Dependency Neighbourhood Size).
The maximum dependency neighbourhood size $D_{n}$ satisfies $D_{n}=o(n^{d})$ for some $d>0$.
Assumption 7 (Average Dependency Neighbourhood Size).
The average dependency neighbourhood size $d_{n}$ satisfies
  1. (a)
    $d_{n}=o(n)$;
  2. (b)
    $\sup_{n\in\mathbb{N}}d_{n}<\infty$.
Assumptions 6 and 7(a) regulate the growth of dependency neighbourhoods and play a role analogous to sparsity constraints commonly used in the analysis of dependent data. Assumption 7(b) additionally imposes uniform boundedness of the average neighbourhood size. These conditions are standard in the local dependence literature and are stated here explicitly to clarify their role and facilitate reference in our self-contained derivations.
In addition, for a sequence of functions $u_{1},\dots,u_{n}\in L^{p}$, where $p>1$, we define the max-$p$ norm as
$\left\lVert u\right\rVert_{\max,p}^{n}:=\max_{1\leq i\leq n}\left\lVert u_{i}\right\rVert_{p}.$ (3.2)
The sequence $\left\lVert u\right\rVert_{\max,p}^{n}$ is clearly non-decreasing in $n$, and, as such, is bounded if and only if it converges as $n\to\infty$.
We now state a proposition that provides an upper bound on the asymptotic variance, denoted by $\sigma_{n}^{2}:=\mathbb{V}[\hat{\tau}_{n}(z,\omega)]$, of the aggregate Riesz estimator $\hat{\tau}_{n}(z,\omega)$ of the estimand $\tau_{n}$. This result follows from the setup and notation introduced above, and it serves to justify or at least suggest the consistency of the estimator under appropriate conditions.
Proposition 3 (Variance Upper Bound).
Under Assumptions 1 and 35, the following inequality holds
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\,\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2},$ (3.3)
for any $p,q\in[1,\infty]$ satisfying $21/p+1/q=1/$ 2. The inequality holds trivially if either $\|\tilde{y}_{i}\|_{p}$ or $\|\psi_{i}\|_{q}$ has no finite upper bound.
Furthermore, we define $S_{2}:=\left\{(p,q)\in[1,\infty]^{2}:\,\frac{1}{p}+\frac{1}{q}=\frac{1}{2}\right\}$, and then we have
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}d_{n}}{n}\left(\inf_{(p,q)\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$ (3.4)
The inequality hold trivially if $\inf_{p,q\in S}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}$ has no finite upper bound.
Proposition 3 provides an upper bound on the variance of the aggregate estimator. It shows that the variance decreases with sample size under boundedness of unit-level norms and local sparsity. The result highlights the interaction between dependence complexity and unit-level variability in determining inferential precision.
It implies that the variance may decay as the sample size increases, provided that the norms $\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}$ and $\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}$ remain bounded if there exists such a pair $p,q\in[1,\infty]$ satisfying $21/p+1/q=1/$ 2, and that the neighbourhood sizes $\left\lvert N_{i}\right\rvert$ are uniformly small. The bound highlights the importance of controlling both the complexity of dependence and the magnitude of unit-level variation in order to obtain valid inference.
Assumption 8 (Boundedness of Norms).
For some $r\geq 1$, there exist real numbers $p,q\in[r,\infty]$ satisfying $1/p+1/q=1/r$ such that
$\sup_{i\in\mathbb{N}}\left\lVert\tilde{y}_{i}\right\rVert_{p}<\infty\quad\text{and}\quad\sup_{i\in\mathbb{N}}\left\lVert\psi_{i}\right\rVert_{q}<\infty.$ (3.5)
This assumption is formulated to allow flexibility in the choice of $r$. Intuitively, the condition imposes uniform boundedness on the $p$th moments of $\tilde{y}_{i}$ and the $q$th moments of $\psi_{i}$, thereby excluding heavy-tailed behaviour.
Note that the max-$p$ norm satisfies $\left\lVert u\right\rVert_{max,p}^{n}\leq\sup_{i\in\mathbb{N}}\left\lVert u_{i}\right\rVert_{p}$. Therefore, Assumption 8 is equivalent to requiring that the four quantities $\|\tilde{y}\|_{\max,p_{1}}^{n}$, $\|\psi\|_{\max,q_{1}}^{n}$, $\|\tilde{y}\|_{\max,p_{2}}^{n}$, and $\|\psi\|_{\max,q_{2}}^{n}$ remain bounded and converge as $n\to\infty$.
The following result follows naturally from our variance control framework and the assumptions stated above.
Theorem 3 (Consistency).
Under Assumptions 1, 35, 7$(a)$, and 8(r=2), the aggregate Riesz estimator in (2.18) is consistent in mean square. Moreover, if 7$(b)$ holds, then the aggregate Riesz estimator satisfies $O_{p}(n^{-1/2})$.
We should interpret $O_{p}(n^{-1/2})$ with care. This notation means that the sequence $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ is bounded in probability, but it does not imply convergence in distribution to a single non-degenerate law. In fact, the sequence may vary between different limiting distributions with bounded variances.
This result has an important practical implication. Although $\hat{\tau}_{n}$ depends on a single treatment assignment $z$ and latent realisation $\omega$, Theorem 3 guarantees that, under mild conditions, averaging over units suffices for consistent estimation. As $n$ grows, $\hat{\tau}_{n}$ approximates the estimand involving expectations over $z$ and $\omega$, reflecting the aforementioned sample-based ergodicity: averaging across units substitutes for repeated sampling.
So far, we have shown that, under certain conditions, the aggregate Riesz estimator is consistent. However, these conditions alone are not sufficient for establishing asymptotic normality; an observation that aligns with the fact that consistency, in general, does not imply asymptotic normality.
The following non-degeneracy assumption imposes a lower bound on the variance of the aggregate Riesz estimator and plays a critical role in ensuring asymptotic normality. When combined with the previous assumptions, it provides sufficient conditions for the aggregate Riesz estimator to satisfy a central limit theorem.
Assumption 9 (Variance Lower Bound).
There exists $\sigma_{0}>0$ such that $\inf_{n\in\mathbb{N}}\sqrt{n}\sigma_{n}\geq\sigma_{0}$.
This condition imposes a non-degeneracy requirement on the estimator. It permits the asymptotic variance $\sigma_{n}^{2}$ to decay with $n$, but not too rapidly, thus ensuring that the root-$n$ scaled estimator $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ retains a non-degenerate variance in the limit.
We require further the following assumption to establish the asymptotic normality result, which concerns the existence of finite fourth moments.
Assumption 10 (Finite Fourth Moment).
The individual treatment effect estimator satisfies:
  1. (a)
    $\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)^{4}\right]<\infty$;
  2. (b)
    $\sup_{i\in\mathbb{N}}\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)^{4}\right]<\infty$.
Note that condition $(b)$ imposes a stronger requirement than $(a)$, as it demands uniform boundedness of the fourth moments across all units.
We now present the theorem establishing asymptotic normality of the Riesz estimator in the stochastic setting.
Theorem 4 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, and 10$(a)$, the aggregate Riesz estimator defined in (2.18) satisfies the following asymptotic normality
[…]
Both Assumptions 8(r=3) and 8(r=4) are required here, as boundedness at $r=4$ does not imply the same at $r=3$ over $(\mathcal{Z},\Omega)$, due to the potential unboundedness of the joint space.
Theorem 4 provides the basis for statistical inference using the aggregate Riesz estimator under outcome-level randomness. In particular, it justifies asymptotic testing and confidence interval construction for the null hypothesis $H_{0}:\tau_{n}=0$, based on the standardised statistic $T_{n}=\hat{\tau}_{n}(z,\omega)/\sigma_{n}$, assuming that both $\hat{\tau}_{n}(z,\omega)$ and $\sigma_{n}$ are consistently estimable from a single realisation of $z$ and $\omega$.
The required rate condition on the maximum dependency neighbourhood size, $D_{n}=o(n^{1/4})$, is in line with that of HWS. However, the proof here proceeds by bounding Wasserstein distance under mixed expectations over both the treatment assignments $z$ and latent heterogeneity $\omega$. This highlights a key observation: the dependence-degree requirement does not tighten despite the added randomness in potential outcomes.
This result is particularly relevant in empirical settings where potential outcomes are inherently random. For instance, in field experiments involving sensor-based measurements, observed outcomes may be contaminated by random noise due to hardware limitations or environmental fluctuations. Similarly, in panel experiments with repeated units, treatment effects may be estimated via machine learning algorithms that introduce randomness in the form of first-stage fitted values. In both cases, the random component of the outcome arises naturally and cannot be ignored. The asymptotic normality result ensures that inference remains valid under such randomness.
4 Variance Estimation under Local Dependence
The test statistic supported by Theorem 4 requires an estimate of the variance of the aggregate Riesz estimator $\hat{\tau}_{n}$, which is generally unknown in finite samples. In the presence of local dependence, such as that arising from network interference, exposure mappings, or latent factors, standard i.i.d. variance estimators are invalid, and specialised methods must be employed.
Several approaches have been proposed. Some yield consistent estimators, but only under strong structural assumptions. For example, Aronow and Samii (2017) rely on known exposure mappings and joint inclusion probabilities; Yu et al. (2022) assume bounded-degree network structures; and Liu and Hudgens (2014) consider two-stage designs with bounded group sizes. Other methods construct variance bounds rather than consistent estimators: Chin (2019) derives conservative expressions using Stein’s method, while Chen and Shao (2004) propose block bootstrap procedures that are valid under fixed neighbourhood size and finite moment assumptions.
Of particular relevance to our setting, HWS construct an unbiased estimator for an upper bound on the variance using a tensor-product Riesz representation. More recently, Harshaw et al. (2024) develop a constrained optimisation framework to tighten such conservative bounds across a wide range of experimental designs. However, neither method formally proves that their bound converges to the true variance, except in special cases. Without sharpness, these bounds may overestimate the variance and lead to overly conservative inference.
In this paper, we take a different approach. Rather than estimating a variance upper bound, we propose a direct estimator for the true asymptotic variance of the aggregate Riesz estimator under local dependence. This estimator is shown to be consistent under conditions that essentially match those required for asymptotic normality. The key requirement is knowledge of the dependency neighbourhood structure: not only the maximal neighbourhood size $D_{n}$, but also which specific units are likely to be dependent. This may seem restrictive, but it is consistent with the broader literature, where the variance is generally considered non-identifiable from a single realisation without further assumptions.
Our recommendation is practical: given $\zeta_{i}=\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$, we define the empirical variance estimator by retaining only the terms that correspond to known (or assumed) dependent pairs. The rest are set to zero. This enables consistent estimation of the variance and allows us to construct feasible and asymptotically valid test statistics.
Let $\vec{\zeta}_{n}=(\zeta_{1},\dots,\zeta_{n})^{\prime}$ denote a vector of centred random variables with covariance matrix $\Sigma_{n}=\mathbb{E}[\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}]$. From a single realisation, we observe only the outer product $\hat{\Sigma}_{n}=\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}$, and cannot estimate $\Sigma_{n}$ consistently without further structure. Even under independence, each $\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}$ typically has variance of order one, so
$\mathbb{E}\bigl{[}\|\hat{\Sigma}_{n}-\Sigma_{n}\|_{F}^{2}\bigr{]}\;\geq\;cn^{2}$ (4.1)
for some constant $c>0$, where $\left\lVert\cdot\right\rVert_{F}$ stands for the Frobenius norm for a matrix, indicating that the Frobenius error diverges with $n$.
Consistency can be recovered by imposing stronger assumptions, for example: observing multiple independent copies of $\vec{\zeta}_{n}$, enabling standard sample covariance estimators; time ordering and weak stationarity allowing the use of banded estimators or kernel-smoothed autocovariances; structural models such as factor structures (low-rank plus sparse decomposition), enabling regularised estimation (e.g., POET); vanishing signal assuming that $\mathbb{V}[\zeta_{n}]\to 0$ as $n\to\infty$, so that the norm $\|\Sigma_{n}\|_{F}$ itself converges to zero; among others.
In the absence of such assumptions, the full matrix $\Sigma_{n}$ is not consistently estimable. Our proposed approach sidesteps this by targeting only those components of the variance that correspond to known dependent pairs.
Suppose the dependency structure is given by
$\mathcal{E}_{n}=\{(i,j)\in\{1,\dots,n\}^{2}:\zeta_{i}\text{ and }\zeta_{j}\text{ are dependent}\}.$ (4.2)
We define the local-dependence variance estimator
$\hat{\sigma}_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j},$ (4.3)
and the population variance
$\sigma_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\mathbb{E}[\zeta_{i}\zeta_{j}].$ (4.4)
The following theorem shows that, under mild regularity conditions, $\hat{\sigma}_{n}^{2}$ consistently estimates $\sigma_{n}^{2}$.
Theorem 5 (Consistency of the Variance Estimator).
Under Assumptions 1, 35, and 10$(b)$, $n\hat{\sigma}_{n}^{2}-n\sigma_{n}^{2}=O_{p}(n^{-1/2}D_{n}^{3/2})$. In particular, under Assumption 6(d=1/3), $n\hat{\sigma}_{n}^{2}-n\sigma_{n}^{2}$ converges to zero in mean square .
Corollary 1 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, and 10$(b)$, the aggregate Riesz estimator defined in (2.18) satisfies the following asymptotic normality
[…]
Note that Assumption 6(d=1/4) is stronger than Assumption 6(d=1/3) and therefore suffices to imply both consistency and asymptotic normality.
Corollary 1 establishes that the aggregate Riesz estimator is asymptotically normal under local dependence, provided that the dependency neighbourhoods are sufficiently sparse and the variance estimator $\hat{\sigma}_{n}^{2}$ is properly constructed.
A natural question is whether one can use a reduced dependency structure based solely on correlation. In many experimental settings, it is difficult to explicitly define or justify a full dependency neighbourhood structure $\mathcal{E}_{n}$ based on latent interference or unobserved design constraints. However, domain knowledge or structural assumptions may suggest which unit-level estimators $\zeta_{i}$ and $\zeta_{j}$ are likely to be uncorrelated, even if their full dependence is unknown. This motivates constructing a variance estimator by summing only over pairs believed to exhibit non-negligible second-order dependence, resulting in a conservative correlation-based estimator. Specifically, consider the set
$\mathcal{E}_{n}^{c}=\{(i,j)\in\{1,\dots,n\}^{2}:\mathbb{E}[\zeta_{i}\zeta_{j}]\neq 0\},$ (4.6)
and define the corresponding correlation-based variance estimator
$\hat{\sigma}_{cn}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}^{c}}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}.$ (4.7)
The answer is affirmative: using this reduced set still yields a consistent variance estimator under the same local dependence framework.
Theorem 6 (Consistency of the Variance Estimator).
Under Assumptions 1, 35, and 10$(b)$, $n\hat{\sigma}_{cn}^{2}-n\sigma_{n}^{2}=O_{p}(n^{-1/2}D_{n}^{3/2})$. In particular, under Assumption 6(d=1/3), $n\hat{\sigma}_{cn}^{2}-n\sigma_{n}^{2}$ converges to zero in mean square .
Corollary 2 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, and 10$(b)$, the aggregate Riesz estimator defined in (2.18) satisfies
[…]
Although the correlation-based variance estimator omits many cross-terms in the full variance formula, it yields valid inference when the omitted pairs correspond to approximately uncorrelated units. Importantly, the assumption that certain unit-level estimators $\zeta_{i}$ and $\zeta_{j}$ are uncorrelated cannot be directly verified from a single dataset. Nonetheless, this assumption often aligns with realistic structural knowledge available to practitioners, such as spatial adjacency, blocking in experimental design, or known network structure, which can be used to construct a conservative approximation of the correlation graph. In this way, the correlation-based approach offers a practically viable and interpretable alternative to full dependency modelling.
Although these variance estimators are developed within our stochastic setting, they are equally applicable to the fixed-outcome setting as a special case. Indeed, the classical framework can be recovered by taking the latent space $\Omega$ to be a singleton, in which case the only source of randomness is the treatment assignment. Under this simplification, the local dependence structure reduces to that of the assignment mechanism alone, typically known by design (see Assumption 2). Consequently, the variance estimators proposed here can be directly applied to previous frameworks such as HWS, and are in fact easier to implement in such settings due to the absence of outcome-level variation. This extends the practical utility of our estimators beyond the stochastic setting, providing consistent inference tools for a broad class of design-based estimators.
In the remainder of the paper, we adopt the correlation-based variance estimator $\hat{\sigma}_{cn}^{2}$, while emphasising that the overall methodology still relies on the local dependency assumption, even though it is not explicitly invoked in the variance formula. Crucially, while uncorrelatedness justifies omitting second-order cross terms in the estimator, it does not imply the vanishing of higher-order joint moments across units. Therefore, a corresponding correlation-based neighbourhood structure cannot be meaningfully defined for the full dependency graph.
In practice, this means we continue to assume a local dependency structure satisfying Assumption 6 with $d=1/4$, while approximating $\sigma_{n}^{2}$ by summing only over pairs $(i,j)$ for which $\zeta_{i}$ and $\zeta_{j}$ are correlated.
Suppose that, in practice, the correlation structure is approximated by a set of index pairs ~$\tilde{\mathcal{E}}_{n}\subset\{1,\dots,n\}^{2}$, serving as a practical substitute for (4.6). To ensure valid inference, the construction of ~$\tilde{\mathcal{E}}_{n}$ should be conservative, satisfying
Assumption 11 (Conservative Set of Index Pairs).
$\mathcal{E}_{n}^{c}\subset\tilde{\mathcal{E}}_{n}\subset\mathcal{E}_{n}$.
Based on this, we define a sparse matrix $\hat{\Sigma}_{n}^{d}\in\mathbb{R}^{n\times n}$ that retains only the entries corresponding to the index set ~$\tilde{\mathcal{E}}_{n}$, representing the estimated dependency structure.
$(\hat{\Sigma}_{n}^{d})_{ij}\;=\;\begin{cases}\zeta_{i}\zeta_{j}&\text{if }(i,j)\in\tilde{\mathcal{E}}_{n},\\ 0&\text{otherwise}.\end{cases}$ (4.9)
The corresponding variance estimator is then given by
$\tilde{\sigma}_{n}^{2}\;=\;\nu_{n}^{\prime}\hat{\Sigma}_{n}^{d}\nu_{n},$ (4.10)
where $\nu_{n}=(\nu_{n1},\dots,\nu_{nn})^{\prime}$ is the vector of aggregation weights.
By appropriately selecting the conservative set of index pairs, the corresponding variance estimator remains consistent.
Theorem 7 (Consistency of the Variance Estimator).
Under Assumptions 1, 35, 10$(b)$, and 11, $n\tilde{\sigma}_{n}^{2}-n\sigma_{n}^{2}=O_{p}(n^{-1/2}D_{n}^{3/2})$. In particular, under Assumption 6(d=1/3), $n\tilde{\sigma}_{n}^{2}-n\sigma_{n}^{2}$ converges to zero in mean square .
Corollary 3 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, 10$(b)$, and 11, the aggregate Riesz estimator defined in (2.18) satisfies
[…]
Note that consistency is preserved even if some independent pairs are mistakenly included in ~$\tilde{\mathcal{E}}_{n}$, provided that the total number of dependent pairs, including both genuinely dependent and erroneously included independent pairs, remains within the sparsity condition $D_{n}=o(n^{1/4})$.
In the special case where all units are believed to be mutually independent, i.e., each $\zeta_{i}$ is independent of every other $\zeta_{j}$, ~$\tilde{\mathcal{E}}_{n}$ consists only of diagonal pairs $(i,i)$. Then $\hat{\Sigma}_{n}^{d}$ reduces to a diagonal matrix, and the variance estimator and the true variances simplify to
$\tilde{\sigma}_{n}^{2}=\sum_{i=1}^{n}\nu_{ni}^{2}\zeta_{i}^{2},\quad\text{and}\quad\sigma_{n}^{2}=\sum_{i=1}^{n}\nu_{ni}^{2}\mathbb{E}[\zeta_{i}^{2}],$ (4.12)
respectively. This corresponds to the classical form used in standard inference under independence. Corollaries 1 and 2 thus recover the familiar asymptotic normality result as a special case of the more general locally dependent setting treated here.
From a functional perspective, the target variance $\sigma_{n}^{2}$ can be regarded as a continuous linear functional of a collection of second-order moments, specifically, $\Sigma_{n}$ or the covariances $\operatorname{Cov}(\zeta_{i},\zeta_{j})$ for pairs $(i,j)$, with norm $\sum_{i=1}^{n}\nu_{ni}^{2}$. Since these covariances are not directly observable from a single experimental realisation, our estimator proceeds by substituting the population moments with observed products $\zeta_{i}\zeta_{j}$, yielding the expression in (4.3). In this sense, the estimator operates analogously to a plug-in estimator: it replaces the unobservable moments with empirical counterparts derived from the data. However, unlike classical plug-in estimators that substitute a distributional parameter with an empirical measure, see for example van der Vaart (2000, Section 25.8), this construction relies on direct moment substitution guided by a known dependency graph. This perspective clarifies that the variance estimators approximates a functional over unknown moments, and its consistency is guaranteed by structural assumptions on the dependence neighbourhood and boundedness of higher-order terms.
Taken together, Theorems 37 and Corollaries 13 convey a central message: although the estimand $\tau_{n}$ is defined as an average of individual treatment effects evaluated over the joint probability space $\mathcal{Z}\times\Omega$, the entire inference procedure, from estimation to variance approximation, can be carried out using just a single realisation of $z\in\mathcal{Z}$ and $\omega\in\Omega$. This is made possible by the structure of the Riesz representation framework and the use of carefully constructed estimators that remain valid under local dependence. As the sample size $n$ grows, one draw from the treatment and outcome distribution contains sufficient information for reliable estimation and inference, provided that the underlying conditions hold. In this way, the results provide a rigorous justification for making population-level claims based on one observed experiment.
5 Riesz Representer Estimation in Practice
The theoretical results in earlier sections show that the aggregate Riesz estimator in the stochastic setting can be constructed using only a single realisation of the underlying data-generating process. Specifically, the Riesz estimates are computed over the randomisation distribution of $z$, conditional on a fixed (but unobserved) $\omega_{0}$. Crucially, this suffices to identify the estimand, which is defined as an expectation over both $z$ and $\omega$.
This structure has concrete implications for how we estimate the individual Riesz representers $\psi_{i}(z,\omega)$. By Theorem 1, for each unit $i$, there exists a unique representer satisfying (2.10) with (2.6). Hence, the representer $\psi_{i}(z,\omega_{0})$ is identified conditionally and can be estimated directly from the realised function $\tilde{y}_{i}(z,\omega_{0})$.
The setting considered by HWS arises as a special case of our framework when $\Omega=\{\omega_{0}\}$, so that potential outcomes reduce to deterministic functions of the treatment assignment.
5.1 Computation in Finite-Dimensional Model Spaces
In finite-dimensional model spaces $\mathcal{M}_{i}$, we can estimate $\psi_{i}(z,\omega_{0})$ via basis expansion and moment matching. Let $\{g_{i,1},\dots,g_{i,m}\}$ be a basis for $\mathcal{M}_{i}$ and $\mathrm{dim}(\mathcal{M}_{i})=m$. We solve for coefficients $\bm{\beta}_{i}\in\mathbb{R}^{m}$ such that
$\hat{\psi}_{i}(z,\omega_{0})=\sum_{k=1}^{m}\beta_{i,k}\,g_{i,k}(z,\omega_{0})$ (5.1)
satisfies the identity $\theta_{i}(\tilde{y}_{i})=\langle\tilde{y}_{i},\hat{\psi}_{i}\rangle^{0}$. Algorithm 1 summarises this procedure. Note that it is in line with the approach described by HWS (preprint version), where Riesz representers are computed via basis expansion and moment matching under the randomisation distribution. we provide here a few concrete examples of basis choices to clarify practical implementation.
When the treatment space $\mathcal{Z}$ is finite with $K$ distinct levels, a simple and interpretable basis consists of treatment indicators. In this case, one may define each basis function $g_{i,k}(z,\omega)$ to equal 1 if $z$ corresponds to the $k$-th treatment condition and 0 otherwise. This yields a saturated model with a separate coefficient for each treatment arm and does not rely on any structural assumptions about the response surface.
When $z$ is continuous or high-dimensional, one can instead use low-order basis functions such as polynomials or splines in $z$, possibly interacted with pre-specified functions of $\omega$. For instance, if $z\in\mathbb{R}$, a natural choice is the span of $\{1,z,z^{2},\omega,z\cdot\omega\}$, which captures basic nonlinearities and interactions between treatment and latent outcome variation. These basis sets offer a balance between expressiveness and computational tractability and can be tailored to the structure of the experiment.
This procedure is fully offline in the sense that it depends only on the design distribution for $z$, the basis functions, and the known functional $\theta_{i}$. It does not require access to observed outcomes $Y_{i}=\tilde{y}_{i}(z_{0},\omega_{0})$.
This construction relies on Assumption 2, which ensures that the treatment assignment $z$ is independent of the latent variable $\omega$. In particular, this allows expectations over $z$ conditional on a fixed realisation $\omega_{0}$ to be taken with respect to the known design distribution $\mu$. Without this assumption, the conditional distribution of $z\mid\omega_{0}$ could differ from $\mu$, and the estimated representer would no longer correspond to the true Riesz representation defined under the design. It is at this stage, during the offline construction of $\hat{\psi}_{i}(z,\omega_{0})$, that Assumption 2 is invoked. All subsequent inference proceeds without requiring independence between treatment assignment and outcome-level randomness.
In symmetric settings with common $\mathcal{M}_{i}$ and $\theta_{i}$, the same representer $\psi$ can be reused across units.
5.2 Approximation in Infinite-Dimensional Model Spaces
In many practical applications, the model space $\mathcal{M}_{i}$ is infinite-dimensional. This arises naturally when either the treatment assignment $z$ or the latent variable $\omega$ lies in a continuous domain, or when potential outcomes depend on complex functional relationships. To compute the Riesz representer in such settings, we project the problem onto a sequence of growing finite-dimensional subspaces.
Our setting explicitly accounts for a random latent variable $\omega$, and the Riesz representation is defined with respect to an inner product $\langle\cdot,\cdot\rangle^{0}$ taken over the treatment assignment $z$, conditional on the fixed value $\omega_{0}\in\Omega$. Although our estimation routine shares a computational structure with existing methods, constructing a Gram matrix and solving a linear system, our framework generalises the representer computation to a richer class of problems involving random potential outcomes.
Assumption 12 (Separable Model Space).
The model space $\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega)$, equipped with the inner product $\langle\cdot,\cdot\rangle^{0}$ defined in (2.6) for some $\omega_{0}\in\Omega$, is a separable Hilbert space.
By Assumption 12, the space $\mathcal{M}_{i}$ admits a countable orthonormal basis $\{e_{k}\}_{k\geq 1}$. For each dimension $m\in\mathbb{N}$, define the subspace
$\mathcal{M}_{i}^{(m)}=\operatorname{span}\{e_{1},\dots,e_{m}\},$ (5.2)
and let $P_{m}:\mathcal{M}_{i}\to\mathcal{M}_{i}^{(m)}$ denote the orthogonal projection. Then the projected representer
$\psi_{i}^{(m)}:=P_{m}\psi_{i}=\sum_{k=1}^{m}\langle\psi_{i},e_{k}\rangle^{0}\,e_{k}$ (5.3)
converges to $\psi_{i}$ in norm as $m\to\infty$, by the completeness of the Hilbert space $\mathcal{M}_{i}$, since $\{e_{k}\}_{k\geq 1}$ forms a total orthonormal system and $P_{m}$ is the orthogonal projection onto the span of the first $m$ basis elements.
The estimation of $\psi_{i}^{(m)}$ proceeds by solving the finite-dimensional moment equation system that results from projecting the Riesz representation identity onto the span of the first $m$ basis functions. This yields a linear system involving a Gram matrix and a target vector, as shown in Algorithm 2.
This procedure is still fully offline in the sense that both $G_{i}^{(m)}$ and $T_{i}^{(m)}$ can be computed offline from the design distribution and the known linear functional $\theta_{i}$. It does not require access to observed outcomes $Y_{i}=\tilde{y}_{i}(z_{0},\omega_{0})$. In symmetric settings with common $\mathcal{M}_{i}$ and $\theta_{i}$, the same representer $\psi$ can be reused across units.
The procedure for estimating $\psi_{i}(z,\omega_{0})$ described above relies on projecting onto a finite-dimensional subspace spanned by a pre-specified basis and solving empirical moment equations derived from the known randomisation design and linear functional $\theta_{i}$. This approach follows the classical sieve estimation framework and is theoretically well understood; see, e.g., Chen (2007); Chen and Pouzo (2012); van der Vaart (2000).
While consistency of $\hat{\psi}_{i}^{(m)}\to\psi_{i}$ under appropriate regularity conditions can be rigorously established using a standard bias-variance decomposition, this is not the focus of the present paper. Our main contribution lies in the theory developed in earlier sections, namely, the design-based asymptotic normality of the aggregate estimator and the construction of consistent variance estimators under local dependence.
For completeness, we note that under standard conditions, such as an orthonormal basis, eigenvalues of the Gram matrix bounded away from zero, and sufficient smoothness of the representer, the estimation error satisfies
$\|\widehat{\psi}_{i}^{(m)}-\psi_{i}\|=O_{p}\left(m^{-s}+\sqrt{m/n}\right)$ (5.4)
for some $s>0$. This rate is minimised by setting the truncation level $m$ proportional to $n^{1/(2s+1)}$, which balances the bias and variance contributions. We omit the full proof, as it follows standard arguments from the sieve estimation literature.
The orthonormal system $\{e_{k}\}$ can either be specified a priori, using classical bases such as Fourier series, wavelets, or splines, or derived from the data through methods like principal components or kernel eigenfunctions. Additional practical examples of basis construction in deterministic settings are discussed in the preprint by HWS, to which we refer the interested reader for further details.
6 Simulation Study
To illustrate the finite-sample behaviour, we carry out a simulation study to evaluate the finite-sample performance of the Riesz estimator and the variance estimators. The simulation is designed to reflect a typical experimental setting with local dependence induced by latent block-level effects. All potential outcomes, covariates, and errors are modelled as functions of a common latent variable, consistent with our framework.
For each unit $i\in\{1,\dots,n\}$, we simulate treatment assignment $z_{i}\sim\mathrm{Bernoulli}(0.5)$ independently. Covariates $x_{i}(\omega)$ and outcome-level parameters $(\alpha_{i},\beta_{i},\delta_{i})$ are drawn i.i.d. from $\mathcal{N}(0,1)$. The potential outcome is generated via
$\tilde{y}_{i}(z,\omega)=\alpha_{i}+\beta_{i}z+\delta_{i}zx_{i}+\varepsilon_{i}(\omega),$ (6.1)
where $\varepsilon_{i}(\omega)$ is a locally dependent noise term described below. The observed outcome is $Y_{i}=\tilde{y}_{i}(z_{i},\omega)$.
Local dependence is introduced by partitioning the sample into blocks, where units in the same block share a common latent shock. Specifically, the number of blocks is $B_{n}=\lfloor n^{1-d}\rfloor$ for a fixed parameter $d\in[0,0.3]$. The error for unit $i$ is
$\varepsilon_{i}(\omega)=\gamma_{i}\eta_{b(i)}+\nu_{i},$
where $b(i)$ is the block to which unit $i$ is assigned, $\eta_{b}\sim\mathcal{N}(0,\sigma^{2})$ is the shared block-level shock (with $\sigma^{2}=1$), $\nu_{i}\sim\mathcal{N}(0,1)$ is an idiosyncratic noise term, and $\gamma_{i}\in\{-1,1\}$ is a random sign (independent Rademacher) to allow both positive and negative correlations between units in the same block. This construction yields a dependency graph $\mathcal{E}_{n}$ with within-block dependence and across-block independence.
We estimate the average treatment effect using the aggregate Riesz estimator. The representer $\psi_{i}$ is computed as
$\psi_{i}(z,\omega_{0})=\frac{z}{\mu_{1}}-\frac{1-z}{\mu_{0}},\quad\text{with }\mu_{1}=\mathbb{E}[z_{i}]=0.5,$
which corresponds to the efficient estimator under the known design. The estimator is
$\hat{\tau}_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}\psi_{i}(z_{i},\omega_{0}),$
and the variance estimator follows (4.3)
$\hat{\sigma}_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j},\quad\text{where }\zeta_{i}=Y_{i}\psi_{i}(z_{i})-\theta_{i}(\tilde{y}_{i})\text{ and }\nu_{ni}=1/n.$
That is, we sum pairwise products of residuals over all unit pairs within the same block.
In this simulation setting, the variance estimators $\hat{\sigma}_{cn}^{2}$ and $\tilde{\sigma}_{n}^{2}$ coincide with the local-dependence estimator $\hat{\sigma}_{n}^{2}$. This is because the dependency structure is fully captured by shared block-level shocks, and all non-zero cross-moments correspond to block-sharing units. Therefore, all simulation results reflect the performance of both estimators. In general, however, the correlation-based estimator yields improved efficiency in settings where weakly dependent pairs can be excluded without loss of accuracy.
We consider sample sizes $n\in\{100,200,500,1000\}$ and dependency growth parameters $d\in\{0.0,0.1,0.2,0.25,0.3\}$. For each $(n,d)$ pair, we run $2000$ independent replications. For each replication, we compute the estimator $\hat{\tau}_{n}$, the standard error $\hat{\sigma}_{n}$, and test statistics for Wald-type inference. We record the following metrics:
  • Empirical bias: $\mathbb{E}[\hat{\tau}_{n}-\tau_{n}]$;
  • Root mean squared error (RMSE): $\sqrt{\mathbb{E}[(\hat{\tau}_{n}-\tau_{n})^{2}]}$;
  • Coverage of 95% confidence intervals: $\mathbb{P}(|\hat{\tau}_{n}-\tau_{n}|\leq 1.96\hat{\sigma}_{n})$;
  • Rejection rates at 1%, 5%, and 10% significance levels.
The complete codebase, including replication scripts and plotting routines, is publicly available at github.com/yukai-yang/RieszRE_Experiments under the MIT license.
Figure 1 depicts the empirical bias and root mean squared error (RMSE) of the Riesz estimator as functions of the dependency growth rate $d$, stratified by sample size. Across all settings, the estimator exhibits negligible bias: empirical values remain close to zero, typically below 0.01 in absolute magnitude. The RMSE decreases steadily as the sample size increases, from approximately 0.37 at $n=100$ to around 0.10 at $n=1000$, indicating improved precision in larger samples. These results offer strong empirical support for the consistency of $\hat{\tau}_{n}$ under local dependence, as established in Theorem 3.
Figure 2 shows the empirical coverage rates of nominal 95% confidence intervals across varying levels of dependency growth $d$ and sample sizes $n$. Coverage rates are uniformly at or above the nominal level in all scenarios. For $n=100$, coverage ranges from 94.9% to 95.6%, while for $n=500$ and $n=1000$, it stabilises in the range of 95.5% to 96.8%. These results indicate that the variance estimator is slightly conservative but well-calibrated, yielding confidence intervals with reliable coverage even in the presence of moderate local dependence.
Table 2 reports the empirical rejection rates of two-sided Wald tests at the 1%, 5%, and 10% significance levels across all combinations of $n$ and $d$. At the 5% level, rejection rates generally lie within the 4%–5% range. For instance, when $n=1000$, the empirical size ranges from 3.4% to 4.5%, depending on the degree of local dependence. Similar patterns hold at the 1% and 10% levels, with empirical rejection rates closely aligned with nominal values across the board. These findings indicate that the standardised test statistic
$T_{n}=\frac{\hat{\tau}_{n}-\tau_{n}}{\hat{\sigma}_{n}}$
is approximately standard normal under the null, consistent with the asymptotic normality established in Corollary 1.
The simulation confirms the unbiasedness, consistency, and inferential validity of the Riesz estimator in finite samples. The variance estimator successfully accounts for local dependence induced by block-level shocks, even when correlations are signed and heterogeneous. The method performs well across all sample sizes and dependency regimes considered, providing empirical support for its practical application.
7 Conclusion
This paper extends the classical design-based framework for causal inference to accommodate random potential outcomes. By introducing latent randomness through a variable $\omega$, we allow for outcome-level variability while preserving the core identification principle of randomised treatment assignment. This formulation retains the key strength of design-based methods, minimal reliance on outcome modelling, while generalising the framework to address the stochastic nature of modern experimental data. In doing so, it contributes toward a broader synthesis between randomised experimental design and inferential generalisation.
We develop a general asymptotic theory for the aggregate Riesz estimator under local dependence in the stochastic setting. A central strength of the results lies in their generality: neither the estimator $\hat{\tau}_{n}(z,\omega)$ nor the variance $\sigma_{n}^{2}$ is required to converge to a fixed distribution or limiting value. Instead, the theory accommodates growing network complexity, heterogeneous variance behaviour, and non-stationary dependence structures.
Our work departs from the fixed-outcome framework by modelling treatment effect functionals in a random outcome space. This yields a distinct theoretical foundation that supports valid inference under local dependence from a single observed dataset. In addition, we develop consistent variance estimators tailored to this setting. The contribution lies not in additional technical abstraction, but in providing a rigorous, design-respecting methodology that accommodates the outcome-level randomness inherent in many modern experiments.
Although our variance estimators are developed within the stochastic setting, they remain valid in the classical fixed-outcome setting as a special case. In particular, when outcome-level randomness is absent, the local dependence structure reduces to that of the assignment mechanism alone. This structure is typically known by design. As a result, our variance estimators can be readily applied in earlier frameworks such as HWS, where they provide a consistent and implementable alternative to the conservative variance bounds previously available.
Theorems 37, along with Corollaries 13, establish a robust framework for inference using the Riesz estimator in the stochastic setting. For instance, Theorem 4 shows that the standardised statistic $\sigma_{n}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)$ is asymptotically standard normal without requiring convergence of the variance sequence. Similarly, Theorems 57 show that the variance estimators $n\hat{\sigma}_{n}^{2}$, $n\hat{\sigma}_{cn}^{2}$, and $\tilde{\sigma}_{n}^{2}$ consistently estimate $n\sigma_{n}^{2}$, even when the latter does not stabilise. These results enable valid inference in experiments with increasing dependence, network size, or covariate complexity.
Theorem 3 establishes mean-square consistency under mild moment and neighbourhood growth assumptions. When the average neighbourhood size is uniformly bounded, the estimator achieves an $O_{p}(n^{-1/2})$ convergence rate without relying on classical limiting distributions. Corollaries 13 provide feasible, fully data-driven test statistics that converge to the standard normal under regularity conditions and consistent variance estimation.
To evaluate finite-sample behaviour, we conduct a simulation study using locally dependent data generated via signed block-level shocks. The results confirm the theoretical properties: the Riesz estimator is unbiased and consistent, with RMSE declining as sample size increases. Confidence intervals based on the variance estimator exhibit accurate or slightly conservative coverage across dependency settings. Empirical rejection rates for two-sided Wald tests closely match nominal levels, even in moderate samples, supporting the method’s practical applicability.
In summary, the Riesz representation approach remains valid and powerful in the presence of outcome-level randomness and structured dependence. Our framework is robust and flexible, enabling valid inference without restrictive asymptotic assumptions. It is particularly well suited to modern experiments involving sensors, adaptive interventions, high-dimensional treatments, or biological systems, where random variation is intrinsic rather than incidental.
While the correlation-based variance estimator $\hat{\sigma}_{cn}^{2}$ offers a more parsimonious alternative to the full local-dependence estimator, it rests on untestable assumptions about which unit-level estimators are uncorrelated. Its finite-sample performance depends on how well the assumed correlation structure approximates the true dependency graph. Nonetheless, simulations suggest that under reasonable constructions, this approach may yield tighter intervals and improved efficiency while maintaining valid coverage.
Future research may extend the framework in several directions. Natural next steps include incorporating non-linear or non-smooth functionals, adapting the methodology to observational studies (e.g., through covariate balancing or doubly robust adjustment), and generalising to high-dimensional or non-parametric spaces. While this paper focuses on latent randomness under local dependence, extensions to dynamic treatments, time-varying interference, or structured measurement error remain open and promising. We believe these tools lay the foundation for a broader class of estimators that remain faithful to randomisation principles while accommodating the stochastic complexity of modern data.
We hope that this framework will serve as a useful addition to the design-based toolkit, especially in experiments where outcome-level randomness cannot be ignored.
Acknowledgements
I am grateful to Fredrik Sävje for his valuable discussions and support.
Appendix A Notation Summary
Table 1: Summary of commonly used notation throughout the paper.
Symbol Description
$n$ Number of units in the experiment
$z\in\mathcal{Z}$ Treatment assignment vector
$\mu$ Probability measure for randomisation
$\omega\in\Omega$ Latent variable representing outcome-level randomness
$P$ Probability measure for the latent random variable $\omega$
$\mathcal{F}_{x}$ Sigma-algebra generated by the random variable $x$
$\tilde{y}_{i}(z,\omega)$ Potential outcome of unit $i$ under treatment $z$ and latent variable $\omega$
$\theta_{i}(\tilde{y}_{i})$ Linear functional defining the treatment effect for unit $i$
$\hat{\theta}_{i}(z,\omega)$ Riesz estimator for the treatment effect for unit $i$
$\zeta_{i}$ the difference $\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$
$\tau_{n}$ Aggregate estimand (e.g., average treatment effect)
$\hat{\tau}_{n}$ Aggregate Riesz estimator for $\tau_{n}$
$\psi_{i}(z,\omega)$ Riesz representer for unit $i$
$\hat{\psi}_{i}(z,\omega)$ Estimated Riesz representer (e.g., via basis expansion)
$\mathcal{M}_{i}$ Model space for unit $i$’s outcome function
$N_{i}$ Dependency neighbourhood of unit $i$
$D_{n}$ Maximum neighbourhood size across units
$d_{n}$ Average neighbourhood size across units
$\sigma_{n}^{2}$ Variance of $\hat{\tau}_{n}$
$\mathcal{E}_{n}$ set of dependent pairs $(i,j)$ for $\zeta_{i}$ and $\zeta_{j}$
$\hat{\sigma}_{n}^{2}$ Variance estimator of $\hat{\tau}_{n}$ using $\mathcal{E}_{n}$
$\mathcal{E}_{n}^{c}$ set of correlated pairs $(i,j)$ for $\zeta_{i}$ and $\zeta_{j}$
$\hat{\sigma}_{cn}^{2}$ Variance estimator of $\hat{\tau}_{n}$ using $\mathcal{E}_{n}^{c}$
~$\tilde{\mathcal{E}}_{n}$ conservative set of index pairs
$\tilde{\sigma}_{n}^{2}$ Variance estimator of $\hat{\tau}_{n}$ using ~$\tilde{\mathcal{E}}_{n}$
Appendix B Algorithms
Algorithm 1 Computing the Riesz representer $\psi_{i}(z,\omega_{0})$, following the procedure of HWS
Inputs:
$\bullet$ Basis functions $\{g_{i,1},\dots,g_{i,m}\}$ for $\mathcal{M}_{i}$
$\bullet$ Known linear functional $\theta_{i}$
$\bullet$ Randomisation distribution for $z$
Step 1: Compute Gram Matrix $\hat{\mathbf{G}}_{i}$
$[\hat{\mathbf{G}}_{i}]_{\ell,k}=\mathbb{E}_{z}\bigl{[}g_{i,\ell}(z,\omega_{0})g_{i,k}(z,\omega_{0})\bigr{]}$
Step 2: Compute Target Vector $\hat{\mathbf{T}}_{i}$
$[\hat{\mathbf{T}}_{i}]_{\ell}=\theta_{i}(g_{i,\ell})$
Step 3: Solve Linear System
$\hat{\mathbf{G}}_{i}\bm{\beta}_{i}=\hat{\mathbf{T}}_{i}\quad\Rightarrow\quad\bm{\beta}_{i}=\hat{\mathbf{G}}_{i}^{-1}\hat{\mathbf{T}}_{i}$
Output: Estimated Riesz representer $\hat{\psi}_{i}(z,\omega_{0})$
Algorithm 2 Approximating the Riesz representer $\psi_{i}(z,\omega_{0})$ in infinite-dimensional settings
Inputs:
$\bullet$ Truncated orthonormal basis $\{e_{1},\dots,e_{m}\}$ of $\mathcal{M}_{i}$
$\bullet$ Known linear functional $\theta_{i}$
$\bullet$ Randomisation distribution for $z$
Step 1: Compute Gram Matrix $\hat{\mathbf{G}}_{i}^{(m)}$
$[\hat{\mathbf{G}}_{i}^{(m)}]_{\ell,k}=\mathbb{E}_{z}\bigl{[}e_{\ell}(z,\omega_{0})\,e_{k}(z,\omega_{0})\bigr{]}$
Step 2: Compute Target Vector $\hat{\mathbf{T}}_{i}^{(m)}$
$[\hat{\mathbf{T}}_{i}^{(m)}]_{\ell}=\theta_{i}(e_{\ell})$
Step 3: Solve Regularised Linear System
$\hat{\mathbf{G}}_{i}^{(m)}\bm{\beta}_{i}=\hat{\mathbf{T}}_{i}^{(m)}\quad\Rightarrow\quad\bm{\beta}_{i}=\left(\hat{\mathbf{G}}_{i}^{(m)}\right)^{-1}\hat{\mathbf{T}}_{i}^{(m)}$
Output: Estimated Riesz representer
$\hat{\psi}_{i}^{(m)}(z,\omega_{0})=\sum_{k=1}^{m}\beta_{i,k}\,e_{k}(z,\omega_{0})$
Appendix C Figures and Table
Figure 1: Bias and RMSE of estimators across varying neighbourhood growth rates $d$.
[…]
Figure 2: Coverage of 95% confidence intervals under varying dependency growth rates $d$.
0.920.940.960.9800.10.20.250.3Dependency growth rate $d$$n=$1002005001000
Table 2: Empirical rejection rates at 1%, 5%, and 10% significance levels across sample sizes $n$ for each dependency growth rate $d$. Based on 2000 replications using the variance estimator.
Sig. levels $d$ $n=100$ $n=200$ $n=500$ $n=1000$
1% 0.00 0.0075 0.0060 0.0060 0.0050
0.10 0.0100 0.0050 0.0050 0.0070
0.20 0.0100 0.0065 0.0045 0.0085
0.25 0.0115 0.0100 0.0075 0.0055
0.30 0.0155 0.0100 0.0100 0.0060
5% 0.00 0.0470 0.0345 0.0325 0.0380
0.10 0.0440 0.0380 0.0410 0.0335
0.20 0.0450 0.0410 0.0390 0.0450
0.25 0.0455 0.0480 0.0425 0.0345
0.30 0.0510 0.0480 0.0415 0.0435
10% 0.00 0.0930 0.0820 0.0785 0.0825
0.10 0.0870 0.0810 0.0830 0.0795
0.20 0.0865 0.0855 0.0855 0.0850
0.25 0.0865 0.0950 0.0930 0.0760
0.30 0.0955 0.0905 0.0940 0.0920
Appendix D Lemmas and Proofs
Proof of Proposition 1.
The linearity for both the two examples follows directly from the facts that integrals are linear and that we are taking differences of linear functionals.
For continuity, we need to rewrite (2.8) clearly
$\theta_{i}(u)=\iint_{\mathcal{Z}\times\Omega}\left(\frac{\mathbf{I}_{A}(z)}{\mu(A)}-\frac{\mathbf{I}_{B}(z)}{\mu(B)}\right)u(z,\omega)\,\mu(\mathrm{d}z)\,P(\mathrm{d}\omega)=\left\langle u,\,\varphi\right\rangle,$
where $\mathbf{I}$ stands for indicator function, and
$\varphi(z,\omega)=\frac{\mathbf{I}_{A}(z)}{\mu(A)}-\frac{\mathbf{I}_{B}(z)}{\mu(B)}.$
We can see that $\varphi\in L^{2}(\mathcal{Z}\times\Omega)$ as
$\|\varphi\|_{L^{2}}^{2}$ $=$ $\int_{\Omega}\int_{\mathcal{Z}}\left|\frac{\mathbf{I}_{A}(z)}{\mu(A)}-\frac{\mathbf{I}_{B}(z)}{\mu(B)}\right|^{2}\,\mu(\mathop{}\!\mathrm{d}z)\,P(\mathop{}\!\mathrm{d}\omega)$
$=$ $\int_{\mathcal{Z}}\left|\frac{\mathbf{I}_{A}(z)}{\mu(A)}-\frac{\mathbf{I}_{B}(z)}{\mu(B)}\right|^{2}\,\mu(\mathrm{d}z)=\frac{1}{\mu(A)}+\frac{1}{\mu(B)}<\infty.$
Now continuity is immediate due to the Cauchy-Schwarz inequality
$|\theta_{i}(u)|=|\langle u,\varphi\rangle|\leq\left\lVert u\right\rVert\,\left\lVert\varphi\right\rVert.$
Proof of Proposition 2.
Derivatives and expectation are linear, which establishes linearity.
Since $u\in L^{2}(\Omega;H^{1}(\mathcal{Z}))$, it follows that for almost every $\omega\in\Omega$, the function $z\mapsto u(z,\omega)\in H^{1}(\mathcal{Z})$, and hence the weak derivative $\partial u/\partial z$ exists and is continuous almost everywhere in $z$. Moreover, for any interior point $z_{0}\in\mathcal{Z}$, the mapping $\omega\mapsto\partial u/\partial z(z_{0},\omega)$ belongs to $L^{2}(\Omega)$, with
$\left\|\frac{\partial u}{\partial z}(z_{0},\cdot)\right\|_{L^{2}(\Omega)}\leq C\left\|\frac{\partial u}{\partial z}\right\|,$
for some constant $C>0$, by the Sobolev embedding theorem and the continuity of the point evaluation map on $H^{1}(\mathcal{Z})$. See Adams and Fournier (2003, Theorem 5.36) and related discussion of pointwise evaluation in $H^{1}$ spaces.
Applying the Cauchy-Schwarz inequality gives
$|\theta_{i}(u)|=\left|\int_{\Omega}\frac{\partial u}{\partial z}(z_{0},\omega)\,P(\mathrm{d}\omega)\right|\leq\left\|\frac{\partial u}{\partial z}(z_{0},\cdot)\right\|_{L^{2}(\Omega)},$
and this completes the proof. ∎
Proof of Theorem 1.
See Riesz (1907). ∎
Proof of Theorem 2.
By definition of the Riesz estimator in (2.14) and the linearity of expectation, we have
$\mathbb{E}\bigl{[}\hat{\theta}_{i}(z,\omega)\bigr{]}=\theta_{i}(\tilde{y}_{i}),$
which further implies
$\mathbb{E}\bigl{[}\hat{\tau}_{n}(z,\omega)\bigr{]}=\sum_{i=1}^{n}\nu_{ni}\,\mathbb{E}\bigl{[}\hat{\theta}_{i}(z,\omega)\bigr{]}=\sum_{i=1}^{n}\nu_{ni}\,\theta_{i}(\tilde{y}_{i})=\tau_{n}.$
This equality follows from the construction of the estimator and does not require any integrability assumptions. In particular, it holds even if $\tilde{y}_{i}(z,\omega)\,\psi_{i}(z,\omega)\notin L^{1}$. ∎
Lemma 1.
Suppose the weights $\nu_{ni}$ satisfy Assumption 5. Then for any integer $r\geq 1$, the following bound holds:
$\sum_{i=1}^{n}\nu_{ni}^{r}=O(n^{1-r}).$ (D.1)
Proof.
By Assumption 5, we have
$\sum_{i=1}^{n}\nu_{ni}^{r}\leq\sum_{i=1}^{n}\left(\frac{\bar{\nu}}{n}\right)^{r}=n\cdot\left(\frac{\bar{\nu}}{n}\right)^{r}=\bar{\nu}^{r}\cdot n^{1-r},$
which implies the claimed order bound. ∎
Lemma 2.
Under Assumptions 1, 3, and 4, the $r$th central moment of the treatment effect Riesz estimator, for some integer $r\geq 1$, if it exists, satisfies
$\mathbb{E}\left|\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right|^{r}\leq\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p}\,\left\lVert\psi_{i}\right\rVert_{q}\right)^{r},$ (D.2)
for any $p,q\in[1,\infty]$ satisfying $1/p+1/q=1/r$. In particular, when $r=2$,
$\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{2}\leq\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2}.$ (D.3)
The inequalities (D.2) and (D.3) hold trivially if either $\|\tilde{y}_{i}\|_{p}$ or $\|\psi_{i}\|_{q}$ has no finite upper bound.
Proof.
Inequality (D.2) is a direct extension of Lemma C.3 in HWS, which handles the case without the additional randomness $\omega$. The proof carries through immediately by applying Hölder’s inequality on the product space $\mathcal{Z}\times\Omega$, since the additional integration over $\omega$ preserves the inequality structure.
Regarding inequality (D.3),
$\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{2}$ $=\mathbb{E}[\hat{\theta}_{i}(z,\omega)]^{2}-(\theta_{i}(\tilde{y}_{i}))^{2}\leq\mathbb{E}\left[\tilde{y}_{i}(z,\omega)\psi_{i}(z,\omega)\right]^{2}$
$=\left\lVert\tilde{y}_{i}(z,\omega)\psi_{i}(z,\omega)\right\rVert^{2}\leq\left(\left\lVert\tilde{y}_{i}(z,\omega)\right\rVert_{p}\,\left\lVert\psi_{i}(z,\omega)\right\rVert_{q}\right)^{2}.$ (Hölder)
Proof of Proposition 3.
Let $\zeta_{i}=\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$ and we observe that $\mathbb{E}[\zeta_{i}]=0$ by Theorem 2. Then,
$\sigma_{n}^{2}$ $=\mathbb{V}[\hat{\tau}_{n}(z,\omega)-\tau_{n}]=\mathbb{V}\left[\sum_{i=1}^{n}\nu_{ni}\zeta_{i}\right]=\mathbb{E}\left[\sum_{i=1}^{n}\nu_{ni}\zeta_{i}\right]^{2}$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\nu_{ni}\nu_{nj}\mathbf{I}(i\in N_{j})\mathbf{I}(j\in N_{i})\mathbb{E}\left[\zeta_{i}\zeta_{j}\right]$
$\leq\sum_{i=1}^{n}\sum_{j=1}^{n}\nu_{ni}\nu_{nj}\mathbf{I}(i\in N_{j})\mathbf{I}(j\in N_{i})\left\lVert\zeta_{i}\right\rVert\,\left\lVert\zeta_{j}\right\rVert$ (Cauchy-Schwartz)
$\leq\sum_{i=1}^{n}\nu_{ni}\sum_{j=1}^{n}\frac{1}{2}\nu_{nj}\left(\mathbf{I}(j\in N_{i})\left\lVert\zeta_{i}\right\rVert^{2}+\mathbf{I}(i\in N_{j})\left\lVert\zeta_{j}\right\rVert^{2}\right)$ (AM-GM)
$=\sum_{i=1}^{n}\nu_{ni}\left(\sum_{j\in N_{i}}\nu_{nj}\right)\left\lVert\zeta_{i}\right\rVert^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\left\lVert\zeta_{i}\right\rVert^{2}$ (Lemma 1 (2.17))
$\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\,\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2}$ (Lemma 2 (D.3))
Note that AM-GM stands for arithmetic-geometric mean inequality.
To obtain the uniform bound, we apply the max-$p$ norm, as defined in (3.2), and note that $d_{n}=n^{-1}\sum_{i=1}^{n}|N_{i}|$, yielding
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\left(\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert=\frac{\bar{\nu}^{2}d_{n}}{n}\left(\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$
Since this inequality holds for any $p,q\in[1,\infty]$ satisfying $21/p+1/q=1/$ 2, we have a more conservative upper bound
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}d_{n}}{n}\left(\inf_{p,q\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$
Proof of Theorem 3.
The result follows directly from Proposition 3. Specifically, the bound in (3.4), together with Assumption 8 with $r=2$, implies that
$\mathbb{E}\left[\hat{\tau}_{n}(z,\omega)-\tau_{n}\right]^{2}\to 0,$
as $n\to\infty$, establishing mean square consistency.
Moreover, suppose $\sup_{n\in\mathbb{N}}d_{n}<\infty$, that is, there exists some $d_{0}\in(0,\infty)$ such that $d_{n}\leq d_{0}$ for all $n$. Since $\hat{\tau}_{n}-\tau_{n}$ follows some distribution with zero mean and variance $\sigma_{n}^{2}$, it follows that $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ has variance $n\sigma_{n}^{2}$. By the bound in Proposition 3, we obtain
$n\sigma_{n}^{2}\leq\bar{\nu}^{2}d_{0}\left(\inf_{p,q\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}<\infty,$
under the stated assumptions. Hence, the aggregate Riesz estimator satisfies $O_{p}(n^{-1/2})$. ∎
Lemma 3.
[…]
Proof.
We will show that $\mathbb{E}[f(W_{n})]\to\mathbb{E}[f(Z)]$ for all bounded uniformly continuous functions $f:X\to\mathbb{R}$, which characterises weak convergence; see Billingsley (1999, Theorem 2.1).
Let $f\in\mathrm{BUC}(X)$, the space of all bounded uniformly continuous functions. Since Lipschitz functions are dense in $\mathrm{BUC}(X)$ under the supremum norm, for any $\varepsilon>0$, there exists a Lipschitz function $f_{\varepsilon}:X\to\mathbb{R}$ such that
$\|f-f_{\varepsilon}\|_{\infty}<\frac{\varepsilon}{3}.$
The function $f_{\varepsilon}$ is Lipschitz with some finite constant $L_{\varepsilon}$. Therefore,
$\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|\leq L_{\varepsilon}\cdot d_{W}(W_{n},Z).$
Since $d_{W}(W_{n},Z)\to 0$, there exists $N(\varepsilon)\in\mathbb{N}$ such that
$\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|<\frac{\varepsilon}{3},\quad\text{for all }n>N(\varepsilon).$
Using the triangle inequality, we have
$\left|\mathbb{E}[f(W_{n})]-\mathbb{E}[f(Z)]\right|$ $\leq\left|\mathbb{E}[f(W_{n})]-\mathbb{E}[f_{\varepsilon}(W_{n})]\right|$
$\quad+\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|$
$\quad+\left|\mathbb{E}[f_{\varepsilon}(Z)]-\mathbb{E}[f(Z)]\right|$
$<\frac{\varepsilon}{3}+\frac{\varepsilon}{3}+\frac{\varepsilon}{3}=\varepsilon.$
[…]
Proof of Theorem 4.
The proof proceeds in two steps. First, we establish Wasserstein convergence for a sequence of probability measures. The key argument in this step is an application of Stein’s method (Stein, 1972), specifically Theorem 3.5 of Ross (2011) for a dependency-graph-based variant, following the strategy also employed in HWS. We then adapt the norm-bounding technique of HWS to our stochastic setting, where both the model space and the structure of dependence differ. Second, we show that Wasserstein convergence implies weak convergence of the same sequence.
Clearly $\mathbb{E}[\hat{\theta}_{i}(z,\omega)]^{4}<\infty$ is equivalent to $\mathbb{E}[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})]^{4}<\infty$. Define $X_{i}=\nu_{ni}(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i}))$, and then $\mathbb{E}[X_{i}^{4}]<\infty$. By Theorem 2, we have also $\mathbb{E}[X_{i}]=0$.
Note that $\sum_{i=1}^{n}X_{i}=\hat{\tau}_{n}(z,\omega)-\tau_{n}$. Define $W_{n}=\sum_{i=1}^{n}X_{i}/\sigma_{n}$. Since the collection $\{X_{1},X_{2},...,X_{n}\}$ have dependency neighborhoods $N_{i}$, $i=1,...,n$, with $D_{n}=\max_{1\leq i\leq n}|N_{i}|$. Then we can now apply Theorem 3.5 in Ross (2011) and obtain
$d_{W}(W_{n},Z)\leq\frac{D_{n}^{2}}{\sigma_{n}^{3}}\sum_{i=1}^{n}\mathbb{E}|X_{i}|^{3}+\frac{\sqrt{26}D_{n}^{3/2}}{\sqrt{\pi}\sigma_{n}^{2}}\sqrt{\sum_{i=1}^{n}\mathbb{E}[X_{i}^{4}]},$ (D.4)
where $d_{W}(W_{n},Z)$ denotes the Wasserstein distance between $W_{n}$ and a standard normal one $Z$. Observe that (D.4) gives an upper bound of the distance.
First, we examine the first term in (D.4)
$\sum_{i=1}^{n}\mathbb{E}\left\lvert X_{i}\right\rvert^{3}$ $=\sum_{i=1}^{n}\mathbb{E}\left\lvert\nu_{ni}\left(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right)\right\rvert^{3}$
$=\sum_{i=1}^{n}\nu_{ni}^{3}\mathbb{E}\left\lvert\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right\rvert^{3}$
$\leq\sqrt{\sum_{i=1}^{n}\nu_{ni}^{6}}\cdot\sqrt{\sum_{i=1}^{n}\left(\mathbb{E}\left\lvert\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right\rvert^{3}\right)^{2}}$ (Cauchy-Schwartz)
$\leq n^{-5/2}\bar{\nu}^{3}\cdot\sqrt{\sum_{i=1}^{n}\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p_{1}}\,\left\lVert\psi_{i}\right\rVert_{q_{1}}\right)^{6}}$ (Lemma 1 and 2)
$\leq n^{-2}\bar{\nu}^{3}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}\right)^{3}.$
Next, we examine the second term in (D.4).
$\sum_{i=1}^{n}\mathbb{E}[X_{i}]^{4}$ $=\sum_{i=1}^{n}\mathbb{E}\left[\nu_{ni}\left(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right)\right]^{4}$
$=\sum_{i=1}^{n}\nu_{ni}^{4}\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{4}$
$\leq\sqrt{\sum_{i=1}^{n}\nu_{ni}^{8}}\cdot\sqrt{\sum_{i=1}^{n}\left(\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{4}\right)^{2}}$ (Cauchy-Schwartz)
$\leq n^{-7/2}\bar{\nu}^{4}\cdot\sqrt{\sum_{i=1}^{n}\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p_{2}}\,\left\lVert\psi_{i}\right\rVert_{q_{2}}\right)^{8}}$ (Lemma 1 and 2)
$\leq n^{-3}\bar{\nu}^{4}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}\right)^{4}.$
Thus, we obtain
$d_{W}(W_{n},Z)$ $\leq\frac{\bar{\nu}^{3}D_{n}^{2}}{n^{2}\sigma_{n}^{3}}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}\right)^{3}+\frac{\sqrt{26}\bar{\nu}^{2}D_{n}^{3/2}}{\sqrt{\pi}n^{3/2}\sigma_{n}^{2}}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}\right)^{2}$
$=O\left(\frac{D_{n}^{2}}{n^{2}\sigma_{n}^{3}}\right)+O\left(\frac{D_{n}^{3/2}}{n^{3/2}\sigma_{n}^{2}}\right),$ (D.5)
under Assumption 8, which states that $\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}$, $\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}$, $\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}$, and $\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}$ converge for some $p_{1},q_{1},p_{2},q_{2}\in[1,\infty]$ satisfying $31/p_{1}+1/q_{1}=1/$ 3 and $41/p_{2}+1/q_{2}=1/$ 4.
Given Assumption 9, (D.5) can be further bounded by
$d_{W}(W_{n},Z)\leq O\left(\frac{D_{n}^{2}}{n^{1/2}}\right)+O\left(\frac{D_{n}^{3/2}}{n^{1/2}}\right),$ (D.6)
which implies that, if $D_{n}=o(n^{1/4})$, then $d_{W}(W_{n},Z)\to 0$.
Lemma 3 shows that Wasserstein convergence implies weak convergence, which completes the proof. ∎
Lemma 4.
Let $\{\zeta_{i}\}_{i=1}^{n}$ be random variables such that
$\mathbb{E}[\zeta_{i}]=0\quad\text{and}\quad\sup_{i\in\mathbb{N}}\mathbb{E}[\zeta_{i}^{4}]<\infty.$
For deterministic weights $\nu_{ni}$ satisfying Assumption 5, define
$\Delta_{n}\;=\;n\sum_{(i,j)\in\mathcal{E}_{n}}\Bigl{(}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}-\nu_{ni}\nu_{nj}\,\mathbb{E}[\zeta_{i}\zeta_{j}]\Bigr{)}.$ (D.7)
Then
$\Delta_{n}\;=\;O_{p}\bigl{(}n^{-1/2}\,D_{n}^{3/2}\bigr{)}.$ (D.8)
Proof.
Set
$X_{ij}\;:=\;n\,\nu_{ni}\nu_{nj}\bigl{(}\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}]\bigr{)},\qquad(i,j)\in\mathcal{E}_{n},$
so that $\Delta_{n}=\sum_{(i,j)\in\mathcal{E}_{n}}X_{ij}$ and
$\mathbb{V}[\Delta_{n}]=\sum_{(i,j),(l,k)\in\mathcal{E}_{n}}\operatorname{Cov}[X_{ij},X_{lk}].$ (D.9)
Note that $\operatorname{Cov}[X_{ij},X_{lk}]=0$ whenever $X_{ij}$ and $X_{lk}$ are independent, so the sum involves only a restricted number of dependent pairs.
Since $\nu_{ni}\leq\bar{\nu}/n$, we have $n\,\nu_{ni}\nu_{nj}\leq\bar{\nu}^{2}/n$. Then, for any dependent $(i,j),(l,k)\in\mathcal{E}_{n}$, we apply Cauchy-Schwarz and the uniform fourth-moment bound $\sup_{i}\mathbb{E}[\zeta_{i}^{4}]<\infty$ to obtain
$\operatorname{Cov}[X_{ij},X_{lk}]\leq\left(\frac{\bar{\nu}^{2}}{n}\right)^{2}\cdot\mathbb{E}\Big{|}\big{(}\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}]\big{)}\big{(}\zeta_{l}\zeta_{k}-\mathbb{E}[\zeta_{l}\zeta_{k}]\big{)}\Big{|}\leq\frac{\bar{\nu}^{4}C}{n^{2}}$ (D.10)
for some constant $C>0$. This bound also applies when $(i,j)=(l,k)$.
Fix $(i,j)\in\mathcal{E}_{n}$. A second pair $(k,\ell)$ can be dependent with $(i,j)$ only if it lies within the corresponding neighbourhoods, that is, $(k,\ell)\cap(N_{i}\cup N_{j})\neq\emptyset$. Each unit has at most $D_{n}$ neighbours, so there are at most $D_{n}^{2}-1=O(D_{n}^{2})$ such pairs. Hence every $X_{ij}$ is dependent with at most $C_{1}D_{n}^{2}$ other summands for some constant $C_{1}>0$.
There are at most $nD_{n}$ ordered pairs in $\mathcal{E}_{n}$, so
$\mathbb{V}[\Delta_{n}]\;\leq\;C_{1}D_{n}^{2}\cdot nD_{n}\cdot\frac{\bar{\nu}^{4}C}{n^{2}}\;=\;C_{2}\,\bar{\nu}^{4}\,\frac{D_{n}^{3}}{n},$ (D.11)
for a constant $C_{2}>0$.
By Chebyshev’s inequality, for any $\varepsilon>0$,
$\mathbb{P}\bigl{(}|\Delta_{n}|>\varepsilon\bigr{)}\;\leq\;\frac{\operatorname{Var}(\Delta_{n})}{\varepsilon^{2}}\;\leq\;\frac{C_{2}\,\bar{\nu}^{4}}{\varepsilon^{2}}\,\frac{D_{n}^{3}}{n}.$
Since the right-hand side is of order $n^{-1}\,D_{n}^{3},$ we have
$\Delta_{n}\;=\;O_{p}\bigl{(}n^{-1/2}\,D_{n}^{3/2}\bigr{)},$
which completes the proof. ∎
Proof of Theorem 5.
By Theorem 2, we have $\mathbb{E}[\zeta_{i}]=0$, where $\zeta_{i}:=\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$. Assumption 10$(b)$ ensures that $\sup_{i\in\mathbb{N}}\mathbb{E}[\hat{\theta}_{i}(z,\omega)^{4}]<\infty$, which is equivalent to $\sup_{i\in\mathbb{N}}\mathbb{E}[\zeta_{i}^{4}]<\infty$, since $\theta_{i}(\tilde{y}_{i})$ is constant.
Lemma 4 together with Assumption 5 implies that
$n\hat{\sigma}_{n}^{2}-n\sigma_{n}^{2}=n\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\left(\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}]\right)=O_{p}(n^{-1/2}D_{n}^{3/2}).$
If Assumption 6 holds with $D_{n}=o(n^{1/3})$, (D.11) implies that this expression converges to zero in mean square. ∎
Proof of Corollary 1.
The result follows directly from Theorems 4 and 5. We write
$\hat{\sigma}_{n}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)=\frac{\sqrt{n}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)}{\sqrt{n}\sigma_{n}}\cdot\frac{\sqrt{n}\sigma_{n}}{\sqrt{n}\hat{\sigma}_{n}}.$
The first factor converges in distribution to $\mathcal{N}(0,1)$ by Theorem 4, and the second factor converges to 1 in probability by Theorem 5, together with Assumption 9, which guarantees that $\sqrt{n}\sigma_{n}$ is bounded away from zero.
Since the product of a sequence converging in distribution and another converging in probability (to a constant) also converges in distribution, Slutsky’s theorem yields the desired result. ∎
Lemma 5.
$\Delta_{n}^{c}\;=\;O_{p}\bigl{(}n^{-1/2}\,D_{n}^{3/2}\bigr{)}$.
Proof of Lemma 5.
Define
$\Delta_{n}^{c}\;=\;n\sum_{(i,j)\in\mathcal{E}_{n}^{c}}\left(\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}-\nu_{ni}\nu_{nj}\,\mathbb{E}[\zeta_{i}\zeta_{j}]\right),$ (D.12)
so that $n\hat{\sigma}_{cn}^{2}-n\sigma_{n}^{2}=\Delta_{n}^{c}$. To bound the variance of $\Delta_{n}^{c}$, observe that
$\mathbb{V}[\Delta_{n}^{c}]=\sum_{(i,j),(l,k)\in\mathcal{E}_{n}^{c}}\operatorname{Cov}[X_{ij},X_{lk}].$ (D.13)
Since $\mathcal{E}_{n}^{c}\subseteq\mathcal{E}_{n}$,
$\mathbb{V}[\Delta_{n}]-\mathbb{V}[\Delta_{n}^{c}]=\sum_{(i,j),(l,k)\in\mathcal{E}_{n}/\mathcal{E}_{n}^{c}}\operatorname{Cov}[X_{ij},X_{lk}].$
This difference can be positive when $\sum_{(i,j),(l,k)\in\mathcal{E}_{n}/\mathcal{E}_{n}^{c}}\operatorname{Cov}[X_{ij},X_{lk}]\geq 0$, but this cannot be guaranteed in general, as some of the covariance terms may be negative.
Nonetheless, we can assert that the number of terms in the upper bound of $\mathbb{V}[\Delta_{n}^{c}]$ is no greater than $C_{1}D_{n}^{2}$, and that the bound in (D.11) applies as well, since it is derived by summing the absolute values of the covariances. Therefore,
$\Delta_{n}^{c}\;=\;O_{p}\bigl{(}n^{-1/2}\,D_{n}^{3/2}\bigr{)},$
as claimed. ∎
Proof of Theorem 6.
By Lemma 5, $\Delta_{n}^{c}=O_{p}\left(n^{-1/2}D_{n}^{3/2}\right)$. In particular, if Assumption 6 holds with $D_{n}=o(n^{1/3})$, then $\hat{\sigma}_{cn}^{2}-\sigma_{n}^{2}$ goes to zero in mean square. ∎
Proof of Corollary 2.
The proof follows the same argument as in Corollary 1, replacing $\hat{\sigma}_{n}^{2}$ with $\hat{\sigma}_{cn}^{2}$. ∎
Proof of Theorem 7.
Define
$\tilde{\Delta}_{n}\;=\;n\sum_{(i,j)\in\tilde{\mathcal{E}}_{n}}\left(\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}-\nu_{ni}\nu_{nj}\,\mathbb{E}[\zeta_{i}\zeta_{j}]\right).$ (D.14)
Obviously, by Assumption 11, the upper bound of $\mathbb{V}[\tilde{\Delta}_{n}]$ lies between those of $\mathbb{V}[\Delta_{n}^{c}]$ and $\mathbb{V}[\Delta_{n}]$, and the result follows directly. ∎
Proof of Corollary 3.
The proof follows the same argument as in Corollary 1, replacing $\hat{\sigma}_{n}^{2}$ with $\tilde{\sigma}_{n}^{2}$. ∎