### v6

# Design-Based Inference under Random Potential Outcomes

Abstract
We introduce a design-based framework for causal inference that accommodates random potential outcomes without introducing outcome models, thereby extending the classical Neyman–Rubin paradigm in which outcomes are treated as fixed. Each unit’s potential outcome is modelled as a structural mapping $\tilde{y}_{i}(z,\omega)$, where $z$ denotes the treatment assignment and $\omega$ represents latent outcome-level randomness. Inspired by recent work linking design-based inference to the Riesz representation theorem, we embed random potential outcomes in a Hilbert space and define treatment effects as linear functionals, leading to estimators characterised by their Riesz representers. This approach preserves the core identification logic of randomised assignment while enabling valid inference under stochastic outcome variation. We establish consistency and asymptotic normality under local dependence, and develop feasible variance estimators that remain valid under weaker structural assumptions, including partially known dependence. A simulation study illustrates the robustness and finite-sample behaviour of the estimators. The proposed framework extends design-based causal inference to settings in which outcome-level randomness and local dependence are intrinsic features of the data.
Keywords: Design-based causal inference; Random potential outcomes; Riesz representation; Local dependence; Variance estimation; Simulation study.
1 Introduction
Randomised experiments remain central to the identification of causal effects. In the classical design-based perspective, pioneered by Neyman (1990) and Rubin (1974), potential outcomes are treated as fixed quantities, with all randomness attributed to the treatment assignment mechanism. This paradigm underpins influential estimators such as the Horvitz-Thompson estimator, whose identification relies exclusively on the experimental design. At a deeper level, the fixed-potential-outcomes assumption is not merely a technical convenience but an ontological stance regarding the nature of causal responses, one that becomes increasingly strained in stochastic environments. Importantly, the randomness of potential outcomes in our framework is not introduced to justify averaging across hypothetical populations, but to formalise outcome-level variability that is already implicit in many design-based estimands.
Building on this foundation, a substantial body of recent work has extended design-based inference to more complex experimental settings. In particular, methods have been developed to accommodate interference and dependence among units, including network and spatial designs (Aronow and Samii, 2017; Sävje et al., 2021; Athey et al., 2021). These approaches preserve the fixed-potential-outcomes assumption while modelling dependence through exposure mappings or structured interference. Other related contributions study robustness or misspecification in observational settings (Abadie et al., 2020; Imbens, 2004), but achieve identification through outcome modelling rather than randomisation.
A fundamental limitation of the classical Neyman–Rubin model is its assumption that potential outcomes are deterministic. This assumption implies that repeating the same experiment under identical conditions would yield identical outcomes, thereby neglecting intrinsic sources of variability such as measurement noise, stochastic biological responses, or environmental fluctuations. It further assigns all randomness to the assignment mechanism, overlooking the often substantial variability generated by the response process itself. More fundamentally, many causal estimands, including average treatment effects, are defined as expectations over potential outcomes or populations rather than over treatment assignments alone. Such definitions implicitly require a probabilistic structure on potential outcomes, a structure that the fixed-outcome paradigm does not formally provide.
This conceptual tension has motivated the widespread adoption of model-assisted approaches that treat potential outcomes as random variables. Seminal examples include Bayesian formulations of causal inference (Rubin, 1978), propensity score methods (Rosenbaum and Rubin, 1983), and recent developments in double machine learning (Chernozhukov et al., 2018). While these approaches relax the deterministic outcome assumption, they typically rely on parametric or semiparametric models for the outcome process, often specified as regressions on treatment and covariates. Such modelling assumptions can be restrictive, particularly in complex or high-dimensional environments. By contrast, classical design-based inference avoids these assumptions by not modelling the outcome surface, relying instead on randomisation for identification.
These considerations reflect a broader reassessment of the scope and limitations of contemporary causal inference. In a recent synthesis, Cinelli et al. (2025) highlight several unresolved challenges, including the need for new identification strategies, experimental designs accommodating complex dependence structures, and frameworks that balance realism with robustness. A central theme in their discussion is the trade-off between capturing intrinsic outcome variability and maintaining minimal modelling assumptions. The present work addresses this tension by extending the classical design-based paradigm to random potential outcomes without introducing outcome models. In this sense, our framework contributes directly to the class of new identification strategies emphasised by Cinelli et al. (2025).
Motivated by these considerations, we propose a design-based framework in which potential outcomes are modelled as random functions of the treatment assignment $z$ and a latent variable $\omega$, written as $\tilde{y}_{i}(z,\omega)$. This formulation reflects a basic principle of statistical inference: although the realised outcomes from a single experiment are fixed, the causal quantities of interest are typically defined as expectations over a broader stochastic environment. By modelling potential outcomes as random functions, we make this probabilistic structure explicit while retaining identification through randomised treatment assignment.
Our analysis is confined to experimental settings with known randomisation schemes. We do not consider observational designs, where identification requires unconfoundedness assumptions and covariate adjustment. The framework allows for spillover effects and continuous treatments.
Recent contributions, notably by Harshaw, Wang, and Sävje (2022, hereafter HWS), have reformulated aspects of classical design-based inference using the Riesz representation theorem. By representing potential outcomes as elements of a Hilbert space induced by the experimental design and defining treatment effects as linear functionals on that space, they provide a functional-analytic perspective that encompasses standard weighting estimators and accommodates settings with interference or continuous treatments. Their asymptotic analysis of consistency and normality draws on ideas from Stein’s method and the theory of dependency graphs (Stein, 1972; Ross, 2011), adapted to the design-based setting.
The present paper draws methodological inspiration from this Riesz-based framework but departs from it in a fundamental respect. While HWS operate within the classical fixed-potential-outcomes paradigm, we allow potential outcomes to be random functions defined on a latent probability space. This change in the object of inference necessitates a corresponding reformulation of causal estimands and asymptotic arguments, while preserving identification through randomised treatment assignment.
Despite this extension, the Riesz representation theorem remains central. Under appropriate regularity conditions, it continues to ensure that treatment effects can be expressed as inner products with a unique representer function. This shared structure provides a coherent foundation for analysing both fixed and random potential outcome settings within a unified design-based framework, with the classical model emerging as a degenerate special case.
The proposed framework yields several contributions. First, it extends classical design-based inference to random potential outcomes without introducing outcome models, thereby preserving identification through randomisation while accommodating intrinsic outcome variability. Second, it reformulates the Riesz representation approach in a stochastic potential outcome setting and provides sufficient conditions for consistency and asymptotic normality under local dependence, making explicit several convergence arguments that are typically left implicit in earlier analyses. Third, it admits feasible variance estimation under local dependence, compatible with both random and fixed potential outcome frameworks, enabling valid hypothesis testing based on a single experimental realisation without relying on overly conservative variance bounds. Finally, the framework offers a conceptual interpretation in terms of sample-based ergodicity, clarifying how unit-level averaging can recover population-level causal quantities in stochastic environments. Finite-sample simulations further demonstrate that the proposed estimator attains valid coverage and remains robust across a range of dependency structures.
Throughout, we use “random” to refer to outcome-level variation due to latent variables, and “stochastic” to describe modelling frameworks that explicitly incorporate such randomness.
The paper is organised as follows. Section 2 introduces the model and basic assumptions. Section 3 establishes large-sample properties of the estimator under local dependence. Section 4 presents feasible variance estimators and discusses their consistency. Section 5 covers computational aspects of estimating the Riesz representer in both finite- and infinite-dimensional settings. Section 6 provides a simulation study. Section 7 concludes with a discussion of the broader implications and possible extensions.
2 Model Setup
2.1 The Potential Outcomes
We begin by considering a setting with $n$ units, where
Assumption 1 (The Stochastic Setting).
each unit’s potential outcome is modelled as a general measurable function $y_{i}(z,x_{i},\epsilon_{i})$ with the following arguments: $z=(z_{1},\dots,z_{n})$ denoting the vector of treatment assignments; $x_{i}=x_{i}(\omega)$ representing the possibly observed covariates for unit $i$; $\epsilon_{i}=\epsilon_{i}(\omega)$ denoting the idiosyncratic error or unobserved heterogeneity for unit $i$. Here, $\omega\in\Omega$ is a latent random element defined on a common probability space $(\Omega,\mathcal{F}_{\omega},P)$ shared by all units.
Furthermore, we define the structural mapping from the latent space $\Omega$ to variables $(x_{i},\epsilon_{i})$ and represent the potential outcome via the composition
$\tilde{y}_{i}(z,\omega)=y_{i}(z,x_{i}(\omega),\epsilon_{i}(\omega)),$ (2.1)
since $x_{i}$ and $\epsilon_{i}$ are measurable functions of $\omega$.
We refer to this as the stochastic setting because the potential outcomes $\tilde{y}_{i}(z,\omega)$ are explicitly modelled as random functions of a latent variable $\omega$, introducing outcome-level randomness into the design-based framework. This formulation enables the analysis of causal effects under a stochastic data-generating process and bridges the gap between classical design-based inference and super-population perspectives.
Although $\omega$ is not unit-specific, its structure permits different subcomponents to influence different units, thereby being able to preserve unit-specific variability and independence and accommodating heterogeneous potential outcomes.
The potential outcome for unit $i$ depends on the entire treatment vector $z$, so that interference (or spillover effects) is permitted. Notably, one may also consider the special case of no spillover effects, whereby
$y_{i}(z,x_{i},\epsilon_{i})=y_{i}(z_{i},x_{i},\epsilon_{i}),$ (2.2)
which corresponds to a version of the Stable Unit Treatment Value Assumption (SUTVA) incorporating consistency, as introduced by Rubin (1980). In the present work we do not impose the no-spillover condition a priori; rather, we allow for general interference and propose to test for its presence.
In practice, it typically suffices to allow heterogeneity across units only through $x_{i}$ and $\epsilon_{i}$, making it unnecessary to distinguish $y_{i}$ explicitly. In such cases, one may work with a common structural function $y$, while still obtaining heterogeneous mappings $\tilde{y}_{i}(z,\omega)=y(z,x_{i},\epsilon_{i})$ across units.
The structural mapping (2.1) is the primary object of interest in what follows.
Denoting $\mathcal{Z}$ as the treatment assignment space, equipped with its Borel sigma-algebra $\mathcal{F}_{z}$, we presume that the structural mapping $\tilde{y}_{i}:\mathcal{Z}\times\Omega\to\mathbb{R}$ is measurable with respect to the sigma-algebra $\mathcal{F}_{z}\otimes\mathcal{F}_{\omega}$ on the domain and the Borel sigma-algebra $\mathcal{B}(\mathbb{R})$ on the codomain. This ensures that the relevant random variables and their expectations are well-defined, and it also serves as a prerequisite for defining inner products and applying Hilbert space tools in the sequel. Note that the measurability of the covariates is sometimes required to ensure that conditional expectations and probabilities given $x_{i}$ are well-defined within the underlying probability space, although we do not explicitly rely on this property in the present paper.
Following randomisation principles, we assume that
Assumption 2 (Randomisation).
the treatment assignment vector $z$ is drawn from a known randomisation distribution that is independent of the outcome-generating process given the latent variable $\omega$.
Assumption 2 is not required for the unbiasedness or asymptotic validity of the Riesz estimator. Its purpose is operational rather than inferential: it ensures that the design distribution $\mu$ remains valid when conditioning on $\omega$, which is essential for defining the Riesz representer via the inner product induced by $\mu$. Once the representer $\psi_{i}$ is constructed, all subsequent inference, including estimation and asymptotic analysis, proceeds without requiring independence between $z$ and $\omega$. This assumption will be invoked again in Section 5, where the Gram matrix is computed in order to estimate the Riesz representer of the potential outcomes.
This assumption also justifies the use of the product measure $\mu(\mathop{}\!\mathrm{d}z)\,P(\mathop{}\!\mathrm{d}\omega)$ in the corresponding integrals. It allows expectations of the form $\mathbb{E}_{z,\omega}[\cdot]$ to be written as iterated expectations $\mathbb{E}_{z}\mathbb{E}_{\omega}[\cdot]$. Throughout, when writing integrals or expectations in this way, we implicitly invoke this assumption.
It is worth noting that this assumption corresponds to the assumption of ignorability or unconfoundedness, as articulated in Rubin (1978) and further developed in Rosenbaum and Rubin (1983) for observational studies. Under this formulation, $\omega$ governs all outcome-related randomness across units, while the randomisation mechanism for $z$ remains externally controlled and separate (see Neyman (1990) and Fisher (1935)).
Conceptually, once a particular realisation of $\omega$ is fixed, each unit’s potential outcome function $\tilde{y}_{i}(z,\omega)$ becomes deterministic in $z$; hence, the only source of randomness in the observed outcomes from the randomised assignment of treatments. This approach preserves the fundamental idea that each unit has a well-defined potential outcome under every treatment assignment, capturing the core principle of SUTVA (or, more generally, the notion of well-defined potential outcomes), even as we permit these outcomes to vary randomly across realisations of $\omega$.
This formulation is flexible and expressive in that it captures both unconfoundedness and confounding, depending on the relationship between $z$ and the latent variable $\omega$. If $z$ and $\omega$ are independent, then $\tilde{y}_{i}(z^{0},\omega)$ and $\tilde{y}_{i}(z^{1},\omega)$ are independent of $z$. More generally, if $z$ and $\omega$ are independent given $x\in X_{0}$, then $y_{i}(z^{0},x,\epsilon_{i})$ and $y_{i}(z^{1},x,\epsilon_{i})$ are independent of $z\mid x\in X_{0}$, for any measurable $X_{0}\subset\mathcal{F}_{x}:=\{A\subseteq x_{i}(\Omega):x_{i}^{-1}(A)\in\mathcal{F}_{\omega}\}$. In randomised experiments, the design ensures that $z$ is independent of $(x_{i},\epsilon_{i})$, mitigating confounding. In observational studies, however, it is imperative to adjust for $x_{i}$ to control for confounding. In either case, this framework models potential outcomes as random rather than fixed.
2.2 Model Space and Inner Product
Let $L^{p}(\mathcal{Z}\times\Omega)$ denote the space of all measurable functions $u:\mathcal{Z}\times\Omega\to\mathbb{R}$, satisfying $\mathbb{E}[|u(z,\omega)|^{p}]<\infty$, for some integer $p\geq 1$. For notational convenience, we henceforth write $L^{p}$ in place of $L^{p}(\mathcal{Z}\times\Omega)$, whenever the meaning is clear from context. Furthermore, we denote the $p$-norm of $u\in L^{p}$ by
$\|u\|_{p}:=\left(\int_{\mathcal{Z}\times\Omega}\left|u(z,\omega)\right|^{p}\mu(\mathop{}\!\mathrm{d}z)P(\mathop{}\!\mathrm{d}\omega)\right)^{1/p}=\left(\mathbb{E}\left|u(z,\omega)\right|^{p}\right)^{1/p},$ (2.3)
where $\mu$ denotes the probability measure for randomisation. Under Assumption 2, the joint law of $(z,\omega)$ factorises as $\mu(\mathop{}\!\mathrm{d}z)P(\mathop{}\!\mathrm{d}\omega)$, which permits rewriting the $L^{p}$-norm as an expectation. We write $\|u\|:=\|u\|_{2}$ for brevity if $u\in L^{2}$. Note that $p$-norm characterises the $p$-moment of the random variable $u$.
We model each unit’s potential outcome function, via the structural mapping (2.1), as an element of a model space $\mathcal{M}_{i}$, which is a subspace of $L^{2}$. Formally, we write
Assumption 3 (Model Space).
$\tilde{y}_{i}(z,\omega)\in\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega).$ (2.4)
Our construction is formulated in a distinct functional setting that explicitly separates treatment assignments from latent variation induced by unobserved random elements. The model space incorporates both an observable assignment vector $z$ and a latent random element $\omega$, capturing two sources of variation. This separation underpins our interpretation of potential outcomes as random functions and motivates a stochastic modelling framework. In particular, the space $\mathcal{M}_{i}$ is defined over random potential outcomes, and the associated inner product reflects variation due to both assignment and latent structure. Importantly, the spaces from which $z$ and $\omega$ are drawn may differ in structure: while $\omega$ is modelled as a general element of a latent probability space, $z$ is often a finite-dimensional or even discrete vector representing treatment assignments. This structural heterogeneity is explicitly accommodated in our formulation of the product space $\mathcal{Z}\times\Omega$ and the associated function space $\mathcal{M}_{i}$.
The framework of HWS can be viewed as a special case of our model, corresponding to the degenerate setting where $\Omega=\{\omega_{0}\}$ for some $\omega_{0}$. In this case, the latent variable $\omega_{0}$ is effectively fixed, and the realised data serve as the entire outcome-generating mechanism. Thus, the potential outcome function (2.1) becomes a deterministic function of $z$, recovering the fixed-outcome setting in the Neyman–Rubin tradition, as adopted in their work.
We equip $\mathcal{M}_{i}$ with the inner product
$\langle u,v\rangle:=\mathbb{E}_{\omega}\Bigl[\mathbb{E}_{z}\bigl[u(z,\omega)\,v(z,\omega)\bigr]\Bigr]=\mathbb{E}\bigl[u(z,\omega)\,v(z,\omega)\bigr],$ (2.5)
where the second equality follows from the law of iterated expectations. The associated norm is given by $\|u\|=\sqrt{\langle u,u\rangle}$, which coincides with the standard $L^{2}$-norm.
In the degenerate case where $\Omega=\{\omega_{0}\}$, the inner product reduces to
$\langle u,v\rangle^{0}=\mathbb{E}_{z}\bigl[u(z,\omega_{0})\,v(z,\omega_{0})\bigr],$ (2.6)
which coincides with the inner product employed in HWS. This structure enables us to apply the Riesz representation theorem to construct estimators for treatment effects, even in the presence of interference.
2.3 Treatment Effect Functionals and Riesz Representation
While our model space $\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega)$ may not be complete under the inner product defined in (2.5), this poses no issue in what follows. Any inner product space admits a Hilbert space completion, and the Riesz representation theorem applies in that completed space. Hence, throughout this paper, whenever necessary, we implicitly work with the closure of $\mathcal{M}_{i}$ in $L^{2}(\mathcal{Z}\times\Omega)$.
Let $\mathcal{M}_{i}^{\prime}$ denote the dual space of $\mathcal{M}_{i}$.
Assumption 4 (Dual Representability of the Treatment Effect).
The treatment effect $\theta_{i}$ for each unit $i$ can be represented as a continuous linear functional on the model space of potential outcomes, viz., $\theta_{i}:\mathcal{M}_{i}\to\mathbb{R}$ with $\theta_{i}\in\mathcal{M}_{i}^{\prime}$.
This assumption reflects a standard construction in functional analysis, where a treatment effect is represented as a continuous linear functional on a function space of potential outcomes. Such a formulation enables the use of the Riesz representation theorem, which yields an estimator expressed as an inner product with a representer function. This approach applies to both classical fixed-outcome frameworks and to stochastic settings where potential outcomes depend on latent variables.
Note that functions in an $L^{2}$ space are identified through equivalence classes, meaning two functions differing only on a set of measure zero are considered identical. Consequently, a well-defined linear functional must assign the same value to all functions within the same equivalence class.
A potential pitfall arises when attempting to use evaluation functionals, such as $\theta(u)=u(x_{0})$, that assign values based on a single point. Such evaluation maps generally are not well-defined in an $L^{2}$ context because they may yield different values for functions identical almost everywhere, violating linearity and the identification of elements in the same equivalence class.
In our analysis, we explicitly avoid these issues by requiring that the treatment effect functional $\theta_{i}$ be defined consistently on equivalence classes. More specifically, if two functions differ by a function of zero $L^{2}$-norm (i.e., distance zero), the functional $\theta_{i}$ shall assign them the same value.
From Assumption 4, we also require that the linear functional $\theta_{i}$ on a normed space is continuous or equivalently bounded. A heuristic justification is that requiring the treatment effect functional to be continuous ensures that small perturbations in the potential outcome function (as measured by the $L^{2}$ norm) result in small changes in the treatment effect. This continuity is essential for invoking the Riesz representation theorem, which in turn guarantees that $\theta$ can be uniquely represented by an element of the Hilbert space.
In all cases, $\theta_{i}$ is a predetermined linear functional on $\mathcal{M}_{i}$, defined by the researcher to encode how treatment effects are to be evaluated from potential outcome functions. While $\theta_{i}$ may differ across individuals, it is fixed throughout the analysis and not learned from data.
Researchers need not be concerned about such technicalities in practice, as they are always free to define or adjust their treatment effect functional to comply with the requirements in Assumptions 4. For instance, one might consider the following examples of continuous linear functionals.
Example 1.
Expected treatment effect for binary treatment
$\theta_{i}(u)\;=\;\mathbb{E}_{\omega}\bigl[u(z^{(1)},\omega)-u(z^{(0)},\omega)\bigr]$ (2.7)
representing the expected outcome difference when shifting treatment assignment from $z^{(0)}$ to $z^{(1)}$.
Example 2.
Expected treatment effect for treatment intervals or regions
$\theta_{i}(u)\;=\;\mathbb{E}_{\omega}\Bigl[\mathbb{E}_{z\in A}\bigl[u(z,\omega)\bigr]\Bigr]-\mathbb{E}_{\omega}\Bigl[\mathbb{E}_{z\in B}\bigl[u(z,\omega)\bigr]\Bigr],$ (2.8)
where $A,B\subset\mathcal{Z}$ are measurable sets with $A\cap B=\emptyset$, representing the expected outcome difference between treatment groups $A$ and $B$.
Example 3.
Expected partial derivative
$\theta_{i}(u)=\mathbb{E}_{\omega}\!\left[\partial_{z}u(z_{0},\omega)\right],$ (2.9)
at some point $z_{0}\in\mathcal{Z}$, representing the expected marginal effect at $z_{0}$.
Such functionals arise naturally when treatment intensity is continuous and marginal effects are of primary interest.
Example 4.
Test function weighted expected partial derivative
$\theta_{i}(u)=\mathbb{E}_{\omega}\!\left[\int_{\mathcal{Z}}\partial_{z}u(z,\omega)\,\varphi(z)\,\mathop{}\!\mathrm{d}z\right],$ (2.10)
where $\varphi$ is a predetermined test function that weights the partial derivative over $z\in\mathcal{Z}$. In particular, taking $\varphi=\mathbf{1}_{A}$ for a measurable set $A\subset\mathcal{Z}$ yields
$\theta_{i}(u)=\mathbb{E}_{\omega}\!\left[\int_{A}\partial_{z}u(z,\omega)\,\mathop{}\!\mathrm{d}z\right],$ (2.11)
representing the expected marginal effect averaged over region $A$.
The following propositions show that these functionals, under certain conditions, can be linear and continuous.
Proposition 1.
Suppose that $\mu(A)>0$ and $\mu(B)>0$. The functional (2.8) is linear and continuous on $\mathcal{M}_{i}$.
Note that Example (2.7) is a special case of Example (2.8), obtained by setting $A=\{z^{(1)}\}$ and $B=\{z^{(0)}\}$. The condition $\mu(A)>0$ and $\mu(B)>0$ is commonly referred to as the positivity or overlap assumption in the causal inference literature. It ensures that both treatment groups are represented in the randomisation distribution, which guarantees that the corresponding estimand is well-defined and estimable. In our case, the positivity assumption plays a crucial role in verifying that the functional is both linear and continuous.
Proposition 2.
[…]
This result implies that, for the functional to satisfy the necessary continuity condition, the model space must be further restricted to $\mathcal{M}_{i}\subset L^{2}(\Omega;H^{s}(\mathcal{Z}))\subset L^{2}(\mathcal{Z}\times\Omega)$. Within this space, functions admit $C^{1}$ representatives in the $z$-variable almost surely, so that pointwise evaluation of derivatives is well-defined and continuous.
Proposition 3.
Let $\mathcal{Z}\subset\mathbb{R}^{d}$ be open and bounded and let $\varphi\in L^{2}(\mathcal{Z})$. Assume $u\in L^{2}(\Omega;H^{1}(\mathcal{Z}))$. Then the functional (2.10) is linear and continuous on $L^{2}(\Omega;H^{1}(\mathcal{Z}))$.
This proposition is deliberately permissive in that the test function $\varphi$ is only required to belong to $L^{2}(\mathcal{Z})$ and is therefore not restricted to indicator functions such as $\mathbf{1}_{A}$ in (2.11). More general weights, including discontinuous or piecewise-defined functions, are admissible. In contrast, Dirac-type test functions corresponding to point evaluation are not elements of $L^{2}(\mathcal{Z})$ and hence fall outside the scope of this result. Such pointwise functionals require additional regularity and are instead covered by Proposition 2.
Theorem 1 (Riesz Representation Theorem in $\mathcal{M}_{i}$).
Under Assumptions 1, 3, and 4, for every $\theta_{i}\in\mathcal{M}_{i}^{\prime}$, there exists a unique element $\psi_{i}\in\mathcal{M}_{i}$ such that
$\theta_{i}(u)=\langle u,\psi_{i}\rangle\quad\text{for all }u\in\mathcal{M}_{i}.$ (2.12)
This theorem follows as a special case of the classical Riesz representation theorem when $\mathcal{M}_{i}$ is a Hilbert space. The resulting Riesz representer $\psi_{i}$ plays a central role in our estimation framework.
This approach retains the mathematical elegance of Riesz-based identification while extending it to accommodate outcome-level stochasticity. It lays the foundation for valid estimation in experiments where heterogeneity arises not only from treatment variation but also from latent random effects.
By constructing the Riesz representer $\psi_{i}$ corresponding to the treatment effect functional $\theta_{i}$, we only consider those treatment effects that are representable in the dual space $\mathcal{M}_{i}^{\prime}$. Consequently, we estimate the treatment effect
$\theta_{i}(\tilde{y}_{i})=\langle\tilde{y}_{i},\psi_{i}\rangle=\mathbb{E}\Bigl[\tilde{y}_{i}(z,\omega)\,\psi_{i}(z,\omega)\Bigr],$ (2.13)
by using the corresponding Riesz estimator
$\hat{\theta}_{i}(z,\omega)=\tilde{y}_{i}(z,\omega)\,\psi_{i}(z,\omega),$ (2.14)
as, in practice, we observe only one realisation of the outcome $\tilde{y}_{i}(z,\omega)$ for certain $z$ and $\omega$. It follows immediately that the Riesz estimator in (2.14) for each unit is unbiased for its treatment effect given in (2.13).
Definition 1 (Aggregate Estimand).
The finite-sample estimand we consider in this paper is defined as a weighted average of individual treatment effects
$\tau_{n}=\sum_{i=1}^{n}\nu_{ni}\theta_{i}(\tilde{y}_{i}),$ (2.15)
where the weights $\nu_{ni}\in\mathbb{R}$ satisfy $\nu_{ni}\geq 0$.
This quantity corresponds to an average treatment effect under a general weighting scheme, aligning with the broader framework of causal inference, where estimands are typically defined as weighted averages of unit-level causal effects across different designs and sampling schemes. Our setup allows for general fixed weighting schemes $\nu_{i}$, providing greater flexibility in defining estimands.
More importantly, $\tau_{n}$ reflects an integration over latent randomness $\omega$, and thus captures variation across a distribution of potential outcomes, effectively averaging over “parallel worlds” indexed by $\omega$. In contrast to the classical Neyman–Rubin framework, where potential outcomes are fixed and comparisons are defined across units within a single universe, our formulation treats outcomes as random functions and estimands as expectations over both treatment assignments and latent heterogeneity. This shift from fixed to random potential outcomes leads to a richer class of estimands, grounded in the observed design but expressive enough to reflect underlying stochastic complexity. It also conceptually aligns with super-population inference while maintaining the rigour of design-based identification.
The estimand $\tau_{n}$ does not rely on Assumption 2; it is mathematically well-defined regardless of whether $z$ and $\omega$ are independent. However, the dependence structure between $z$ and $\omega$ influences the potential outcomes $\tilde{y}_{i}$ and thus affects the value of the estimand. In the causal inference literature, selecting the right estimand involves specifying a target quantity that captures the scientific question of interest under the assumed data-generating process. Different assumptions about the relationship between $z$ and $\omega$ lead to different estimands, each carrying its own causal or associational interpretation.
The estimand considered by HWS corresponds to the special case of our formulation,
$\tau_{n}^{0}=\frac{1}{n}\sum_{i=1}^{n}\langle\tilde{y}_{i},\psi_{i}\rangle^{0}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}_{z}\Bigl[\tilde{y}_{i}(z,\omega_{0})\,\psi_{i}(z,\omega_{0})\Bigr],$ (2.16)
where $\Omega=\{\omega_{0}\}$ is a singleton probability space. In this special case, the treatment effect is deterministic and the average is equally weighted.
A common choice for the weights is the uniform weighting $\nu_{ni}=1/n$, which yields the sample average treatment effect (SATE). More generally, this formulation accommodates subgroup-specific or covariate-adjusted averages, depending on the structure of the weights. We impose the following assumption to control the magnitude of the weights $\nu_{ni}$.
Assumption 5 (Uniform Weight Upper Bound).
There exists a constant $\bar{\nu}>0$ such that
$n\nu_{ni}\leq\bar{\nu}\quad\text{for all }i\text{ and all }n\in\mathbb{N}.$ (2.17)
It guarantees that each weight is bounded by a constant multiple of $1/n$, uniformly in $i$, without requiring the individual sequences $\{n\nu_{ni}\}_{n\geq i}$ to converge. Consequently, no finite subset of indices can receive an asymptotically dominant share of the total weight. Note that Assumption 5 implies $\sup_{n\to\infty}\sum_{i=1}^{n}\nu_{ni}<\infty$ but not necessarily one, which is quite general. In the continuous analogue this corresponds to the integrability of the weight function.
Given the unit-level Riesz estimators (2.14) and the weights,
Definition 2 (Aggregate Riesz Estimator).
we define the corresponding aggregate Riesz estimator for the finite-sample estimand
$\hat{\tau}_{n}(z,\omega)=\sum_{i=1}^{n}\nu_{ni}\hat{\theta}_{i}(z,\omega).$ (2.18)
Thus we have the unbiasedness of the Riesz estimator in the following.
Theorem 2 (Unbiasedness of the Riesz Estimator).
The Riesz estimator defined in (2.14) is unbiased for the individual treatment effect, i.e., $\mathbb{E}\bigl[\hat{\theta}_{i}(z,\omega)\bigr]=\theta_{i}(\tilde{y}_{i})$. Consequently, the aggregate Riesz estimator in (2.18) is unbiased for the finite-sample estimand (2.15), i.e., $\mathbb{E}\Bigl[\hat{\tau}_{n}(z,\omega)\Bigr]=\tau_{n}$.
3 Large-Sample Properties of the Aggregate Riesz Estimator
In this section, we clarify the notion of local independence required in our framework, which is a concept that has been widely adopted in recent literature on causal inference under interference. Informally, local dependence refers to the idea that subsets of random variables are conditionally independent of those outside their respective neighbourhoods. This structure is often represented using a dependency graph, in which vertices correspond to units and edges encode potential statistical dependence. For further information about the dependence structure of the variables, see for example Chen and Shao (2004).
Definition 3 (Dependency Neighbourhoods).
For each unit $i$, let $\{M_{ik}\}_{k\in I_{i}}$ be the collection of all subsets $M_{ik}\subset\{1,2,\dots,n\}$ such that the pair of random variables $(\tilde{y}_{i},\psi_{i})$ is independent of the set $\left\{(\tilde{y}_{j},\psi_{j}):j\notin M_{ik}\right\}$. We define the dependency neighbourhood of unit $i$ as the intersection over all such sets:
$N_{i}:=\bigcap_{k\in I_{i}}M_{ik}.$ (3.1)
This definition ensures that $N_{i}$ is the smallest subset of units for which independence holds. Formally, for all such $j\notin N_{i}$, and for all measurable sets $A_{i}\in\mathcal{F}_{\tilde{y}_{i}}$, $B_{i}\in\mathcal{F}_{\psi_{i}}$, $A_{j}\in\mathcal{F}_{\tilde{y}_{j}}$, and $B_{j}\in\mathcal{F}_{\psi_{j}}$, we have
$\mathbb{P}\left((A_{i}\cap B_{i})\cap(A_{j}\cap B_{j})\right)=\mathbb{P}(A_{i}\cap B_{i})\,\,\mathbb{P}(A_{j}\cap B_{j}).$ (3.2)
This definition corresponds to Definition 3.1 in Ross (2011) and is a weaker version of Definition 5.1 in HWS, as it only requires independence of the potential outcome $\tilde{y}_{i}$ and the Riesz representer $\psi_{i}$ from those outside $N_{i}$, rather than full functional independence across the spaces $\mathcal{M}_{i}$ and $\mathcal{M}_{j}$.
In contrast to HWS, who assume independence of the full model spaces across units, our framework only requires independence between the observable random pairs $(\tilde{y}_{i},\psi_{i})$. This relaxation is justified by the introduction of a latent random variable $\omega$, under which potential outcomes are modelled as conditionally random functions. Importantly, our formulation permits cases where the model spaces coincide, i.e. $\mathcal{M}_{i}=\mathcal{M}_{j}$, so long as the associated outcome representer pairs remain independent. Consequently, we do not assume functional independence at the space level, but merely require that the realised outcomes and their associated Riesz representers be independent across sufficiently separated units. This distinction both broadens the scope of our analysis and clarifies the assumptions necessary for variance control and asymptotic normality in a design-based setting.
We denote by $\left\lvert N_{i}\right\rvert$ the number of units in the dependency neighbourhood of unit $i$, which we refer to as the neighbourhood size for unit $i$. Let $D_{n}=\max_{1\leq i\leq n}\left\lvert N_{i}\right\rvert$ and $d_{n}=n^{-1}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert$ denote, respectively, the maximum and average neighbourhood sizes across the sample. These quantities will play a central role in the asymptotic analysis to follow.
In practice, the specification of neighbourhoods $N_{i}$ depends on substantive knowledge of the experimental context. In network experiments, for example, $N_{i}$ might include unit $i$ and its observed social or physical connections. In spatial designs, it is common to define $N_{i}$ via fixed-radius rules or $k$-nearest-neighbour relationships based on geographic distance. When no such structure is explicitly available, one can conservatively define $N_{i}$ to include all units within the same covariate stratum or treatment block as unit $i$. These strategies allow researchers to encode plausible channels of dependence without imposing strong model-based assumptions.
The following two assumptions restrict the growth rate of the dependency neighbourhoods.
Assumption 6 (Maximum Dependency Neighbourhood Size).
The maximum dependency neighbourhood size $D_{n}$ satisfies $D_{n}=o(n^{d})$ for some $d>0$.
Assumption 7 (Average Dependency Neighbourhood Size).
The average dependency neighbourhood size $d_{n}$ satisfies
  1. (a)
    $d_{n}=o(n)$;
  2. (b)
    $\sup_{n\in\mathbb{N}}d_{n}<\infty$.
Assumptions 6 and 7(a) regulate the growth of dependency neighbourhoods and play a role analogous to sparsity constraints commonly used in the analysis of dependent data. Assumption 7(b) additionally imposes uniform boundedness of the average neighbourhood size. These conditions are standard in the local dependence literature and are stated here explicitly to clarify their role and facilitate reference in our self-contained derivations.
In addition, for a sequence of functions $u_{1},\dots,u_{n}\in L^{p}$, where $p>1$, we define the max-$p$ norm as
$\left\lVert u\right\rVert_{\max,p}^{n}:=\max_{1\leq i\leq n}\left\lVert u_{i}\right\rVert_{p}.$ (3.3)
The sequence $\left\lVert u\right\rVert_{\max,p}^{n}$ is clearly non-decreasing in $n$, and, as such, is bounded if and only if it converges as $n\to\infty$.
We now state a proposition that provides an upper bound on the asymptotic variance, denoted by $\sigma_{n}^{2}:=\mathbb{V}[\hat{\tau}_{n}(z,\omega)]$, of the aggregate Riesz estimator $\hat{\tau}_{n}(z,\omega)$ of the estimand $\tau_{n}$. This result follows from the setup and notation introduced above, and it serves to justify or at least suggest the consistency of the estimator under appropriate conditions.
Proposition 4 (Variance Upper Bound).
Under Assumptions 1 and 35, the following inequality holds
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\,\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2},$ (3.4)
for any $p,q\in[1,\infty]$ satisfying $1/p+1/q=1/2$. The inequality holds trivially if either $\|\tilde{y}_{i}\|_{p}$ or $\|\psi_{i}\|_{q}$ has no finite upper bound.
Furthermore, we define $S_{2}:=\left\{(p,q)\in[1,\infty]^{2}:\,\frac{1}{p}+\frac{1}{q}=\frac{1}{2}\right\}$, and then we have
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}d_{n}}{n}\left(\inf_{(p,q)\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$ (3.5)
The inequality hold trivially if $\inf_{p,q\in S}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}$ has no finite upper bound.
Proposition 4 provides an upper bound on the variance of the aggregate estimator. It shows that the variance decreases with sample size under boundedness of unit-level norms and local sparsity. The result highlights the interaction between dependence complexity and unit-level variability in determining inferential precision.
It implies that the variance may decay as the sample size increases, provided that the norms $\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}$ and $\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}$ remain bounded if there exists such a pair $p,q\in[1,\infty]$ satisfying $1/p+1/q=1/2$, and that the neighbourhood sizes $\left\lvert N_{i}\right\rvert$ are uniformly small. The bound highlights the importance of controlling both the complexity of dependence and the magnitude of unit-level variation in order to obtain valid inference.
Assumption 8 (Boundedness of Norms).
For some $r\geq 1$, there exist real numbers $p,q\in[r,\infty]$ satisfying $1/p+1/q=1/r$ such that
$\sup_{i\in\mathbb{N}}\left\lVert\tilde{y}_{i}\right\rVert_{p}<\infty\quad\text{and}\quad\sup_{i\in\mathbb{N}}\left\lVert\psi_{i}\right\rVert_{q}<\infty.$ (3.6)
This assumption is formulated to allow flexibility in the choice of $r$. Intuitively, the condition imposes uniform boundedness on the $p$th moments of $\tilde{y}_{i}$ and the $q$th moments of $\psi_{i}$, thereby excluding heavy-tailed behaviour.
Note that the max-$p$ norm satisfies $\left\lVert u\right\rVert_{max,p}^{n}\leq\sup_{i\in\mathbb{N}}\left\lVert u_{i}\right\rVert_{p}$. Therefore, Assumption 8 is equivalent to requiring that the four quantities $\|\tilde{y}\|_{\max,p_{1}}^{n}$, $\|\psi\|_{\max,q_{1}}^{n}$, $\|\tilde{y}\|_{\max,p_{2}}^{n}$, and $\|\psi\|_{\max,q_{2}}^{n}$ remain bounded and converge as $n\to\infty$.
The following result follows naturally from our variance control framework and the assumptions stated above.
Theorem 3 (Consistency).
Under Assumptions 1, 35, 7(a)(a), and 8(r=2), the aggregate Riesz estimator in (2.18) is consistent in mean square. Moreover, if 7(b)(b) holds, then the aggregate Riesz estimator satisfies $O_{p}(n^{-1/2})$.
We should interpret $O_{p}(n^{-1/2})$ with care. This notation means that the sequence $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ is bounded in probability, but it does not imply convergence in distribution to a single non-degenerate law. In fact, the sequence may vary between different limiting distributions with bounded variances.
This result has an important practical implication. Although the aggregate Riesz estimator $\hat{\tau}_{n}$ is computed from a single realised treatment assignment $z$ and a single latent realisation $\omega$, Theorem 3 shows that, under mild regularity conditions, averaging across units is sufficient for consistent estimation.
More precisely, the theorem implies a form of sample-based ergodicity: for almost every realisation $(z,\omega)$, the estimation error of $\hat{\tau}_{n}(z,\omega)$ is of order $O_{p}(n^{-1/2})$, even though no repeated sampling of $(z,\omega)$ is available. As the cross-sectional sample size grows, averaging across units substitutes for averaging over hypothetical repetitions of the treatment assignment and the latent environment.
This interpretation clarifies that consistency is achieved not through repeated experimentation, but through structural regularity and sparsity in cross-unit dependence. Consequently, a single realised experiment contains sufficient information for reliable estimation, provided that the local dependence conditions are satisfied.
This sample-based ergodicity perspective serves as a standing interpretative principle for all subsequent large-sample results in the paper. In particular, the asymptotic normality and variance estimation results that follow are to be understood in the same way.
So far, we have shown that, under certain conditions, the aggregate Riesz estimator is consistent. However, these conditions alone are not sufficient for establishing asymptotic normality; an observation that aligns with the fact that consistency, in general, does not imply asymptotic normality.
The following non-degeneracy assumption imposes a lower bound on the variance of the aggregate Riesz estimator and plays a critical role in ensuring asymptotic normality. When combined with the previous assumptions, it provides sufficient conditions for the aggregate Riesz estimator to satisfy a central limit theorem.
Assumption 9 (Variance Lower Bound).
There exists $\sigma_{0}>0$ such that $\inf_{n\in\mathbb{N}}\sqrt{n}\sigma_{n}\geq\sigma_{0}$.
This condition imposes a non-degeneracy requirement on the estimator. It permits the asymptotic variance $\sigma_{n}^{2}$ to decay with $n$, but not too rapidly, thus ensuring that the root-$n$ scaled estimator $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ retains a non-degenerate variance in the limit.
We require further the following assumption to establish the asymptotic normality result, which concerns the existence of finite fourth moments.
Assumption 10 (Finite Fourth Moment).
The individual treatment effect estimator satisfies:
  1. (a)
    $\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)^{4}\right]<\infty$;
  2. (b)
    $\sup_{i\in\mathbb{N}}\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)^{4}\right]<\infty$.
Note that condition $(b)$ imposes a stronger requirement than $(a)$, as it demands uniform boundedness of the fourth moments across all units.
We now present the theorem establishing asymptotic normality of the Riesz estimator in the stochastic setting.
Theorem 4 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, and 10(a)(a), the aggregate Riesz estimator defined in (2.18) satisfies the following asymptotic normality
$\sigma_{n}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)\xrightarrow{d}\mathcal{N}(0,1).$ (3.7)
Both Assumptions 8(r=3) and 8(r=4) are required here, as boundedness at $r=4$ does not imply the same at $r=3$ over $(\mathcal{Z},\Omega)$, due to the potential unboundedness of the joint space.
Theorem 4 provides the basis for statistical inference using the aggregate Riesz estimator under outcome-level randomness. In particular, it justifies asymptotic testing and confidence interval construction for the null hypothesis $H_{0}:\tau_{n}=0$, based on the standardised statistic $T_{n}=\hat{\tau}_{n}(z,\omega)/\sigma_{n}$, assuming that both $\hat{\tau}_{n}(z,\omega)$ and $\sigma_{n}$ are consistently estimable from a single realisation of $z$ and $\omega$.
The required rate condition on the maximum dependency neighbourhood size, $D_{n}=o(n^{1/4})$, is in line with that of HWS. However, the proof here proceeds by bounding Wasserstein distance under mixed expectations over both the treatment assignments $z$ and latent heterogeneity $\omega$. This highlights a key observation: the dependence-degree requirement does not tighten despite the added randomness in potential outcomes.
This result is particularly relevant in empirical settings where potential outcomes are inherently random. For instance, in field experiments involving sensor-based measurements, observed outcomes may be contaminated by random noise due to hardware limitations or environmental fluctuations. Similarly, in panel experiments with repeated units, treatment effects may be estimated via machine learning algorithms that introduce randomness in the form of first-stage fitted values. In both cases, the random component of the outcome arises naturally and cannot be ignored. The asymptotic normality result ensures that inference remains valid under such randomness.
4 Variance Estimation under Local Dependence
The key message of this section is that, under local dependence, consistent variance estimation is possible without repeated experiments, provided that the dependence structure is sufficiently sparse and partially known. The test statistic supported by Theorem 4 requires an estimate of the variance of the aggregate Riesz estimator $\hat{\tau}_{n}$, which is generally unknown in finite samples. In the presence of local dependence, such as that arising from network interference, exposure mappings, or latent factors, standard i.i.d. variance estimators are invalid, and specialised methods must be employed.
Several approaches have been proposed. Some yield consistent estimators, but only under strong structural assumptions. For example, Aronow and Samii (2017) rely on known exposure mappings and joint inclusion probabilities; Yu et al. (2022) assume bounded-degree network structures; and Liu and Hudgens (2014) consider two-stage designs with bounded group sizes. Other methods construct variance bounds rather than consistent estimators: Chin (2019) derives conservative expressions using Stein’s method, while Chen and Shao (2004) propose block bootstrap procedures that are valid under fixed neighbourhood size and finite moment assumptions.
Of particular relevance to our setting, HWS construct an unbiased estimator for an upper bound on the variance using a tensor-product Riesz representation. More recently, Harshaw et al. (2024) develop a constrained optimisation framework to tighten such conservative bounds across a wide range of experimental designs. However, neither method formally proves that their bound converges to the true variance, except in special cases. Without sharpness, these bounds may overestimate the variance and lead to overly conservative inference.
In this paper, we take a different approach. Rather than estimating a variance upper bound, we propose a direct estimator for the true asymptotic variance of the aggregate Riesz estimator under local dependence. This estimator is shown to be consistent under conditions that essentially match those required for asymptotic normality. The key requirement is knowledge of the dependency neighbourhood structure: not only the maximal neighbourhood size $D_{n}$, but also which specific units are likely to be dependent. This may seem restrictive, but it reflects a fundamental limitation of single-realisation variance estimation.
Our recommendation is practical. The functional $\theta_{i}(\cdot)$ is known, and given $\psi_{i}(\cdot,\omega_{0})$ the quantity $\hat{\theta}_{i}(z,\omega_{0})$ is directly computable from the observed data via (2.14). In practice, $\psi_{i}(\cdot,\omega_{0})$ is approximated by $\hat{\psi}_{i}(\cdot,\omega_{0})$ constructed in the following section; all subsequent results are formulated in terms of $\psi_{i}$, and the approximation error induced by $\hat{\psi}_{i}$ is controlled separately. We allow for sharp null hypotheses under which $\theta_{i}(\tilde{y}_{i})$ is specified, and under such nulls, the centred quantity
$\zeta_{i}(z,\omega)\;=\;\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$ (4.1)
is well-defined and its realised value $\zeta_{i}(z,\omega_{0})$ is available for variance estimation. We therefore retain only those cross terms corresponding to known or assumed dependent pairs, setting all others to zero, which yields a feasible variance estimator that is consistent under local dependence and supports asymptotically valid inference.
Let $\vec{\zeta}_{n}=(\zeta_{1},\dots,\zeta_{n})^{\prime}$ denote a vector of centred random variables with covariance matrix $\Sigma_{n}=\mathbb{E}[\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}]$. From a single realisation, we observe only the outer product $\hat{\Sigma}_{n}=\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}$, and cannot estimate $\Sigma_{n}$ consistently without further structure. Even under independence, each $\vec{\zeta}_{n}\vec{\zeta}_{n}^{\prime}$ typically has variance of order one, so
$\mathbb{E}\bigl[\|\hat{\Sigma}_{n}-\Sigma_{n}\|_{F}^{2}\bigr]\;\geq\;cn^{2}$ (4.2)
for some constant $c>0$, where $\left\lVert\cdot\right\rVert_{F}$ stands for the Frobenius norm for a matrix, indicating that the Frobenius error diverges with $n$.
Consistency can be recovered by imposing stronger assumptions, for example: observing multiple independent copies of $\vec{\zeta}_{n}$, enabling standard sample covariance estimators; time ordering and weak stationarity, allowing the use of banded estimators or kernel-smoothed autocovariances; structural models such as factor structures (low-rank plus sparse decomposition), enabling regularised estimation (e.g., POET); or vanishing second-order signal, for instance when $\|\Sigma_{n}\|_{F}\to 0$ as $n\to\infty$; among others.
Our proposed approach sidesteps this by targeting only those components of the variance that correspond to known dependent pairs.
Suppose the dependency structure is given by
$\mathcal{E}_{n}=\{(i,j)\in\{1,\dots,n\}^{2}:\zeta_{i}\text{ and }\zeta_{j}\text{ are dependent}\}.$ (4.3)
We define the local-dependence variance estimator
$\hat{\sigma}_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j},$ (4.4)
and the population variance
$\sigma_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\mathbb{E}[\zeta_{i}\zeta_{j}].$ (4.5)
The following theorem shows that, under mild regularity conditions, $\hat{\sigma}_{n}^{2}$ consistently estimates $\sigma_{n}^{2}$.
Theorem 5 (Consistency of the Variance Estimator).
Under Assumptions 1, 35, and 10(b)(b), $n\hat{\sigma}_{n}^{2}-n\sigma_{n}^{2}=O_{p}(n^{-1/2}D_{n}^{3/2})$. In particular, under Assumption 6(d=1/3), $n\hat{\sigma}_{n}^{2}-n\sigma_{n}^{2}$ converges to zero in mean square .
Corollary 1 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, and 10(b)(b), the aggregate Riesz estimator defined in (2.18) satisfies the following asymptotic normality
$\hat{\sigma}_{n}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)\xrightarrow{d}\mathcal{N}(0,1).$ (4.6)
Note that Assumption 6(d=1/4) is stronger than Assumption 6(d=1/3) and therefore suffices to imply both consistency and asymptotic normality.
Corollary 1 establishes that the aggregate Riesz estimator is asymptotically normal under local dependence, provided that the dependency neighbourhoods are sufficiently sparse and the variance estimator $\hat{\sigma}_{n}^{2}$ is properly constructed.
A natural question is whether one can use a reduced dependency structure based solely on correlation. In many experimental settings, it is difficult to explicitly define or justify a full dependency neighbourhood structure $\mathcal{E}_{n}$ based on latent interference or unobserved design constraints. However, domain knowledge or structural assumptions may suggest which unit-level estimators $\zeta_{i}$ and $\zeta_{j}$ are likely to be uncorrelated, even if their full dependence is unknown. This motivates constructing a variance estimator by summing only over pairs believed to exhibit non-negligible second-order dependence, resulting in a conservative correlation-based estimator. Specifically, consider the set
$\mathcal{E}_{n}^{c}=\{(i,j)\in\{1,\dots,n\}^{2}:\mathbb{E}[\zeta_{i}\zeta_{j}]\neq 0\},$ (4.7)
and define the corresponding correlation-based variance estimator
$\hat{\sigma}_{cn}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}^{c}}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}.$ (4.8)
The answer is affirmative: using this reduced set still yields a consistent variance estimator under the same local dependence framework.
Theorem 6 (Consistency of the Variance Estimator).
Under Assumptions 1, 35, and 10(b)(b), $n\hat{\sigma}_{cn}^{2}-n\sigma_{n}^{2}=O_{p}(n^{-1/2}D_{n}^{3/2})$. In particular, under Assumption 6(d=1/3), $n\hat{\sigma}_{cn}^{2}-n\sigma_{n}^{2}$ converges to zero in mean square .
Corollary 2 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, and 10(b)(b), the aggregate Riesz estimator defined in (2.18) satisfies
$\hat{\sigma}_{cn}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)\xrightarrow{d}\mathcal{N}(0,1).$ (4.9)
Although the correlation-based variance estimator omits many cross-terms in the full variance formula, it yields valid inference when the omitted pairs correspond to approximately uncorrelated units. Importantly, the assumption that certain unit-level estimators $\zeta_{i}$ and $\zeta_{j}$ are uncorrelated cannot be directly verified from a single dataset. Nonetheless, this assumption often aligns with realistic structural knowledge available to practitioners, such as spatial adjacency, blocking in experimental design, or known network structure, which can be used to construct a conservative approximation of the correlation graph. In this way, the correlation-based approach offers a practically viable and interpretable alternative to full dependency modelling.
Although these variance estimators are developed within our stochastic setting, they are equally applicable to the fixed-outcome setting as a special case. Indeed, the classical framework can be recovered by taking the latent space $\Omega$ to be a singleton, in which case the only source of randomness is the treatment assignment. Under this simplification, the local dependence structure reduces to that of the assignment mechanism alone, typically known by design (see Assumption 2). Consequently, the variance estimators proposed here can be directly applied to previous frameworks such as HWS, and are in fact easier to implement in such settings due to the absence of outcome-level variation. This extends the practical utility of our estimators beyond the stochastic setting, providing consistent inference tools for a broad class of design-based estimators.
In the remainder of the paper, we adopt the correlation-based variance estimator $\hat{\sigma}_{cn}^{2}$, while emphasising that the overall methodology still relies on the local dependency assumption, even though it is not explicitly invoked in the variance formula. Crucially, while uncorrelatedness justifies omitting second-order cross terms in the estimator, it does not imply the vanishing of higher-order joint moments across units. Therefore, a corresponding correlation-based neighbourhood structure cannot be meaningfully defined for the full dependency graph.
In practice, this means we continue to assume a local dependency structure satisfying Assumption 6 with $d=1/4$, while approximating $\sigma_{n}^{2}$ by summing only over pairs $(i,j)$ for which $\zeta_{i}$ and $\zeta_{j}$ are correlated.
Suppose that, in practice, the correlation structure is approximated by a set of index pairs $\tilde{\mathcal{E}}_{n}\subset\{1,\dots,n\}^{2}$, serving as a practical substitute for (4.7). To ensure valid inference, the construction of $\tilde{\mathcal{E}}_{n}$ should be conservative, satisfying
Assumption 11 (Conservative Set of Index Pairs).
$\mathcal{E}_{n}^{c}\subset\tilde{\mathcal{E}}_{n}\subset\mathcal{E}_{n}$.
Based on this, we define a sparse matrix $\hat{\Sigma}_{n}^{d}\in\mathbb{R}^{n\times n}$ that retains only the entries corresponding to the index set $\tilde{\mathcal{E}}_{n}$, representing the estimated dependency structure.
[…]
The corresponding variance estimator is then given by
$\tilde{\sigma}_{n}^{2}\;=\;\nu_{n}^{\prime}\hat{\Sigma}_{n}^{d}\nu_{n},$ (4.11)
where $\nu_{n}=(\nu_{n1},\dots,\nu_{nn})^{\prime}$ is the vector of aggregation weights.
By appropriately selecting the conservative set of index pairs, the corresponding variance estimator remains consistent.
Theorem 7 (Consistency of the Variance Estimator).
Under Assumptions 1, 35, 10(b)(b), and $1$, $n\tilde{\sigma}_{n}^{2}-n\sigma_{n}^{2}=O_{p}(n^{-1/2}D_{n}^{3/2})$. In particular, under Assumption 6(d=1/3), $n\tilde{\sigma}_{n}^{2}-n\sigma_{n}^{2}$ converges to zero in mean square .
Corollary 3 (Asymptotic Normality).
Under Assumptions 1, 35, 6(d=1/4), 8(r=3), 8(r=4), 9, 10(b)(b), and $1$, the aggregate Riesz estimator defined in (2.18) satisfies
$\tilde{\sigma}_{n}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)\xrightarrow{d}\mathcal{N}(0,1).$ (4.12)
Note that consistency is preserved even if some independent pairs are mistakenly included in $\tilde{\mathcal{E}}_{n}$, provided that the total number of dependent pairs, including both genuinely dependent and erroneously included independent pairs, remains within the sparsity condition $D_{n}=o(n^{1/4})$.
In the special case where all units are believed to be mutually independent, i.e., each $\zeta_{i}$ is independent of every other $\zeta_{j}$, $\tilde{\mathcal{E}}_{n}$ consists only of diagonal pairs $(i,i)$. Then $\hat{\Sigma}_{n}^{d}$ reduces to a diagonal matrix, and the variance estimator and the true variances simplify to
$\tilde{\sigma}_{n}^{2}=\sum_{i=1}^{n}\nu_{ni}^{2}\zeta_{i}^{2},\quad\text{and}\quad\sigma_{n}^{2}=\sum_{i=1}^{n}\nu_{ni}^{2}\mathbb{E}[\zeta_{i}^{2}],$ (4.13)
respectively. This corresponds to the classical form used in standard inference under independence. Corollaries 1 and 2 thus recover the familiar asymptotic normality result as a special case of the more general locally dependent setting treated here.
From a functional perspective, the target variance $\sigma_{n}^{2}$ can be regarded as a continuous linear functional of a collection of second-order moments, specifically, $\Sigma_{n}$ or the covariances $\operatorname{Cov}(\zeta_{i},\zeta_{j})$ for pairs $(i,j)$, with norm $\sum_{i=1}^{n}\nu_{ni}^{2}$. Since these covariances are not directly observable from a single experimental realisation, our estimator proceeds by substituting the population moments with observed products $\zeta_{i}\zeta_{j}$, yielding the expression in (4.4). In this sense, the estimator operates analogously to a plug-in estimator: it replaces the unobservable moments with empirical counterparts derived from the data. However, unlike classical plug-in estimators that substitute a distributional parameter with an empirical measure, see for example van der Vaart (2000, Section 25.8), this construction relies on direct moment substitution guided by a known dependency graph. This perspective clarifies that the variance estimators approximates a functional over unknown moments, and its consistency is guaranteed by structural assumptions on the dependence neighbourhood and boundedness of higher-order terms.
Taken together, Theorems 37 and Corollaries 13 convey a central message: although the estimand $\tau_{n}$ is defined as an average of individual treatment effects evaluated over the joint probability space $\mathcal{Z}\times\Omega$, the entire inference procedure, from estimation to variance approximation, can be carried out using just a single realisation of $z\in\mathcal{Z}$ and $\omega\in\Omega$. This is made possible by the structure of the Riesz representation framework and the use of carefully constructed estimators that remain valid under local dependence. As the sample size $n$ grows, one draw from the treatment and outcome distribution contains sufficient information for reliable estimation and inference, provided that the underlying conditions hold. In this way, the results provide a rigorous justification for making population-level claims based on one observed experiment.
5 Riesz Representer Estimation in Practice
The theoretical results in earlier sections show that the aggregate Riesz estimator in the stochastic setting can be constructed using only a single realisation of the underlying data-generating process. Specifically, the Riesz estimates are computed over the randomisation distribution of $z$, conditional on a fixed (but unobserved) $\omega_{0}$. Crucially, although the representers are constructed conditionally on $\omega_{0}$, the target estimand remains identifiable because aggregation across units recovers the expectation over $\omega$, as established in previous sections.
This structure has concrete implications for how we estimate the individual Riesz representers $\psi_{i}(z,\omega)$. By Theorem 1, for each unit $i$, there exists a unique representer satisfying (2.12) with (2.6). Hence, the representer $\psi_{i}(z,\omega_{0})$ is identified conditionally and can be characterised by the realised potential outcome function $\tilde{y}_{i}(\cdot,\omega_{0})$.
The setting considered by HWS arises as a special case of our framework in which the latent space collapses to a singleton, $\Omega=\{\omega_{0}\}$, so that potential outcomes reduce to deterministic functions of the treatment assignment.
5.1 Computation in Finite-Dimensional Model Spaces
In finite-dimensional model spaces $\mathcal{M}_{i}$, we can estimate $\psi_{i}(z,\omega_{0})$ via basis expansion and moment matching. Let $\{g_{i,1},\dots,g_{i,m}\}$ be a basis for $\mathcal{M}_{i}$ and $\mathrm{dim}(\mathcal{M}_{i})=m$. We solve for coefficients $\bm{\beta}_{i}\in\mathbb{R}^{m}$ such that
$\hat{\psi}_{i}(z,\omega_{0})=\sum_{k=1}^{m}\beta_{i,k}\,g_{i,k}(z,\omega_{0})$ (5.1)
satisfies the identity $\theta_{i}(\tilde{y}_{i})=\langle\tilde{y}_{i},\hat{\psi}_{i}\rangle^{0}$. Algorithm 1 summarises this procedure. This construction aligns with the approach described by HWS (preprint version), where Riesz representers are computed via basis expansion and moment matching under the randomisation distribution.
Throughout the construction of the Riesz representer, all objects are understood conditionally on the realised latent state $\omega_{0}$. Accordingly, basis functions are evaluated at $\omega_{0}$ and can be treated as functions of the treatment assignment $z$ alone. Formally, this corresponds to working on the fibre $\mathcal{M}_{i}(\omega_{0})\subset L^{2}(\mathcal{Z})$ induced by the Bochner space representation of random potential outcomes.
When the treatment space $\mathcal{Z}$ is finite with $K$ distinct levels, a simple and interpretable basis consists of treatment indicators. In this case, each basis function is defined as a function of $z$ that equals one if $z$ corresponds to the $k$-th treatment condition and zero otherwise. This yields a saturated approximation space with a separate coefficient for each treatment arm and does not rely on any structural assumptions about the response surface.
When $z$ is continuous or high-dimensional, implementable approximation spaces are typically constructed using low-order basis functions of $z$, such as polynomials or splines. These bases offer a balance between expressiveness and computational tractability and can be tailored to the structure of the experimental design. Although the underlying potential outcome functions may depend jointly on $z$ and the latent environment $\omega$, this dependence is fixed at $\omega_{0}$ during the representer construction stage and is therefore absorbed into the realised section of the function space.
This procedure is fully offline in the sense that it depends only on the design distribution for $z$, the basis functions, and the known functional $\theta_{i}$. It does not require access to observed outcomes $Y_{i}=\tilde{y}_{i}(z_{0},\omega_{0})$.
This construction relies on Assumption 2, which ensures that the treatment assignment $z$ is independent of the latent variable $\omega$. In particular, this allows expectations over $z$ conditional on a fixed realisation $\omega_{0}$ to be taken with respect to the known design distribution $\mu$. Without this assumption, the conditional distribution of $z\mid\omega_{0}$ could differ from $\mu$, and the estimated representer would no longer correspond to the true Riesz representation defined under the design. It is at this stage, during the offline construction of $\hat{\psi}_{i}(z,\omega_{0})$, that Assumption 2 is invoked. Independence between the treatment assignment $z$ and the latent variable $\omega$ is required solely to identify the Riesz representer under the design. Once the representer has been constructed, all subsequent inference relies only on the realised data and does not require independence between treatment assignment and outcome-level randomness.
In symmetric settings with common $\mathcal{M}_{i}$ and $\theta_{i}$, the same representer $\psi$ can be reused across units.
5.2 Approximation in Infinite-Dimensional Model Spaces
In many practical applications, the model space $\mathcal{M}_{i}$ is infinite-dimensional. This arises naturally when either the treatment assignment $z$ or the latent variable $\omega$ lies in a continuous domain, or when potential outcomes depend on complex functional relationships. To compute the Riesz representer in such settings, we project the problem onto a sequence of growing finite-dimensional subspaces.
Our setting explicitly accounts for a random latent variable $\omega$, and the Riesz representation is defined with respect to an inner product $\langle\cdot,\cdot\rangle^{0}$ taken over the treatment assignment $z$, conditional on the fixed value $\omega_{0}\in\Omega$. Although our estimation routine shares a computational structure with existing methods, constructing a Gram matrix and solving a linear system, our framework generalises the representer computation to a richer class of problems involving random potential outcomes.
Assumption 12 (Separable Model Space).
The model space $\mathcal{M}_{i}\subset L^{2}(\mathcal{Z}\times\Omega)$, equipped with the inner product $\langle\cdot,\cdot\rangle^{0}$ defined in (2.6) for some $\omega_{0}\in\Omega$, is a separable Hilbert space.
By Assumption 12, the space $\mathcal{M}_{i}$ admits a countable orthonormal basis $\{e_{k}\}_{k\geq 1}$. For each dimension $m\in\mathbb{N}$, define the subspace
$\mathcal{M}_{i}^{(m)}=\operatorname{span}\{e_{1},\dots,e_{m}\},$ (5.2)
and let $P_{m}:\mathcal{M}_{i}\to\mathcal{M}_{i}^{(m)}$ denote the orthogonal projection. Then the projected representer
$\psi_{i}^{(m)}:=P_{m}\psi_{i}=\sum_{k=1}^{m}\langle\psi_{i},e_{k}\rangle^{0}\,e_{k}$ (5.3)
converges to $\psi_{i}$ in norm as $m\to\infty$, by the completeness of the Hilbert space $\mathcal{M}_{i}$, since $\{e_{k}\}_{k\geq 1}$ forms a total orthonormal system and $P_{m}$ is the orthogonal projection onto the span of the first $m$ basis elements.
The estimation of $\psi_{i}^{(m)}$ proceeds by solving the finite-dimensional moment equation system that results from projecting the Riesz representation identity onto the span of the first $m$ basis functions. This yields a linear system involving a Gram matrix and a target vector, as shown in Algorithm 2.
This procedure is still fully offline in the sense that both $G_{i}^{(m)}$ and $T_{i}^{(m)}$ can be computed offline from the design distribution and the known linear functional $\theta_{i}$. It does not require access to observed outcomes $Y_{i}=\tilde{y}_{i}(z_{0},\omega_{0})$. In symmetric settings with common $\mathcal{M}_{i}$ and $\theta_{i}$, the same representer $\psi$ can be reused across units.
The procedure for estimating $\psi_{i}(z,\omega_{0})$ described above relies on projecting onto a finite-dimensional subspace spanned by a pre-specified basis and solving empirical moment equations derived from the known randomisation design and linear functional $\theta_{i}$. This approach follows the classical sieve estimation framework and is theoretically well understood; see, e.g., Chen (2007); Chen and Pouzo (2012); van der Vaart (2000).
While consistency of $\hat{\psi}_{i}^{(m)}\to\psi_{i}$ under appropriate regularity conditions can be rigorously established using a standard bias-variance decomposition, this is not the focus of the present paper. Our main contribution lies in the theory developed in earlier sections, namely, the design-based asymptotic normality of the aggregate estimator and the construction of consistent variance estimators under local dependence.
For completeness, we note that under standard conditions, such as an orthonormal basis, eigenvalues of the Gram matrix bounded away from zero, and sufficient smoothness of the representer, the estimation error satisfies
$\|\widehat{\psi}_{i}^{(m)}-\psi_{i}\|=O_{p}\left(m^{-s}+\sqrt{m/n}\right)$ (5.4)
for some $s>0$. This rate is minimised by setting the truncation level $m$ proportional to $n^{1/(2s+1)}$, which balances the bias and variance contributions. We omit the full proof, as it follows standard arguments from the sieve estimation literature.
The orthonormal system $\{e_{k}\}$ can either be specified a priori, using classical bases such as Fourier series, wavelets, or splines, or derived from the data through methods like principal components or kernel eigenfunctions. Additional practical examples of basis construction in deterministic settings are discussed in the preprint by HWS, to which we refer the interested reader for further details.
6 Simulation Study
To illustrate the finite-sample behaviour, we carry out a simulation study to evaluate the finite-sample performance of the Riesz estimator and the variance estimators. The simulation is designed to reflect a typical experimental setting with local dependence induced by latent block-level effects.
For each unit $i\in\{1,\dots,n\}$, we simulate treatment assignment $z_{i}\sim\mathrm{Bernoulli}(0.5)$ independently. Covariates $x_{i}$ and outcome-level parameters $(\alpha_{i},\delta_{i})$ are drawn i.i.d. from $\mathcal{N}(0,1)$, while $\beta_{i}$, representing the unit-level treatment effect, is drawn i.i.d. from $\mathcal{N}(1,1)$. This specification induces a non-zero average treatment effect, facilitating power assessment in finite samples.
The potential outcome is generated via
$\tilde{y}_{i}(z,\omega)=\tilde{y}_{i}(z_{i})=\alpha_{i}+\beta_{i}z_{i}+\delta_{i}x_{i}+\varepsilon_{i},$ (6.1)
where $\varepsilon_{i}$ is a locally dependent noise term described below. The observed outcome is $Y_{i}=\tilde{y}_{i}(z,\omega)$.
Local dependence is introduced by partitioning the sample into blocks, where units in the same block share a common latent shock. Specifically, the number of blocks is $B_{n}=\lfloor n^{1-d}\rfloor$ for a fixed parameter $d\in\{0,0.1,0.2,0.25,0.3\}$. The error for unit $i$ is
$\varepsilon_{i}=\gamma_{i}\,\eta_{b(i)}+\nu_{i},$ (6.2)
where $b(i)$ denotes the block index to which unit $i$ belongs, $\eta_{b(i)}\sim\mathcal{N}(0,1)$ is the shared block-level shock, $\nu_{i}\sim\mathcal{N}(0,1)$ is an idiosyncratic noise term, and $\gamma_{i}\in\{-1,1\}$ is a random sign (independent Rademacher) to allow both positive and negative correlations between units in the same block. This construction yields a dependency graph $\mathcal{E}_{n}$ with within-block dependence and across-block independence.
In this simulation, we set $\theta_{i}(\tilde{y}_{i}):=\tilde{y}_{i}(1)-\tilde{y}_{i}(0)$. Inference is conducted under sharp null hypotheses that specify these values unit by unit; in particular, for the size experiments we generate data under the sharp null $\theta_{i}(\tilde{y}_{i})\equiv 0$.
We estimate the average treatment effect using the aggregate Riesz estimator. The representer $\psi_{i}$ is computed as
$\psi_{i}(z,\omega_{0})=\frac{z}{\mu_{1}}-\frac{1-z}{\mu_{0}},$
with $\mu_{1}=\mathbb{E}[z_{i}]=0.5$, $\mu_{0}=1-\mu_{1}$, which corresponds to the efficient estimator under the known design. This choice corresponds to the efficient influence function for the average treatment effect under Bernoulli randomisation and is standard in the design-based literature; see, for example, Neyman (1990), Imbens and Rubin (2015), and Aronow and Samii (2017). In the simulation, we work with the true representer $\psi_{i}$ to isolate the impact of local dependence and variance estimation, as the construction and estimation of $\psi_{i}$ in more general settings is not the focus of this paper. The estimator is
$\hat{\tau}_{n}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}\psi_{i}(z_{i},\omega_{0}),$
and the variance estimator follows (4.4)
$\hat{\sigma}_{n}^{2}=\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j},\quad\text{where }\zeta_{i}=Y_{i}\psi_{i}(z_{i})-\theta_{i}(\tilde{y}_{i})\text{ and }\nu_{ni}=1/n.$
That is, we sum pairwise products of residuals over all unit pairs within the same block.
In this simulation setting, the variance estimators $\hat{\sigma}_{cn}^{2}$ and $\tilde{\sigma}_{n}^{2}$ coincide with the local-dependence estimator $\hat{\sigma}_{n}^{2}$. This is because the dependency structure is fully captured by shared block-level shocks, and all non-zero cross-moments correspond to block-sharing units. Therefore, all simulation results reflect the performance of both estimators. In general, however, the correlation-based estimator yields improved efficiency in settings where weakly dependent pairs can be excluded without loss of accuracy.
We consider sample sizes $n\in\{100,200,500,1000\}$ and dependency growth parameters $d\in\{0.0,0.1,0.2,0.25,0.3\}$. For each $(n,d)$ pair, we run $2000$ independent replications. For each replication, we compute the estimator $\hat{\tau}_{n}$ and the standard error $\hat{\sigma}_{n}$, and conduct inference under two specifications of the centring term: the correctly specified sharp null $\theta_{i}(\tilde{y}_{i})$ and the misspecified null $\theta_{i}(\tilde{y}_{i})\equiv 0$. We record the following metrics:
  • Coverage of 95% confidence intervals: $\mathbb{P}(|\hat{\tau}_{n}-\tau_{n}|\leq 1.96\hat{\sigma}_{n})$;
  • Rejection rates at 1%, 5%, and 10% significance levels.
The complete codebase, including replication scripts and plotting routines, is publicly available at github.com/yukai-yang/RieszRE_Experiments under the MIT license.
Figure 1 reports the empirical coverage of nominal 95% confidence intervals across different dependency growth rates $d$ and sample sizes $n$. The dependency parameter increases from left to right, with $d\in\{0,0.1,0.2,0.25,0.3\}$, where the largest value $d=0.3$ lies outside the scope of Assumption 6 (d=1/4) in Theorem 4.
Across all configurations, empirical coverage is slightly above the nominal level of 0.95, indicating mildly conservative inference. For smaller sample sizes such as $n=100$ and $n=200$, no clear monotone pattern with respect to $d$ is visible, which is consistent with higher sampling variability in small samples. As the sample size increases, particularly for $n=1000$, a clearer pattern emerges: coverage increases gradually with the degree of local dependence. This behaviour reflects the increasing conservativeness of the variance estimator as dependence intensifies.
Importantly, when the dependency growth rate satisfies the conditions of Theorem 4, namely $d\leq 1/4$, coverage remains close to the nominal level and is uniformly satisfactory. Overall, these results confirm that the proposed variance estimator delivers well-calibrated confidence intervals under local dependence, with a modest degree of conservativeness that becomes more pronounced as dependence strengthens.
Table 2 reports empirical rejection frequencies at the 1%, 5%, and 10% nominal levels for both the size and power experiments, across all combinations of sample size $n$ and dependency growth rate $d$.
In the size experiment, rejection rates are generally close to their nominal levels across all sample sizes and dependency regimes considered. At the 5% level, empirical rejection frequencies typically fall in the range of approximately 3.6% to 5%, with no evidence of systematic size distortion as $n$ increases. Similar behaviour is observed at the 1% and 10% levels, indicating that the test maintains appropriate size control under local dependence when the conditions of Corollary 1 are satisfied.
Moreover, except for the case $d=0.3$, empirical rejection frequencies tend to lie slightly below their nominal levels. This indicates that the proposed test is mildly conservative within the theoretical regime. This behaviour is fully consistent with the coverage results reported in Figure 1, where confidence intervals slightly over-cover the nominal level. Together, these findings suggest that the conservativeness arises primarily from variance estimation rather than from any systematic distortion of the estimator itself.
Mild over-rejection is observed for larger values of $d$, particularly when $d=0.3$, which lies outside the theoretical range required by Assumption 6(d=1/4)(d=1/4). This pattern is consistent with the coverage results and suggests a gradual loss of calibration as the dependence neighbourhood grows faster than permitted by the theory.
In the power experiment, rejection frequencies increase rapidly with the sample size, approaching one even for moderate $n$. For $n\geq 500$, rejection rates are essentially one across all significance levels and dependency regimes, reflecting the strong signal embedded in the data-generating process. These results confirm that the proposed procedure retains substantial power while maintaining accurate size control within the theoretical regime.
Overall, the simulation results are consistent with the theoretical analysis. Within the dependency regimes covered by the theory, the aggregate Riesz estimator admits reliable inference under local dependence, with finite-sample behaviour consistent with the asymptotic results.
Although the simulation design is based on a parametric data-generating process, the proposed estimation and testing procedures do not exploit any model structure. All inference is conducted in a fully design-based manner, relying solely on the randomisation and dependence structure. The estimand itself is defined as an expectation over random potential outcomes, rather than as a fixed unit-level quantity. This illustrates that valid inference can be achieved from a single realised experiment under local dependence, without outcome modelling, and highlights the practical relevance of the proposed framework.
7 Conclusion
This paper extends the classical design-based framework for causal inference to settings in which potential outcomes are random. By allowing outcomes to depend on a latent random variable $\omega$, we accommodate outcome-level randomness while preserving the identifying role of randomised treatment assignment. In this formulation, the target estimands are defined as expectations over random potential outcomes, rather than as fixed unit-level quantities. This generalises the Neyman–Rubin framework without introducing outcome models or parametric assumptions.
Within this stochastic setting, we study the large-sample behaviour of the aggregate Riesz estimator under local dependence. The theory allows for growing and heterogeneous dependence structures and does not rely on convergence of the estimator or its variance to fixed limits. Instead, inference is conducted directly for the finite-sample target $\tau_{n}$, even when the variance sequence itself does not stabilise. This feature makes the framework well suited to experimental environments with network structure or non-standard dependence.
A key contribution of the paper is the development of consistent and feasible variance estimators that explicitly account for local dependence. Although motivated by random potential outcomes, these estimators also apply to the classical fixed-outcome setting as a special case. In such designs, the dependence structure is typically known in advance, and the proposed estimators provide a practical alternative to conservative variance bounds that have been used in earlier work.
The main theoretical results establish consistency and asymptotic normality of the standardised Riesz estimator under mild moment and neighbourhood growth conditions. Notably, valid inference is obtained from a single realised experiment, without relying on outcome modelling or repeated sampling of the stochastic environment. Under local dependence, averaging across units plays the role traditionally assigned to averaging across repeated experiments.
Simulation results support the theoretical analysis. Across a range of sample sizes and dependence regimes, confidence intervals exhibit accurate or slightly conservative coverage, and rejection frequencies remain close to nominal levels within the theoretical range of dependence growth. When the dependence neighbourhood grows beyond this range, calibration deteriorates gradually, in line with the limits imposed by the assumptions. At the same time, power increases rapidly with sample size, indicating that robustness to dependence does not come at the expense of efficiency.
Taken together, these results show that the Riesz representation approach remains effective for design-based inference in the presence of outcome-level randomness and structured dependence. The framework is particularly relevant for modern experiments in which stochastic variation is intrinsic rather than incidental. Possible extensions include non-smooth functionals, dynamic treatments, and observational designs, while maintaining the central principle that valid inference should rest on the randomisation and dependence structure rather than on outcome models.
Acknowledgements
I am grateful to Fredrik Sävje for his valuable discussions and support.
Appendix A Notation Summary
Table 1: Summary of commonly used notation throughout the paper.
Symbol Description
$n$ Number of units in the experiment
$z\in\mathcal{Z}$ Treatment assignment vector
$\mu$ Probability measure for randomisation
$\omega\in\Omega$ Latent variable representing outcome-level randomness
$P$ Probability measure for the latent random variable $\omega$
$\mathcal{F}_{x}$ Sigma-algebra generated by the random variable $x$
$\tilde{y}_{i}(z,\omega)$ Potential outcome of unit $i$ under treatment $z$ and latent variable $\omega$
$\theta_{i}(\tilde{y}_{i})$ Linear functional defining the treatment effect for unit $i$
$\hat{\theta}_{i}(z,\omega)$ Riesz estimator for the treatment effect for unit $i$
$\zeta_{i}$ the difference $\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$
$\tau_{n}$ Aggregate estimand (e.g., average treatment effect)
$\hat{\tau}_{n}$ Aggregate Riesz estimator for $\tau_{n}$
$\psi_{i}(z,\omega)$ Riesz representer for unit $i$
$\hat{\psi}_{i}(z,\omega)$ Estimated Riesz representer (e.g., via basis expansion)
$\mathcal{M}_{i}$ Model space for unit $i$’s outcome function
$N_{i}$ Dependency neighbourhood of unit $i$
$D_{n}$ Maximum neighbourhood size across units
$d_{n}$ Average neighbourhood size across units
$\sigma_{n}^{2}$ Variance of $\hat{\tau}_{n}$
$\mathcal{E}_{n}$ set of dependent pairs $(i,j)$ for $\zeta_{i}$ and $\zeta_{j}$
$\hat{\sigma}_{n}^{2}$ Variance estimator of $\hat{\tau}_{n}$ using $\mathcal{E}_{n}$
$\mathcal{E}_{n}^{c}$ set of correlated pairs $(i,j)$ for $\zeta_{i}$ and $\zeta_{j}$
$\hat{\sigma}_{cn}^{2}$ Variance estimator of $\hat{\tau}_{n}$ using $\mathcal{E}_{n}^{c}$
$\tilde{\mathcal{E}}_{n}$ conservative set of index pairs
$\tilde{\sigma}_{n}^{2}$ Variance estimator of $\hat{\tau}_{n}$ using $\tilde{\mathcal{E}}_{n}$
Appendix B Algorithms
Algorithm 1 Computing the Riesz representer $\psi_{i}(z,\omega_{0})$, following the procedure of HWS
Inputs:
$\bullet$ Basis functions $\{g_{i,1},\dots,g_{i,m}\}$ for $\mathcal{M}_{i}$
$\bullet$ Known linear functional $\theta_{i}$
$\bullet$ Randomisation distribution for $z$
Step 1: Compute Gram Matrix $\hat{\mathbf{G}}_{i}$
$\hat{\mathbf{G}}_{i}]_{\ell,k}=\mathbb{E}_{z}\bigl[g_{i,\ell}(z,\omega_{0})g_{i,k}(z,\omega_{0})\bigr]$
Step 2: Compute Target Vector $\hat{\mathbf{T}}_{i}$
$\hat{\mathbf{T}}_{i}]_{\ell}=\theta_{i}(g_{i,\ell})$
Step 3: Solve Linear System
$\hat{\mathbf{G}}_{i}\bm{\beta}_{i}=\hat{\mathbf{T}}_{i}\quad\Rightarrow\quad\bm{\beta}_{i}=\hat{\mathbf{G}}_{i}^{-1}\hat{\mathbf{T}}_{i}$
Output: Estimated Riesz representer $\hat{\psi}_{i}(z,\omega_{0})$
Algorithm 2 Approximating the Riesz representer $\psi_{i}(z,\omega_{0})$ in infinite-dimensional settings
Inputs:
$\bullet$ Truncated orthonormal basis $\{e_{1},\dots,e_{m}\}$ of $\mathcal{M}_{i}$
$\bullet$ Known linear functional $\theta_{i}$
$\bullet$ Randomisation distribution for $z$
Step 1: Compute Gram Matrix $\hat{\mathbf{G}}_{i}^{(m)}$
$\hat{\mathbf{G}}_{i}^{(m)}]_{\ell,k}=\mathbb{E}_{z}\bigl[e_{\ell}(z,\omega_{0})\,e_{k}(z,\omega_{0})\bigr]$
Step 2: Compute Target Vector $\hat{\mathbf{T}}_{i}^{(m)}$
$\hat{\mathbf{T}}_{i}^{(m)}]_{\ell}=\theta_{i}(e_{\ell})$
Step 3: Solve Regularised Linear System
$\hat{\mathbf{G}}_{i}^{(m)}\bm{\beta}_{i}=\hat{\mathbf{T}}_{i}^{(m)}\quad\Rightarrow\quad\bm{\beta}_{i}=\left(\hat{\mathbf{G}}_{i}^{(m)}\right)^{-1}\hat{\mathbf{T}}_{i}^{(m)}$
Output: Estimated Riesz representer
$\hat{\psi}_{i}^{(m)}(z,\omega_{0})=\sum_{k=1}^{m}\beta_{i,k}\,e_{k}(z,\omega_{0})$
Appendix C Figures and Table
Figure 1: Coverage of 95% confidence intervals under varying dependency growth rates $d$.
0.920.940.960.9800.10.20.250.3Dependency growth rate ddn=n=1002005001000
Table 2: Empirical rejection rates under local dependence. The left panel reports size experiments under the correctly specified sharp null, while the right panel reports power experiments under heterogeneous treatment effects. Results are based on 2000 replications.
Size experiment Power experiment
level $d$ $n=100$ $200$ $500$ $1000$ $n=100$ $200$ $500$ $1000$
1% 0.00 0.0095 0.0090 0.0070 0.0085 0.340 0.730 0.997 1.000
0.10 0.0090 0.0065 0.0085 0.0110 0.335 0.741 0.994 1.000
0.20 0.0050 0.0065 0.0070 0.0075 0.314 0.705 0.991 1.000
0.25 0.0050 0.0095 0.0060 0.0050 0.285 0.710 0.990 1.000
0.30 0.0060 0.0065 0.0055 0.0055 0.268 0.690 0.992 1.000
5% 0.00 0.0445 0.0395 0.0415 0.0480 0.619 0.894 1.000 1.000
0.10 0.0425 0.0360 0.0410 0.0435 0.618 0.902 1.000 1.000
0.20 0.0400 0.0465 0.0460 0.0435 0.600 0.892 0.999 1.000
0.25 0.0415 0.0420 0.0450 0.0420 0.597 0.894 0.998 1.000
0.30 0.0495 0.0415 0.0400 0.0360 0.572 0.885 0.998 1.000
10% 0.00 0.0865 0.0905 0.0890 0.0935 0.744 0.942 1.000 1.000
0.10 0.0975 0.0875 0.0845 0.0875 0.740 0.948 1.000 1.000
0.20 0.0860 0.1040 0.0890 0.0850 0.714 0.938 1.000 1.000
0.25 0.0950 0.0960 0.0950 0.0905 0.732 0.942 1.000 1.000
0.30 0.0980 0.0905 0.0940 0.0870 0.719 0.944 1.000 1.000
Appendix D Lemmas and Proofs
Proof of Proposition 1.
The linearity for both the two examples follows directly from the facts that integrals are linear and that we are taking differences of linear functionals.
For continuity, we need to rewrite (2.8) clearly
$\theta_{i}(u)=\iint_{\mathcal{Z}\times\Omega}\left(\frac{\mathbf{1}_{A}(z)}{\mu(A)}-\frac{\mathbf{1}_{B}(z)}{\mu(B)}\right)u(z,\omega)\,\mu(\mathrm{d}z)\,P(\mathrm{d}\omega)=\left\langle u,\,\varphi\right\rangle,$
and
$\varphi(z,\omega)=\frac{\mathbf{1}_{A}(z)}{\mu(A)}-\frac{\mathbf{1}_{B}(z)}{\mu(B)}.$
We can see that $\varphi\in L^{2}(\mathcal{Z}\times\Omega)$ as
$\|\varphi\|_{L^{2}}^{2}$ $=$ $\int_{\Omega}\int_{\mathcal{Z}}\left|\frac{\mathbf{1}_{A}(z)}{\mu(A)}-\frac{\mathbf{1}_{B}(z)}{\mu(B)}\right|^{2}\,\mu(\mathop{}\!\mathrm{d}z)\,P(\mathop{}\!\mathrm{d}\omega)$
$=$ $\int_{\mathcal{Z}}\left|\frac{\mathbf{1}_{A}(z)}{\mu(A)}-\frac{\mathbf{1}_{B}(z)}{\mu(B)}\right|^{2}\,\mu(\mathrm{d}z)=\frac{1}{\mu(A)}+\frac{1}{\mu(B)}<\infty.$
Now continuity is immediate due to the Cauchy–Schwarz inequality
$|\theta_{i}(u)|=|\langle u,\varphi\rangle|\leq\left\lVert u\right\rVert\,\left\lVert\varphi\right\rVert.$
Proof of Proposition 2.
Linearity follows from the linearity of differentiation and expectation.
Since $s>1+\frac{d}{2}$, the Sobolev embedding theorem yields a continuous embedding $H^{s}(\mathcal{Z})\hookrightarrow C^{1}(\overline{\mathcal{Z}})$; see, e.g., Adams and Fournier (2003, Chapter 5). Hence, for almost every $\omega\in\Omega$, the function $z\mapsto u(z,\omega)$ admits a $C^{1}(\overline{\mathcal{Z}})$ representative, so that $\partial_{z}u(z_{0},\omega)$ is well-defined. Moreover, by continuity of the embedding, there exists a constant $C>0$ such that
$\bigl|\partial_{z}u(z_{0},\omega)\bigr|\leq C\|u(\cdot,\omega)\|_{H^{s}(\mathcal{Z})}\quad\text{for a.e. }\omega.$
Therefore,
$|\theta_{i}(u)|$ $=\left|\mathbb{E}_{\omega}\!\left[\partial_{z}u(z_{0},\omega)\right]\right|\leq\mathbb{E}_{\omega}\!\left[\bigl|\partial_{z}u(z_{0},\omega)\bigr|\right]$
$\leq C\,\mathbb{E}_{\omega}\!\left[\|u(\cdot,\omega)\|_{H^{s}(\mathcal{Z})}\right]\leq C\Bigl(\mathbb{E}_{\omega}\|u(\cdot,\omega)\|_{H^{s}(\mathcal{Z})}^{2}\Bigr)^{1/2}=C\|u\|_{L^{2}(\Omega;H^{s}(\mathcal{Z}))},$
where the last inequality is Cauchy–Schwarz. This proves continuity. ∎
Proof of Proposition 3.
Linearity is immediate. For a.e. $\omega$, $\partial_{z}u(\cdot,\omega)\in L^{2}(\mathcal{Z})$. Hence the integral is well-defined. By Cauchy–Schwarz,
$\left|\int_{\mathcal{Z}}\partial_{z}u(z,\omega)\,\varphi(z)\,dz\right|\leq\|\partial_{z}u(\cdot,\omega)\|_{L^{2}(\mathcal{Z})}\|\varphi\|_{L^{2}(\mathcal{Z})}.$
Taking expectations and applying Cauchy–Schwarz in $\omega$ yields
$|\theta_{i}(u)|\leq\|\varphi\|_{L^{2}(\mathcal{Z})}\Bigl(\mathbb{E}\|\partial_{z}u(\cdot,\omega)\|_{L^{2}(\mathcal{Z})}^{2}\Bigr)^{1/2}\leq\|\varphi\|_{L^{2}(\mathcal{Z})}\,\|u\|_{L^{2}(\Omega;H^{1}(\mathcal{Z}))}.$
Proof of Theorem 1.
See Riesz (1907). ∎
Proof of Theorem 2.
By definition of the Riesz estimator in (2.14) and the linearity of expectation, we have
$\mathbb{E}\bigl[\hat{\theta}_{i}(z,\omega)\bigr]=\theta_{i}(\tilde{y}_{i}),$
which further implies
$\mathbb{E}\bigl[\hat{\tau}_{n}(z,\omega)\bigr]=\sum_{i=1}^{n}\nu_{ni}\,\mathbb{E}\bigl[\hat{\theta}_{i}(z,\omega)\bigr]=\sum_{i=1}^{n}\nu_{ni}\,\theta_{i}(\tilde{y}_{i})=\tau_{n}.$
This equality follows from the construction of the estimator and does not require any integrability assumptions. In particular, it holds even if $\tilde{y}_{i}(z,\omega)\,\psi_{i}(z,\omega)\notin L^{1}$. ∎
Lemma 1.
Suppose the weights $\nu_{ni}$ satisfy Assumption 5. Then for any integer $r\geq 1$, the following bound holds:
$\sum_{i=1}^{n}\nu_{ni}^{r}=O(n^{1-r}).$ (D.1)
Proof.
By Assumption 5, we have
$\sum_{i=1}^{n}\nu_{ni}^{r}\leq\sum_{i=1}^{n}\left(\frac{\bar{\nu}}{n}\right)^{r}=n\cdot\left(\frac{\bar{\nu}}{n}\right)^{r}=\bar{\nu}^{r}\cdot n^{1-r},$
which implies the claimed order bound. ∎
Lemma 2.
Under Assumptions 1, 3, and 4, the $r$th central moment of the treatment effect Riesz estimator, for some integer $r\geq 1$, if it exists, satisfies
$\mathbb{E}\left|\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right|^{r}\leq\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p}\,\left\lVert\psi_{i}\right\rVert_{q}\right)^{r},$ (D.2)
for any $p,q\in[1,\infty]$ satisfying $1/p+1/q=1/r$. In particular, when $r=2$,
$\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{2}\leq\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2}.$ (D.3)
The inequalities (D.2) and (D.3) hold trivially if either $\|\tilde{y}_{i}\|_{p}$ or $\|\psi_{i}\|_{q}$ has no finite upper bound.
Proof.
Inequality (D.2) is a direct extension of Lemma C.3 in HWS, which handles the case without the additional randomness $\omega$. The proof carries through immediately by applying Hölder’s inequality on the product space $\mathcal{Z}\times\Omega$, since the additional integration over $\omega$ preserves the inequality structure.
Regarding inequality (D.3),
$\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{2}$ $=\mathbb{E}[\hat{\theta}_{i}(z,\omega)]^{2}-(\theta_{i}(\tilde{y}_{i}))^{2}\leq\mathbb{E}\left[\tilde{y}_{i}(z,\omega)\psi_{i}(z,\omega)\right]^{2}$
$=\left\lVert\tilde{y}_{i}(z,\omega)\psi_{i}(z,\omega)\right\rVert^{2}\leq\left(\left\lVert\tilde{y}_{i}(z,\omega)\right\rVert_{p}\,\left\lVert\psi_{i}(z,\omega)\right\rVert_{q}\right)^{2}.$ (Hölder)
Proof of Proposition 4.
Let $\zeta_{i}=\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$ and we observe that $\mathbb{E}[\zeta_{i}]=0$ by Theorem 2. Then,
$\sigma_{n}^{2}$ $=\mathbb{V}[\hat{\tau}_{n}(z,\omega)-\tau_{n}]=\mathbb{V}\left[\sum_{i=1}^{n}\nu_{ni}\zeta_{i}\right]=\mathbb{E}\left[\sum_{i=1}^{n}\nu_{ni}\zeta_{i}\right]^{2}$
$=\sum_{i=1}^{n}\sum_{j=1}^{n}\nu_{ni}\nu_{nj}\mathbf{1}(i\in N_{j})\mathbf{1}(j\in N_{i})\mathbb{E}\left[\zeta_{i}\zeta_{j}\right]$
$\leq\sum_{i=1}^{n}\sum_{j=1}^{n}\nu_{ni}\nu_{nj}\mathbf{1}(i\in N_{j})\mathbf{1}(j\in N_{i})\left\lVert\zeta_{i}\right\rVert\,\left\lVert\zeta_{j}\right\rVert$ (Cauchy–Schwartz)
$\leq\sum_{i=1}^{n}\nu_{ni}\sum_{j=1}^{n}\frac{1}{2}\nu_{nj}\left(\mathbf{1}(j\in N_{i})\left\lVert\zeta_{i}\right\rVert^{2}+\mathbf{1}(i\in N_{j})\left\lVert\zeta_{j}\right\rVert^{2}\right)$ (AM-GM)
$=\sum_{i=1}^{n}\nu_{ni}\left(\sum_{j\in N_{i}}\nu_{nj}\right)\left\lVert\zeta_{i}\right\rVert^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\left\lVert\zeta_{i}\right\rVert^{2}$ (Lemma 1 (2.17))
$\leq\frac{\bar{\nu}^{2}}{n^{2}}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert\,\left\lVert\tilde{y}_{i}\right\rVert_{p}^{2}\,\left\lVert\psi_{i}\right\rVert_{q}^{2}$ (Lemma 2 (D.3))
Note that AM-GM stands for arithmetic-geometric mean inequality.
To obtain the uniform bound, we apply the max-$p$ norm, as defined in (3.3), and note that $d_{n}=n^{-1}\sum_{i=1}^{n}|N_{i}|$, yielding
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}}{n^{2}}\left(\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}\sum_{i=1}^{n}\left\lvert N_{i}\right\rvert=\frac{\bar{\nu}^{2}d_{n}}{n}\left(\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$
Since this inequality holds for any $p,q\in[1,\infty]$ satisfying $1/p+1/q=1/2$, we have a more conservative upper bound
$\sigma_{n}^{2}\leq\frac{\bar{\nu}^{2}d_{n}}{n}\left(\inf_{p,q\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}.$
Proof of Theorem 3.
The result follows directly from Proposition 4. Specifically, the bound in (3.5), together with Assumption 8 with $r=2$, implies that
$\mathbb{E}\left[\hat{\tau}_{n}(z,\omega)-\tau_{n}\right]^{2}\to 0,$
as $n\to\infty$, establishing mean square consistency.
Moreover, suppose $\sup_{n\in\mathbb{N}}d_{n}<\infty$, that is, there exists some $d_{0}\in(0,\infty)$ such that $d_{n}\leq d_{0}$ for all $n$. Since $\hat{\tau}_{n}-\tau_{n}$ follows some distribution with zero mean and variance $\sigma_{n}^{2}$, it follows that $\sqrt{n}(\hat{\tau}_{n}-\tau_{n})$ has variance $n\sigma_{n}^{2}$. By the bound in Proposition 4, we obtain
$n\sigma_{n}^{2}\leq\bar{\nu}^{2}d_{0}\left(\inf_{p,q\in S_{2}}\,\left\lVert\tilde{y}_{i}\right\rVert_{\max,p}^{n}\,\left\lVert\psi_{i}\right\rVert_{\max,q}^{n}\right)^{2}<\infty,$
under the stated assumptions. Hence, the aggregate Riesz estimator satisfies $O_{p}(n^{-1/2})$. ∎
Lemma 3.
Let $(X,d)$ be a metric space. Suppose $W_{n}$ and $Z$ are random variables taking values in $X$, and $d_{W}(W_{n},Z)\to 0$ as $n\to\infty$, where $d_{W}$ denotes the Wasserstein distance. Then $W_{n}\xrightarrow{d}Z$, i.e., $W_{n}$ converges in distribution to $Z$.
Proof.
We will show that $\mathbb{E}[f(W_{n})]\to\mathbb{E}[f(Z)]$ for all bounded uniformly continuous functions $f:X\to\mathbb{R}$, which characterises weak convergence; see Billingsley (1999, Theorem 2.1).
Let $f\in\mathrm{BUC}(X)$, the space of all bounded uniformly continuous functions. Since Lipschitz functions are dense in $\mathrm{BUC}(X)$ under the supremum norm, for any $\varepsilon>0$, there exists a Lipschitz function $f_{\varepsilon}:X\to\mathbb{R}$ such that
$\|f-f_{\varepsilon}\|_{\infty}<\frac{\varepsilon}{3}.$
The function $f_{\varepsilon}$ is Lipschitz with some finite constant $L_{\varepsilon}$. Therefore,
$\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|\leq L_{\varepsilon}\cdot d_{W}(W_{n},Z).$
Since $d_{W}(W_{n},Z)\to 0$, there exists $N(\varepsilon)\in\mathbb{N}$ such that
$\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|<\frac{\varepsilon}{3},\quad\text{for all }n>N(\varepsilon).$
Using the triangle inequality, we have
$\left|\mathbb{E}[f(W_{n})]-\mathbb{E}[f(Z)]\right|$ $\leq\left|\mathbb{E}[f(W_{n})]-\mathbb{E}[f_{\varepsilon}(W_{n})]\right|$
$\quad+\left|\mathbb{E}[f_{\varepsilon}(W_{n})]-\mathbb{E}[f_{\varepsilon}(Z)]\right|$
$\quad+\left|\mathbb{E}[f_{\varepsilon}(Z)]-\mathbb{E}[f(Z)]\right|$
$<\frac{\varepsilon}{3}+\frac{\varepsilon}{3}+\frac{\varepsilon}{3}=\varepsilon.$
Since $\varepsilon>0$ and $f\in\mathrm{BUC}(X)$ were arbitrary, we conclude that $\mathbb{E}[f(W_{n})]\to\mathbb{E}[f(Z)]$, i.e., $W_{n}\xrightarrow{d}Z$. ∎
Proof of Theorem 4.
The proof proceeds in two steps. First, we establish Wasserstein convergence for a sequence of probability measures. The key argument in this step is an application of Stein’s method (Stein, 1972), specifically Theorem 3.5 of Ross (2011) for a dependency-graph-based variant, following the strategy also employed in HWS. We then adapt the norm-bounding technique of HWS to our stochastic setting, where both the model space and the structure of dependence differ. Second, we show that Wasserstein convergence implies weak convergence of the same sequence.
Clearly $\mathbb{E}[\hat{\theta}_{i}(z,\omega)]^{4}<\infty$ is equivalent to $\mathbb{E}[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})]^{4}<\infty$. Define $X_{i}=\nu_{ni}(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i}))$, and then $\mathbb{E}[X_{i}^{4}]<\infty$. By Theorem 2, we have also $\mathbb{E}[X_{i}]=0$.
Note that $\sum_{i=1}^{n}X_{i}=\hat{\tau}_{n}(z,\omega)-\tau_{n}$. Define $W_{n}=\sum_{i=1}^{n}X_{i}/\sigma_{n}$. Since the collection $\{X_{1},X_{2},...,X_{n}\}$ have dependency neighborhoods $N_{i}$, i=1,,ni=1,...,n, with $D_{n}=\max_{1\leq i\leq n}|N_{i}|$. Then we can now apply Theorem 3.5 in Ross (2011) and obtain
$d_{W}(W_{n},Z)\leq\frac{D_{n}^{2}}{\sigma_{n}^{3}}\sum_{i=1}^{n}\mathbb{E}|X_{i}|^{3}+\frac{\sqrt{26}D_{n}^{3/2}}{\sqrt{\pi}\sigma_{n}^{2}}\sqrt{\sum_{i=1}^{n}\mathbb{E}[X_{i}^{4}]},$ (D.4)
where $d_{W}(W_{n},Z)$ denotes the Wasserstein distance between $W_{n}$ and a standard normal one $Z$. Observe that (D.4) gives an upper bound of the distance.
First, we examine the first term in (D.4)
$\sum_{i=1}^{n}\mathbb{E}\left\lvert X_{i}\right\rvert^{3}$ $=\sum_{i=1}^{n}\mathbb{E}\left\lvert\nu_{ni}\left(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right)\right\rvert^{3}$
$=\sum_{i=1}^{n}\nu_{ni}^{3}\mathbb{E}\left\lvert\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right\rvert^{3}$
$\leq\sqrt{\sum_{i=1}^{n}\nu_{ni}^{6}}\cdot\sqrt{\sum_{i=1}^{n}\left(\mathbb{E}\left\lvert\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right\rvert^{3}\right)^{2}}$ (Cauchy–Schwartz)
$\leq n^{-5/2}\bar{\nu}^{3}\cdot\sqrt{\sum_{i=1}^{n}\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p_{1}}\,\left\lVert\psi_{i}\right\rVert_{q_{1}}\right)^{6}}$ (Lemma 1 and 2)
$\leq n^{-2}\bar{\nu}^{3}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}\right)^{3}.$
Next, we examine the second term in (D.4).
$\sum_{i=1}^{n}\mathbb{E}[X_{i}]^{4}$ $=\sum_{i=1}^{n}\mathbb{E}\left[\nu_{ni}\left(\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right)\right]^{4}$
$=\sum_{i=1}^{n}\nu_{ni}^{4}\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{4}$
$\leq\sqrt{\sum_{i=1}^{n}\nu_{ni}^{8}}\cdot\sqrt{\sum_{i=1}^{n}\left(\mathbb{E}\left[\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})\right]^{4}\right)^{2}}$ (Cauchy–Schwartz)
$\leq n^{-7/2}\bar{\nu}^{4}\cdot\sqrt{\sum_{i=1}^{n}\left(2\left\lVert\tilde{y}_{i}\right\rVert_{p_{2}}\,\left\lVert\psi_{i}\right\rVert_{q_{2}}\right)^{8}}$ (Lemma 1 and 2)
$\leq n^{-3}\bar{\nu}^{4}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}\right)^{4}.$
Thus, we obtain
$d_{W}(W_{n},Z)$ $\leq\frac{\bar{\nu}^{3}D_{n}^{2}}{n^{2}\sigma_{n}^{3}}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}\right)^{3}+\frac{\sqrt{26}\bar{\nu}^{2}D_{n}^{3/2}}{\sqrt{\pi}n^{3/2}\sigma_{n}^{2}}\left(2\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}\,\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}\right)^{2}$
$=O\left(\frac{D_{n}^{2}}{n^{2}\sigma_{n}^{3}}\right)+O\left(\frac{D_{n}^{3/2}}{n^{3/2}\sigma_{n}^{2}}\right),$ (D.5)
under Assumption 8, which states that $\left\lVert\tilde{y}\right\rVert_{\max,p_{1}}^{n}$, $\left\lVert\psi\right\rVert_{\max,q_{1}}^{n}$, $\left\lVert\tilde{y}\right\rVert_{\max,p_{2}}^{n}$, and $\left\lVert\psi\right\rVert_{\max,q_{2}}^{n}$ converge for some $p_{1},q_{1},p_{2},q_{2}\in[1,\infty]$ satisfying $1/p_{1}+1/q_{1}=1/3$ and $1/p_{2}+1/q_{2}=1/4$.
Given Assumption 9, (D.5) can be further bounded by
$d_{W}(W_{n},Z)\leq O\left(\frac{D_{n}^{2}}{n^{1/2}}\right)+O\left(\frac{D_{n}^{3/2}}{n^{1/2}}\right),$ (D.6)
which implies that, if $D_{n}=o(n^{1/4})$, then $d_{W}(W_{n},Z)\to 0$.
Lemma 3 shows that Wasserstein convergence implies weak convergence, which completes the proof. ∎
Lemma 4.
Let $\{\zeta_{i}\}_{i=1}^{n}$ be random variables such that
$\mathbb{E}[\zeta_{i}]=0\quad\text{and}\quad\sup_{i\in\mathbb{N}}\mathbb{E}[\zeta_{i}^{4}]<\infty.$
For deterministic weights $\nu_{ni}$ satisfying Assumption 5, define
$\Delta_{n}\;=\;n\sum_{(i,j)\in\mathcal{E}_{n}}\Bigl(\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}-\nu_{ni}\nu_{nj}\,\mathbb{E}[\zeta_{i}\zeta_{j}]\Bigr).$ (D.7)
Then
$\Delta_{n}\;=\;O_{p}\bigl(n^{-1/2}\,D_{n}^{3/2}\bigr).$ (D.8)
Proof.
Set
$X_{ij}\;:=\;n\,\nu_{ni}\nu_{nj}\bigl(\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}]\bigr),\qquad(i,j)\in\mathcal{E}_{n},$
so that $\Delta_{n}=\sum_{(i,j)\in\mathcal{E}_{n}}X_{ij}$ and
$\mathbb{V}[\Delta_{n}]=\sum_{(i,j),(l,k)\in\mathcal{E}_{n}}\operatorname{Cov}[X_{ij},X_{lk}].$ (D.9)
Note that $\operatorname{Cov}[X_{ij},X_{lk}]=0$ whenever $X_{ij}$ and $X_{lk}$ are independent, so the sum involves only a restricted number of dependent pairs.
Since $\nu_{ni}\leq\bar{\nu}/n$, we have $n\,\nu_{ni}\nu_{nj}\leq\bar{\nu}^{2}/n$. Then, for any dependent $(i,j),(l,k)\in\mathcal{E}_{n}$, we apply Cauchy–Schwarz and the uniform fourth-moment bound $\sup_{i}\mathbb{E}[\zeta_{i}^{4}]<\infty$ to obtain
$\operatorname{Cov}[X_{ij},X_{lk}]\leq\left(\frac{\bar{\nu}^{2}}{n}\right)^{2}\cdot\mathbb{E}\Big|\big(\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}]\big)\big(\zeta_{l}\zeta_{k}-\mathbb{E}[\zeta_{l}\zeta_{k}]\big)\Big|\leq\frac{\bar{\nu}^{4}C}{n^{2}}$ (D.10)
for some constant $C>0$. This bound also applies when $(i,j)=(l,k)$.
Fix $(i,j)\in\mathcal{E}_{n}$. A second pair $(k,\ell)$ can be dependent with $(i,j)$ only if it lies within the corresponding neighbourhoods, that is, $(k,\ell)\cap(N_{i}\cup N_{j})\neq\emptyset$. Each unit has at most $D_{n}$ neighbours, so there are at most $D_{n}^{2}-1=O(D_{n}^{2})$ such pairs. Hence every $X_{ij}$ is dependent with at most $C_{1}D_{n}^{2}$ other summands for some constant $C_{1}>0$.
There are at most $nD_{n}$ ordered pairs in $\mathcal{E}_{n}$, so
$\mathbb{V}[\Delta_{n}]\;\leq\;C_{1}D_{n}^{2}\cdot nD_{n}\cdot\frac{\bar{\nu}^{4}C}{n^{2}}\;=\;C_{2}\,\bar{\nu}^{4}\,\frac{D_{n}^{3}}{n},$ (D.11)
for a constant $C_{2}>0$.
By Chebyshev’s inequality, for any $\varepsilon>0$,
$\mathbb{P}\bigl(|\Delta_{n}|>\varepsilon\bigr)\;\leq\;\frac{\operatorname{Var}(\Delta_{n})}{\varepsilon^{2}}\;\leq\;\frac{C_{2}\,\bar{\nu}^{4}}{\varepsilon^{2}}\,\frac{D_{n}^{3}}{n}.$
Since the right-hand side is of order $n^{-1}\,D_{n}^{3},$ we have
$\Delta_{n}\;=\;O_{p}\bigl(n^{-1/2}\,D_{n}^{3/2}\bigr),$
which completes the proof. ∎
Proof of Theorem 5.
By Theorem 2, we have $\mathbb{E}[\zeta_{i}]=0$, where $\zeta_{i}:=\hat{\theta}_{i}(z,\omega)-\theta_{i}(\tilde{y}_{i})$. Assumption 10(b)(b) ensures that $\sup_{i\in\mathbb{N}}\mathbb{E}[\hat{\theta}_{i}(z,\omega)^{4}]<\infty$, which is equivalent to $\sup_{i\in\mathbb{N}}\mathbb{E}[\zeta_{i}^{4}]<\infty$, since $\theta_{i}(\tilde{y}_{i})$ is constant.
Lemma 4 together with Assumption 5 implies that
$n\hat{\sigma}_{n}^{2}-n\sigma_{n}^{2}=n\sum_{(i,j)\in\mathcal{E}_{n}}\nu_{ni}\nu_{nj}\left(\zeta_{i}\zeta_{j}-\mathbb{E}[\zeta_{i}\zeta_{j}]\right)=O_{p}(n^{-1/2}D_{n}^{3/2}).$
If Assumption 6 holds with $D_{n}=o(n^{1/3})$, (D.11) implies that this expression converges to zero in mean square. ∎
Proof of Corollary 1.
The result follows directly from Theorems 4 and 5. We write
$\hat{\sigma}_{n}^{-1}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)=\frac{\sqrt{n}\left(\hat{\tau}_{n}(z,\omega)-\tau_{n}\right)}{\sqrt{n}\sigma_{n}}\cdot\frac{\sqrt{n}\sigma_{n}}{\sqrt{n}\hat{\sigma}_{n}}.$
The first factor converges in distribution to $\mathcal{N}(0,1)$ by Theorem 4, and the second factor converges to 1 in probability by Theorem 5, together with Assumption 9, which guarantees that $\sqrt{n}\sigma_{n}$ is bounded away from zero.
Since the product of a sequence converging in distribution and another converging in probability (to a constant) also converges in distribution, Slutsky’s theorem yields the desired result. ∎
Lemma 5.
$\Delta_{n}^{c}\;=\;O_{p}\bigl(n^{-1/2}\,D_{n}^{3/2}\bigr)$.
Proof of Lemma 5.
Define
$\Delta_{n}^{c}\;=\;n\sum_{(i,j)\in\mathcal{E}_{n}^{c}}\left(\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}-\nu_{ni}\nu_{nj}\,\mathbb{E}[\zeta_{i}\zeta_{j}]\right),$ (D.12)
so that $n\hat{\sigma}_{cn}^{2}-n\sigma_{n}^{2}=\Delta_{n}^{c}$. To bound the variance of $\Delta_{n}^{c}$, observe that
$\mathbb{V}[\Delta_{n}^{c}]=\sum_{(i,j),(l,k)\in\mathcal{E}_{n}^{c}}\operatorname{Cov}[X_{ij},X_{lk}].$ (D.13)
Since $\mathcal{E}_{n}^{c}\subseteq\mathcal{E}_{n}$,
$\mathbb{V}[\Delta_{n}]-\mathbb{V}[\Delta_{n}^{c}]=\sum_{(i,j),(l,k)\in\mathcal{E}_{n}/\mathcal{E}_{n}^{c}}\operatorname{Cov}[X_{ij},X_{lk}].$
This difference can be positive when $\sum_{(i,j),(l,k)\in\mathcal{E}_{n}/\mathcal{E}_{n}^{c}}\operatorname{Cov}[X_{ij},X_{lk}]\geq 0$, but this cannot be guaranteed in general, as some of the covariance terms may be negative.
Nonetheless, we can assert that the number of terms in the upper bound of $\mathbb{V}[\Delta_{n}^{c}]$ is no greater than $C_{1}D_{n}^{2}$, and that the bound in (D.11) applies as well, since it is derived by summing the absolute values of the covariances. Therefore,
$\Delta_{n}^{c}\;=\;O_{p}\bigl(n^{-1/2}\,D_{n}^{3/2}\bigr),$
as claimed. ∎
Proof of Theorem 6.
By Lemma 5, $\Delta_{n}^{c}=O_{p}\left(n^{-1/2}D_{n}^{3/2}\right)$. In particular, if Assumption 6 holds with $D_{n}=o(n^{1/3})$, then $\hat{\sigma}_{cn}^{2}-\sigma_{n}^{2}$ goes to zero in mean square. ∎
Proof of Corollary 2.
The proof follows the same argument as in Corollary 1, replacing $\hat{\sigma}_{n}^{2}$ with $\hat{\sigma}_{cn}^{2}$. ∎
Proof of Theorem 7.
Define
$\tilde{\Delta}_{n}\;=\;n\sum_{(i,j)\in\tilde{\mathcal{E}}_{n}}\left(\nu_{ni}\nu_{nj}\zeta_{i}\zeta_{j}-\nu_{ni}\nu_{nj}\,\mathbb{E}[\zeta_{i}\zeta_{j}]\right).$ (D.14)
Obviously, by Assumption 11, the upper bound of $\mathbb{V}[\tilde{\Delta}_{n}]$ lies between those of $\mathbb{V}[\Delta_{n}^{c}]$ and $\mathbb{V}[\Delta_{n}]$, and the result follows directly. ∎
Proof of Corollary 3.
The proof follows the same argument as in Corollary 1, replacing $\hat{\sigma}_{n}^{2}$ with $\tilde{\sigma}_{n}^{2}$. ∎